### v1

# Assessing (im)balance in signed brain networks

Abstract
Many complex systems - be they financial, natural or social - are composed by units - such as stocks, neurons or agents - whose joint activity can be represented as a multivariate time series. An issue of both practical and theoretical importance concerns the possibility of inferring the presence of a static relationships between any two units solely from their dynamic state. The present contribution aims at providing an answer within the frame of traditional hypothesis testing. Briefly speaking, our suggestion is that of linking any two units if behaving in a sufficiently similar way. To achieve such a goal, we project a multivariate time series onto a signed graph, by i) comparing the empirical properties of the former with those expected under a suitable benchmark and ii) linking any two units with a positive (negative) edge in case the corresponding series share a significantly large number of concordant (discordant) values. To define our benchmarks, we adopt an information-theoretic approach that is rooted into the constrained maximisation of Shannon entropy, a procedure inducing an ensemble of multivariate time series that preserves some of the empirical properties on average while randomising everything else. We showcase the possible applications of our method by addressing one of the most timely issues in the domain of neurosciences, i.e. that of determining if brain networks are frustrated or not - and, in case, to what extent. As our results suggest, this is indeed the case, the structure of the negative subgraph being more prone to inter-subject variability than the complementary, positive subgraph. At the mesoscopic level, instead, the minimisation of the Bayesian Information Criterion instantiated with the Signed Stochastic Block Model reveals that brain areas gather into modules aligning with the statistical variant of the Relaxed Balance Theory.
pacs:
89.75.Fb; 02.50.Tt
I Introduction
Time series are ubiquitous and versatile representations of dynamical complex systems, be they financial, natural or social. A major challenge is that of inferring the underlying dependencies between units from the (often non-linear) dynamics characterising them. Being capable of transforming a time series into a symbolic sequence, while preserving essential information about the ‘original’ generative process, is of practical relevance, as the dynamical properties of the former can be revealed by focusing on the structural properties of the latter Onnela et al. (2004); Sinha et al. (2010); Bouchaud and Potters (2003). To this aim, sequential data are often represented as traditional graphs, with nodes replacing individual units (such as stocks, neurons or agents) and edges reflecting their activity-driven interactions.
This choice plays a major role in the neuroscientific domain, where network-based approaches have revolutionised our understanding of the brain functioning, especially through the study of functional connectivity networks derived from Blood Oxygen Level Dependent (BOLD) fMRI signals. Rooted in graph theory and statistical physics, this perspective, often referred to as network neuroscience, aims at revealing the organisational principles of cognition and behaviour Kulkarni and Bassett (2025) by modelling the brain as an integrated system of interacting regions Bassett and Sporns (2017).
A review of the approaches that have been proposed so far to extract relationships between nodes directly from the time series of the original units is provided below.
I.1 From time series to proximity and from proximity to graphs
Given a set of time series, the most common approach is that of extracting a proximity matrix from it and, then, a graph from the proximity matrix. Although several definitions of proximity have been proposed [^1] We redirect the interested reader to the reviews Cliff et al. (2023); Liu et al. (2024), where hundreds of proximity measures (such as the dynamic time warping, the mutual information and the transfer entropy) are discussed and compared., the commonest choice is identifying it with the notion of pairwise correlation, as returned by the calculation of the Pearson correlation coefficient
$C_{ij}=\frac{\sum_{t=1}^{T}(X_{it}-\underline{X}_{i})\cdot(X_{jt}-\underline{X}_{j})}{\sqrt{\sum_{t=1}^{T}(X_{it}-\underline{X}_{i})^{2}}\cdot\sqrt{\sum_{t=1}^{T}(X_{jt}-\underline{X}_{j})^{2}}},$ (1)
where
$\underline{X}_{i}=\frac{\sum_{t=1}^{T}X_{it}}{T}\quad\text{and}\quad\underline{X}_{j}=\frac{\sum_{t=1}^{T}X_{jt}}{T}$ (2)
represent the average value of series $\mathbf{X}_{i}=\{X_{it}\}_{t=1}^{T}$ and $\mathbf{X}_{j}=\{X_{jt}\}_{t=1}^{T}$, respectively.
A less popular, but still worth to be mentioned, definition of proximity is the one resting upon the notion of partial correlation Hampson et al. (2002); Fransson and Marrelec (2008). If we indicate the matrix of pairwise correlations with $\mathbf{C}$, the generic entry of the matrix of partial correlations, then, reads $\hat{J}_{ik}=J_{ik}/\sqrt{J_{ii}\cdot J_{kk}}$, with $\mathbf{J}=\mathbf{C}^{-1}$ being the so-called precision matrix. Estimations of the functional connectivity (FC) based on partial correlations are affected by a smaller inter-subject variability Brier et al. (2015); Rahim et al. (2019), as they have been shown to resemble more closely the so-called structural connectivity (SC) patterns Huang et al. (2010); Salvador et al. (2005); Lee et al. (2011); Santucci et al. (2025).
I.1.1 Fixed thresholding algorithms
The simplest approach of this kind is based on the interpretation of a correlation matrix (be it partial or not) as a fully connected weighted matrix. In order to obtain non-trivial structures within such a context, sparsification procedures become, thus, necessary, the simplest one being that of thresholding correlations Bullmore and Sporns (2009). This procedure ultimately boils down to remove the weaker connections - usually the most affected by experimental noise Rubinov and Sporns (2010) - while retaining the stronger ones.
In the neuroscientific literature, one of the earliest attempts of this kind was the one pursued in Stam (2004), where measures of synchronisation between brain regions, derived from magnetoencephalographic (MEG) recordings, were considered to extract frequency band-specific networks by fixing their average degree; a similar attempt was pursued in Micheloyannis et al. (2009), where electroencephalographic (EEG) recordings were, instead, considered.
In the financial literature, the network resulting from the application of a threshold is also known with the name of asset graph Sinha et al. (2010); Onnela et al. (2004); Heimo et al. (2009). In Onnela et al. (2004), the authors derived an asset graph from the correlation matrix of a set of stocks by filling an empty graph, edge by edge, starting from the one corresponding to the largest correlation and stopping at the $N$-th largest correlation (with $N$ being the number of nodes). Upon doing so, the considered asset graph is comparable in size with the most popular asset tree, although being less constrained by requests such as the one of forbidding cycles.
As a last comment, we report the concern expressed by the authors of MacMahon and Garlaschelli (2015), according to whom thresholding may prevent the detection of a modular organisation, as the correlations within modules may be cut out even if significantly stronger than those between modules.
Since letting the network density drive the thresholding procedure appears as arbitrary as setting a threshold directly on correlations, it has been proposed to determine it by requiring that the resulting network remains connected. Such a data-driven procedure removes the weaker correlations until the percolation threshold is reached, i.e. the value in correspondence of which the resulting network disconnects Alexander-Bloch et al. (2010); Gallos et al. (2012). Although this kind of analysis has been popularised with the name of percolation analysis, the aforementioned value is, technically speaking, the connectivity threshold.
A different, yet related, criterion prescribes to adopt the threshold that maximises the value $\text{APL}_{\text{ER}}=\ln N/\ln\overline{k}$, with $\overline{k}=\sum_{i=1}^{N}k_{i}/N$, i.e. the (expected) average path length of an Erdös-Rényi network whose connectance matches the empirical one Hayasaka and Laurienti (2010); Marangolo et al. (2016); alternatively, one may adopt the threshold [^2] As the authors themselves explicitly state, their criteria individuate a whole range of link densities, more precisely from 0.4$\simeq$ 0.4≃ 0.4 to 0.5$\simeq$ 0.5≃ 0.5: below 0.4$\simeq$ 0.4≃ 0.4 graphs fragment and above 0.5$\simeq$ 0.5≃ 0.5 their topology becomes increasingly random. induced by the request of having a small-world network with no isolated nodes Bassett et al. (2008); Lynall et al. (2010).
If the previous approaches prescribe to stop pruning correlations once the percolation threshold is reached, a second approach focuses on the mesoscopic organisation of the network and prescribes to consider a correlation-induced distance333 Examples are provided the one reading $d_{ij}=\sqrt{2(1-C_{ij})}$, mapping distances between time series into the interval [0,2][0,2][ 0 , 2 ], and the one reading $d_{ij}=\sqrt{(1-C_{ij})/2}$, mapping distances between time series into the interval [0,1][0,1][ 0 , 1 ]. In the first case, 0 corresponds to linearly correlated time series, 2 corresponds to anti-correlated time series and $\sqrt{2}$ corresponds to the case where a linear correlation is not found; in the second case, 0 corresponds to linearly correlated time series, 1 corresponds to anti-correlated time series and $\sqrt{2}/2$ corresponds to the case where a linear correlation is not found., draw the corresponding dendrogram [^4] A dendrogram is constructed by clustering objects in a hierarchical fashion: to this aim, a distance between clusters is needed. Some of these ‘linkages’ algorithms induce an ultra-metric distance on the data Rammal et al. (1986) that has been argued to induce distances progressively less similar to the original ones as higher levels of the taxonomic tree are climbed: according to this intuition, such a method is more reliable to determine the low-level structure of the tree (induced by the strongest correlations) than the high-level one (induced by the weakest correlations) MacMahon and Garlaschelli (2015); Milligan (1979). and cut it in correspondence of a properly-defined threshold Mantegna and Stanley (1999); Tola et al. (2008). A possible criterion to individuate it is that of maximising the modular character of the resulting network structure, analysed upon running one of the many algorithms that have been proposed to this aim (Louvain, Infomap, Surprise, etc.) Newman (2006); Bordier et al. (2017); Marchese et al. (2022).
Examples from the financial literature are provided by Di Matteo et al. (2004), whose authors cut the dendrogram induced by 343434 weekly time series of interest rates at a distance which is halfway between the one induced by $C_{ij}=0$ and the one induced by $C_{ij}=1$, thus recovering 666 clusters throughout the whole period considered there, i.e. 198219821982-199719971997, and Tola et al. (2008), whose authors compare the performance of a similar clustering algorithm with that of an algorithm based on Random Matrix Theory in singling out groups of assets with comparable returns.
I.1.2 Progressive thresholding algorithms
Since individuating a unique threshold rises doubts about the arbitrariness of such a choice - which may significantly impact the results - alternatives have been proposed to make this approach a bit more principled, such as monitoring the changes in the resulting network structure upon rising the threshold itself. The earliest attempt, pursued in Eguíluz et al. (2005), consisted in identifying three threshold values, i.e. 0.60.60.6, 0.70.70.7 and 0.80.80.8: topological traits (averaged over the three sets of threshold-specific configurations) such as the scale-freeness of the distribution of functional connections and the small-worldness of the resulting network structure were, however, found to be robust against progressive thresholding. In Honey et al. (2009), the authors investigated how the relationships between FC and SC evolve upon progressively removing all FC values, the aim being that of quantifying the extent to which the SC values can be inferred from the ‘surviving’ FC ones. Studying the progressive engagement of brain regions has been proven useful to reveal the hierarchical character of the mouse brain Bardella et al. (2016) as well as distinguish healthy subjects from schizophrenic patients Mastrandrea et al. (2021).
I.1.3 Filtering algorithms
A different class of algorithms attempts to filter a correlation matrix by letting its values ‘dress’ peculiar graph topologies, supposedly isolating the most relevant correlations [^5] A threshold is, thus, still individuated, although only indirectly..
Filtering correlations via the MST. The most popular filtering approach returns the so-called Maximum Spanning Tree (MST), a technique retaining the $N-1$ largest correlations and ensuring that i) loops are discarded; ii) each node is reachable from any other node via a path that is the widest between its endpoints; iii) the total weight is greater than, or equal to, the total weight of any other spanning tree. It can be obtained upon running the Kruskal’s algorithm and is often employed to distinguish the connector areas (i.e. the most central ones) from the provincial areas (i.e. the most peripheral ones) Rammal et al. (1986); Mantegna and Stanley (1999); Caldarelli et al. (2012); Mastrandrea et al. (2017, 2021).
The Kruskal’s algorithm can be slightly modified by adding the extra condition that no edge between any two nodes that are already connected is allowed. Upon doing so, a Maximum Spanning Forest (MSF) is obtained, i.e. a set of disconnected trees usually interpreted as the backbone of a set of communities [^6] Notice that such a technique differs from the one prescribing to run the Kruskal’s algorithm on a disconnected graph, as it discards the edges that would join the different components of the MST, beside those generating loops..
Both the MST and the MSF are rather different from their random counterparts Bardella et al. (2016), their structural organisation providing a topological signature of the arising of diseases such as schizophrenia Mastrandrea et al. (2021). Coming to the financial literature, the authors of Bonanno et al. (2004) filter the correlation matrix induced by a set of daily time series of stock returns from the NYSE throughout the whole period considered there, i.e. 198719871987-199819981998, by calculating the Minimum Spanning Tree defined by the $N-1$ smallest correlation-induced distances (see footnote 3).
Filtering correlations via the PMFG. An alternative approach, discarding less information, returns the so-called Planar Maximally Filtered Graph (PMFG) Aste et al. (2005); Tumminello et al. (2005). The criterion informing such a method prescribes to retain the correlations ensuring that the resulting network is a planar graph, i.e. a graph that can be drawn with no links crossing each other (i.e. intersecting each other not in correspondence of a node). Since the PMFG always contains the MST, the former complements the latter by providing additional (and not just different) information. The PMFG has also been described as the simplest instance of a more general procedure embedding high-dimensional data into lower-dimensional manifolds having a controllable number of ‘holes’ called ‘genus’. The PMFG corresponds to the case in which the genus is zero Tumminello et al. (2005).
Although versatile - the method has been extended in a variety of ways, e.g. to produce a nested hierarchy of time series Song et al. (2011, 2012); Morales et al. (2014) - it suffers from some degree of arbitrariness too, affecting the properties of the postulated, approximating structure: the value of the genus must, in fact, be fixed a priori; besides, there is no obvious reason why time series should find a natural embedding into a bidimensional plane.
I.1.4 Testing correlations
A more principled approach to extract information from a correlation matrix is that of comparing (some of) its properties with the ones expected under a properly defined benchmark, the aim of this class of techniques being that of identifying non-random collective information about the set of time series induced by $\mathbf{C}$. Quite interestingly, these techniques do not discard negative correlations, known to be a consequence of spatial and temporal heterogeneities induced by non-trivial, neural-mediated, hemodynamic mechanisms and contributing to the stability of resting-state brain networks, by affecting the synchronisation of neural activity Saberi et al. (2021).
Although isolated attempts exist - such as the one proposed in Salvador et al. (2005), whose authors test the null hypothesis that the average value of the distribution of partial correlations, for each pair of areas and across subjects, is zero - methods of this kind can be split into those based on Random Matrix Theory (RMT) and those based on Entropy Maximization (EM).
Null models based on RMT. One of the most popular models studied in the context of RMT is the Wishart ensemble Wishart (1948); Livan et al. (2018) which, in its simplest form, leads to the Marcenko-Pastur distribution Marčenko and Pastur (1967), i.e. the distribution of the eigenvalues of a correlation matrix induced by $N$ completely random time series of duration $T$ in the limits $N\rightarrow+\infty$ and $T\rightarrow+\infty$ with $1<T/N<+\infty$, and reads
$\rho(\lambda)=\frac{T}{N}\cdot\frac{\sqrt{(\lambda_{+}-\lambda)(\lambda-\lambda_{-})}}{2\pi\lambda},\quad\lambda_{+}\leq\lambda\leq\lambda_{-}$ (3)
with $\sqrt{\lambda_{\pm}}=1\pm\sqrt{N/T}$ and $\rho(\lambda)=0$ otherwise. As a consequence, while the eigenvalues of an empirical correlation matrix falling within the range $[\lambda_{-},\lambda_{+}]$ can be imputable to random noise, any other eigenvalue can be deemed as representing meaningful structure in the data. As a consequence, the random component of any empirical correlation matrix can be identified with the tensor reading
$|\mathbf{C}^{(r)}=\sum_{i:\lambda_{i}\leq\lambda_{+}}\lambda_{i}\ket{v_{i}}\bra{v_{i}}$ | (4)
where |$\ket{v_{i}}$ is the eigenvector associated with the eigenvalue [^7] The eigenvalues smaller than $\lambda_{-}$ are usually also included. $\lambda_{i}$.
A typical feature of the spectrum of empirical correlation matrices is that of having an eigenvalue $\lambda_{m}$ that is (much) larger than any other eigenvalue and whose corresponding eigenvector |$\ket{v_{m}}$ has all positive components: such a ‘market mode’ contributes to the random component of an empirical correlation matrix $\mathbf{C}$ with a tensor reading $|\mathbf{C}^{(m)}=\lambda_{m}\ket{v_{m}}\bra{v_{m}}$ |.
The structured part of $\mathbf{C}$, reading
$\mathbf{C}^{(g)}=\mathbf{C}-\mathbf{C}^{(r)}-\mathbf{C}^{(m)}=\sum_{i:\lambda_{+}<\lambda_{i}<\lambda_{m}}\lambda_{i}\ket{v_{i}}\bra{v_{i}},$ (5)
is, then, inserted into the novel definition of modularity
$Q(\mathbf{\sigma})=\frac{1}{|\mathbf{C}|}\sum_{i=1}^{N}\sum_{j=1}^{N}C^{(g)}_{ij}\delta_{\sigma_{i}\sigma_{j}}$ (6)
with $|\mathbf{C}|=\sum_{i=1}^{N}\sum_{j=1}^{N}C_{ij}$ and whose maximisation can be proven to lead to clusters of nodes that are positively correlated while nodes belonging to different clusters are negatively correlated MacMahon and Garlaschelli (2015).
In Plerou et al. (2002), the authors employ models of this kind to analyse correlations of time series of stock market returns: by comparing them with null models based on RMT, the authors find that most eigenvalues are compatible with statistical noise, those capturing meaningful correlations (and being linked to economic factors such as market-wide movements or sectoral influences) being the few, large, deviating ones. In Utsugi et al. (2004), the authors find that the distribution of the eigenvalues of the correlation matrix induced by a set of daily time series of stock prices from the TSE is stable, throughout the whole period of time considered there, i.e. 199319931993-200120012001. In Potters et al. (2005), the authors explore how RMT can be applied to better estimate financial risk. In Zema et al. (2025), the authors build upon such a scheme by developing a novel approach for portfolio management.
Null models based on EM. A different, yet related, approach considers null models induced by entropy maximisation Masuda et al. (2018): in this case, one seeks to maximise the differential entropy
[…]
where $\mathbf{X}=\{X_{it}\}$ is an $N\times T$ (i.e. generally rectangular) table encoding information about $N$ standardised time series whose duration is $T$, subject to constraints concerning the empirical covariance matrix, now reading
$\mathbf{C}=\frac{\mathbf{X}\cdot\mathbf{X}^{T}}{T}$ (8)
or, more explicitly, $C_{ij}=\sum_{t=1}^{T}X_{it}X_{jt}/T$; such constraints are the values of the diagonal entries of $\mathbf{C}$, i.e.
$C_{ii}=\frac{\sum_{t=1}^{T}X_{it}^{2}}{T},\quad\forall\>i$ (9)
and its row sums, i.e.
$C_{i}=\sum_{j=1}^{N}C_{ij}=\frac{\sum_{j=1}^{N}\sum_{t=1}^{T}X_{it}X_{jt}}{T}=\frac{\sum_{t=1}^{T}X_{it}X_{t}}{T},\quad\forall\>i$ (10)
with $X_{t}=\sum_{j=1}^{N}X_{jt}$. The resulting distribution $P(\mathbf{X})$ is a multivariate normal distribution defined by a covariance matrix $\mathbf{C}^{(r)}$ obeying a Wishart distribution with $T$ degrees of freedom.
In Kojaku and Masuda (2019), the authors employ the same null model to filter covariance matrices, inducing networks in which any two nodes are linked if their correlation is found to be significantly larger than expected: to this aim, the matrix $\bm{\Delta}=\mathbf{C}-\mathbf{C}^{(r)}$, defined as the difference between the empirical and the random covariance matrix, is numerically determined by maximising the log-likelihood function [^8] To avoid overfitting, the log-likelihood is penalised by an adaptive-Lasso term whose parameters are determined by minimising the Bayesian Information Criterion (BIC). A package named ‘SCOLA’ (an acronym stating for ‘Sparse Networks Construction from Correlational Data with Lasso’) has been also released. $\mathcal{L}=\ln P(\mathbf{X}|\mathbf{C})$.
Bayesian inference. Another way for inferring a sparse graph from a set of time series, while avoiding thresholding, is that of inferring a precision matrix with $\ell_{1}$ penalization, where the value of the lasso regularisation parameter is usually set by cross-validation. This is equivalent to choosing the precision matrix that maximises the posterior distribution given the data, using a Gaussian likelihood and a prior encoding the $\ell_{1}$ constraint Ibáñez Berganza et al. (2023); Santucci et al. (2025); Varoquaux and Craddock (2013). Such an approach leads to a sparse, positive-definite precision matrix, interpretable as an adjacency matrix.
I.2 From univariate time series to graphs
Different (classes of) approaches exist that seek to extract a network from one or more time series while ignoring (the intermediate step represented by) the proximity matrix [^9] We redirect the interested reader to the reviews Silva et al. (2021); Zou et al. (2019), discussing the differences between the existing (classes of) algorithms and comparing their performances..
I.2.1 Visibility-induced mapping
According to such a mapping, the time steps of a univariate time series are interpreted as nodes to be connected if ‘visible’ to each other. Different types of visibility algorithms exist, designed to obtain undirected/directed, binary/weighted graphs.
[…]
A variant of the aforementioned procedure has been proposed in Luque et al. (2009), where the concept of horizontal visibility - as opposed to the one of natural visibility - is defined. A horizontal visibility graph (HVG) is constructed by considering (only) horizontal visibility lines: as a consequence, the HVG is always a subgraph of the VG, hence containing less information, although its construction is faster. HVGs have been employed to analyse seismic signals Telesca and Lovallo (2012), the volatility of returns of financial time series Zhang et al. (2015), heartbeat rates Xie et al. (2019) and for classification tasks Li et al. (2018).
A third variant is defined by implementing the concept of directed visibility, whose introduction has been justified by noticing that time has a natural direction: in Lacasa et al. (2012), the authors adopted the rule according to which any pair of time steps $t_{i}<t_{j}$ populates a directed horizontal visibility graph (DHVG) with an edge $i\rightarrow j$ and employed DHVGs to quantify time series irreversibility.
A fourth variant is defined by implementing the concept of limited penetrable visibility, proposed in Zhou et al. (2012) to reduce the level of noise accompanying data to a larger extent than that achievable by using VGs and HVGs. In this case, any two nodes are connected if the visibility line joining the tops of the two associated bars crosses a number of ‘obstacles’ that is smaller than, or equal to, a fixed value $L$ naturally, the limited penetrable visibility graph (LPVG) reduces to the VG when $L=0$, i.e. when there are no ‘obstacles’. LPVGs have been employed to analyse EEG signals of people in different conditions (affected by Alzheimer’s disease Wang et al. (2016) or under manual acupuncture Pei et al. (2014)).
A fifth variant is defined by implementing the concept of weighted visibility. The method, proposed in Supriya et al. (2016), prescribes to enrich each edge of a DHVG with a weight matching the ‘view angle’ of the observing node, its sign with respect to the horizontal axis indicating if the values of the time series increase (if positive) or decrease (if negative); since such a method allows for negative weights, the absolute value of the angle is often considered. The authors employed the method to identify EEG signals associated with epilepsy.
A last variant is defined by implementing the concept of differential visibility, realised by removing the edge set of the HVG from the edge set of the VG Zhu et al. (2014).
I.2.2 Transition-induced mapping
According to such a mapping, connections are established between (symbols assigned to) different portions of univariate time series, intended to represent states between which transitions occur: methods of this kind, in fact, represent a time series as a directed, weighted graph.
The mapping introduced in Campanharo et al. (2011) is based on a partition of the distribution of the values assumed by a time series in quantiles, to each of which a symbol is assigned - the number of quantiles is an arbitrary parameter whose choice should depend on the length of the series itself Shirazi et al. (2009). The variant of this method introduced in Weng et al. (2017) allows ‘memory networks’ to be constructed by assigning the attribute $t$ to the edge directed from $i$ to $j$ if the transition between the corresponding symbols occurs at time step $t$. A second variant is the one introduced in Campanharo and Ramos (2016), where the concept of renewal intervals, i.e. transitions between symbols posed at various temporal distances, is introduced.
The mapping introduced in Gao and Jin (2009) is named Coarse-Grained Phase Space Graphs and requires the definition of a time window $\omega$ to determine the number of values of the series to be represented as an $\omega$-dimensional vector, whose components are numerically determined by adopting a rule similar to the quantiles-based one.
A different, yet related, mapping is the one introduced in Small (2013), named Ordinal Partition Transition Networks. Here, the time series is embedded in a $\omega$-dimensional space where, starting from each point of the series, the current and the following $\omega-1$ numbers, interspaced by a time delay $\tau$, are considered (for instance, setting $\omega=3$ and $\tau=2$, the $i$-th embedding vector will contain the values at $i$, $i+2$, and $i+4$). The embedding vectors are, then, associated with patterns induced by the rank of their values and each ordinal pattern is represented as a node of the network: nodes $i$ and $j$ are connected if the corresponding patterns occur at subsequent time steps (i.e. if $t_{j}=t_{i}+1$).
A third variant of the same kind of mapping is the one introduced in Stephen et al. (2015) and named Visibility Graphlets Networks. This method requires the definition of a time window $\omega$ to determine the number of values to be considered: the time steps from $t$ to $t+\omega$ are, then, mapped into a directed visibility graph (DVG). Such a procedure is repeated for each time step and the DVGs having the same adjacency matrix are associated with the same symbol, i.e. one of the nodes of the resulting network. Different nodes are linked according to the subsequent occurrences of the related DVGs.
I.2.3 Proximity-induced mapping
According to such a mapping, connections are established between (groups of) time steps of a univariate time series, according to measures of distance/similarity between them.
The first method of this kind was introduced in Zhang et al. (2006) to represent pseudo-periodic time series. First, it segments a time series into cycles (e.g. dictated by a periodic behaviour) but without requiring them to have the same length; then, correlations between pairs of cycles are computed - by shifting the shortest one, if needed - and the largest correlation value is compared against a threshold: if found to be larger, an edge between the two cycles, represented as nodes in the resulting network, is established.
An alternative way of employing correlations to implement a proximity-induced mapping is introduced in Yang and Yang (2008) and prescribes to compute the correlations between the (either overlapping or non-overlapping) portions into which a time series can be partitioned: the nodes representing them are linked if the correlation between the corresponding portions exceeds a certain threshold.
A different class of methods is the one gathering the so-called Recurrence Networks (RNs) and is based on the idea that the points of a time series are recurrent if their distance, in a properly defined embedding space, is ‘short enough’. Similarly to the prescription characterising the Ordinal Partition Transition Networks algorithm, each time step is mapped into a point of an embedding space. Distances are, then, computed by using one of the many norms that can be defined. The way the embedding space is analysed distinguishes RNs in three sub-categories: i) the first one, introduced in Small et al. (2009), is composed by the so-called k-Nearest Neighbour Networks and requires to connect each node with the $k$ nodes whose corresponding points in the embedding space are closest to its corresponding point - to be noticed that such a variant has been defined for directed networks; ii) the second one, introduced in Xu et al. (2008), is composed by the so-called Adaptive Nearest Neighbour Networks and implements the undirected version of the aforementioned recipe; iii) the third one, introduced in Donner et al. (2010), is composed by the so-called $\epsilon$-Recurrence Networks ($\epsilon$-RNs) and prescribes to link the nodes whose corresponding points lie at a distance smaller than $\epsilon$: in symbols, the generic element of an $\epsilon$-RN matrix is defined as
$\mathbf{RN}_{tt^{\prime}}^{(x,x)}=\begin{cases}1,&\text{if\>}\left\|\bm{x}_{t}-\bm{x}_{t^{\prime}}\right\|\leq\epsilon\\ 0,&\text{otherwise}\end{cases}$ (11)
with $\bm{x}_{t}$ and $\bm{x}_{t^{\prime}}$ indicating the points of the chosen embedding space representing the values of the time series $X_{t}$ and $X_{t^{\prime}}$.
I.3 From multivariate time series to graphs
Methods of this kind aim at extracting a network from a multivariate time series. While some of them output single-layer networks, each time series being represented as a node, others output multi-layer networks, each time series being mapped into a different layer by employing any of the methods reviewed in the previous section.
I.3.1 Single-layer mapping
This first kind of mapping collects generalisations of the aforementioned RNs, hereby named Cross Recurrence Networks (CRNs), Inter-System Recurrence Networks (ISRNs) and Joint Recurrence Networks (JRNs) Feldhoff et al. (2012, 2013); Romano et al. (2004).
CRNs are defined for pairs of time series: after mapping the two time series $\mathbf{X}$ and $\mathbf{Y}$ into the same space, a biadjacency matrix is induced as follows:
$\mathbf{CRN}_{tt^{\prime}}^{(x,y)}=\begin{cases}1,&\text{if\>}\left\|\bm{x}_{t}-\bm{y}_{t^{\prime}}\right\|\leq\epsilon\\ 0,&\text{otherwise}\end{cases}$ (12)
with $\bm{x}_{t}$ and $\bm{y}_{t^{\prime}}$ indicating the points of the chosen embedding space representing the values of the time series $X_{t}$ and $Y_{t^{\prime}}$ notice that the two time series do not need to be of the same length, hence inducing a rectangular table Marwan et al. (2002); Zbilut et al. (1998).
Such a recipe can be further extended to deal with multivariate time series: in this case, we have the so-called ISRNs, inducing adjacency matrices like
$\mathbf{ISRN}=\begin{pmatrix}\mathbf{RN}^{(1,1)}&\mathbf{CRN}^{(1,2)}&\dots&\mathbf{CRN}^{(1,m)}\\ \mathbf{CRN}^{(2,1)}&\mathbf{RN}^{(2,2)}&\dots&\mathbf{CRN}^{(2,m)}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{CRN}^{(m,1)}&\mathbf{CRN}^{(m,2)}&\dots&\mathbf{RN}^{(m,m)}\end{pmatrix},$ (13)
constituted by (matrices representing the) RNs along the diagonal and by (matrices representing the) CRNs off the diagonal Feldhoff et al. (2012).
An alternative recipe is represented by the JRNs Romano et al. (2004); Feldhoff et al. (2013), inducing adjacency matrices like the one reading
$\mathbf{JRN}_{tt^{\prime}}^{(x,y)}=\begin{cases}1,&\text{if\>}\left\|\bm{x}_{t}-\bm{x}_{t^{\prime}}\right\|\leq\epsilon_{x}\cap\left\|\bm{y}_{t}-\bm{y}_{t^{\prime}}\right\|\leq\epsilon_{y}\\ 0,&\text{otherwise}\end{cases}$ (14)
that, similarly to what has been observed for the ISRNs, is block-wise, each diagonal block referring to a single time series and the off-diagonal blocks referring to pairs of time series.
A generalisation of the Ordinal Partition Transition Networks for multivariate time series was introduced in Zhang et al. (2017). The method requires a binarisation of the univariate time series, letting an increasing trend [^10] Defined as ‘the value of the series at time $t$ is larger than the value of the series at time $t-1$’. being represented with a 111 and a decreasing trend being represented with a 0: upon doing so, $2^{N}$ combinations, with $N$ being the number of time series, are possible for each time step, $t$; finally, each combination is assigned a node in the resulting network, any two nodes connected by a directed edge representing combinations found at subsequent steps. For deterministic and periodic time series, some of the nodes remain isolated; random processes are, instead, associated with densely (if not completely) connected graphs.
I.3.2 Multiple-layer mapping
This second kind of mapping, instead, associates a multivariate time series with a multi-layer network. A first example of such a mapping is provided by Lacasa et al. (2015): here, a multiplex visibility graph (MVG) is obtained by mapping each univariate time series into a graph by employing a VG; afterwards, the MVG is aggregated into a single-layer weighted graph, each node representing a layer-specific VG and each weight proxying the similarity between the corresponding layers. Such a method has been employed to analyse fMRI data Sannino et al. (2017).
The authors of Eroglu et al. (2018) follow the same pipeline but replace VGs with RNs, thus employing Multiplex Recurrence Networks (MRNs) to study paleoclimatic time series.
I.4 From multivariate time series to
validated graphs
The most recent approaches emphasise the importance of probabilistic (null, as well as generative) models to benchmark network features and disentangle fundamental, structural properties from statistical artifacts Varoquaux and Craddock (2013); Váša and Mišić (2022). These models play a crucial role in understanding whether empirical features (e.g. a modular structure, a small-world architecture, a rich-club organisation) are indicators of a genuine degree of self-organisation or simple by-products of lower-order constraints [^11] More explicitly, the practice of employing graph-theoretical estimations for inter-subject analyses without comparing them with a proper benchmark gives ‘[…] a fairly unspecific characterisation of the brain, being fragile to noise […] Another caveat is that […] correlation matrices display small-world properties […] by construction […] This observation highlights the need for well-defined null hypotheses […] but also for controlled recovery of brain functional connectivity going beyond empirical correlation matrices […]’ Varoquaux and Craddock (2013)..
In the present contribution, we focus on the issue of representing sequential data as relational data, by approaching the question how can a set of time series be represented as a traditional graph? from the perspective of traditional hypothesis testing. After comparing the statistical properties of the former against those expected under a suitable benchmark, our suggestion is that of projecting the set of time series onto a signed graph, by linking any two series with a positive (negative) edge in case they share a statistically significant number of concordant (discordant) values. More formally, we adopt an information-theoretic approach that is rooted into the constrained maximisation of Shannon entropy, a procedure inducing an ensemble of multivariate time series that preserves some of their empirical properties on average while randomising everything else Park and Newman (2004); Garlaschelli and Loffredo (2008); Squartini and Garlaschelli (2011).
Our method deals with signed data Gallo et al. (2024a, b, 2025): one may, thus, wonder how relevant is the information provided by signs within the neuroscientific context. Several studies have recently investigated the role played by the negative correlations characterising resting-state fMRI data (rs-fMRI): the authors of Saberi et al. (2021) show that negative correlations are essential for the stability of resting-state brain networks against perturbations or changes; the authors of Saberi et al. (2022) notice that the interplay between positive and negative correlations induces frustrated patterns preventing the brain from reaching a state of minimum energy and enhancing its adaptability to dynamical changes; finally, the authors of Demertzi et al. (2022) show that the default mode network (DMN) works ‘in opposition’ with the various task-positive networks, a supposedly crucial feature to maintain cognitive functions and consciousness.
Let us, however, stress that pre-processing procedures like the Global Signal Regression may artificially induce negative correlations, an evidence further stressing the importance of such a preliminary phase when cleaning neuroscientific data Murphy et al. (2009).
II Setting up the formalism
Let us start by setting up the formalism. Since a generic times series can be represented as $\mathbf{X}_{i}=\{X_{it}\}_{t=1}^{T}$, a set of $N$ time series whose duration is $T$ will be represented as an $N\times T$ rectangular matrix $\mathbf{X}=\{X_{it}\}$, whose generic entry $X_{it}$ represents the value of the $i$-th time series at time $t$. In what follows, we will standardise our time series and replace $X_{it}$ with
$w_{it}\equiv\frac{X_{it}-\underline{X}_{i}}{\text{std}[X_{i}]},\quad\forall\>i,t$ (15)
where
$\underline{X}_{i}=\frac{\sum_{t=1}^{T}X_{it}}{T}\quad\text{and}\quad\text{std}[X_{i}]=\sqrt{\frac{\sum_{t=1}^{T}(X_{it}-\underline{X}_{i})^{2}}{T}}$ (16)
are the average value and standard deviation of series $i$. Since $w_{it}$ can be either positive or negative, to ease mathematical manipulations let us employ Iverson’s brackets - a notation ensuring all quantities of interest to be non-negative Gallo et al. (2024a) - and define
$w_{it}^{-}$ $\equiv|w_{it}|\cdot[w_{it}<0],\quad\forall\>i,t$ (17)
$w_{it}^{+}$ $\equiv|w_{it}|\cdot[w_{it}>0],\quad\forall\>i,t$ (18)
i.e. mutually exclusive variables that induce the non-negative matrices $\mathbf{W}^{+}$ and $\mathbf{W}^{-}$ obeying
$w_{it}$ $=w_{it}^{+}-w_{it}^{-},\quad\forall\>i,t$ (19)
$|w_{it}|$ $=w_{it}^{+}+w_{it}^{-},\quad\forall\>i,t$ (20)
in an entry-wise fashion - or, more compactly, $\mathbf{W}\equiv\mathbf{W}^{+}-\mathbf{W}^{-}$ and $|\mathbf{W}|\equiv\mathbf{W}^{+}+\mathbf{W}^{-}$. Their purely binary counterparts read
$b_{it}^{-}$ $\equiv\left[w_{it}<0\right],\quad\forall\>i,t$ (21)
$b_{it}^{+}$ $\equiv\left[w_{it}>0\right],\quad\forall\>i,t$ (22)
in turn, inducing the non-negative matrices $\mathbf{B}^{+}$ and $\mathbf{B}^{-}$ that obey the analogous relationships
$b_{it}$ $=b_{it}^{+}-b_{it}^{-},\quad\forall\>i,t$ (23)
$|b_{it}|$ $=b_{it}^{+}+b_{it}^{-},\quad\forall\>i,t$ (24)
in an entry-wise fashion - or, more compactly, $\mathbf{B}\equiv\mathbf{B}^{+}-\mathbf{B}^{-}$ and $|\mathbf{B}|\equiv\mathbf{B}^{+}+\mathbf{B}^{-}$. The present analysis solely focuses on binarised time series: beside easing the mathematical tractability of the latter, binarisation has been also shown to lead to better clustering results Thomas Yeo et al. (2011).
Figure 1: Infographics illustrating our analysis pipeline. Pictorial representation of the pipeline we follow in the present contribution, from the registration of brain activity to the identification of communities in statistically validated signed projections: (a) rs-fMRI signals are recorded; (b) pairwise interactions between brain regions are considered; (c) the scalar product of each pair of (standardised and binarised) time series is calculated to quantify their signature; (d) its empirical value is validated against a null model to remove statistical noise; (e) a significantly large positive (negative) signature induces a positive (negative) link between the two involved brain regions; (f) a BIC-based community detection reveals the modules partitioning our statistically validated signed projections.
II.1 Local properties and dependencies of time series
In analogy with the quantities considered for the study of bipartite signed networks, let us define the total number of positive values as
$B^{+}\equiv\sum_{i=1}^{N}\sum_{t=1}^{T}b_{it}^{+}=\sum_{i=1}^{N}k_{i}^{+}=\sum_{t=1}^{T}\kappa_{t}^{+};$ (25)
the expression above suggests the definition of two (sets of) local positive quantities, i.e.
$k_{i}^{+}\equiv\sum_{t=1}^{T}b_{it}^{+},$ (26)
the positive degree of series $i$, capturing the total number of positive entries characterising the $i$-th series and providing information about its auto-correlation, and
$\kappa_{t}^{+}\equiv\sum_{i=1}^{N}b_{it}^{+},$ (27)
the positive degree at time $t$, capturing the total number of positive entries characterising the $t$-th time step and providing information about the correlations between the members of the whole set of time series; analogously,
$B^{-}\equiv\sum_{i=1}^{N}\sum_{t=1}^{T}b_{it}^{-}=\sum_{i=1}^{N}k_{i}^{-}=\sum_{t=1}^{T}\kappa_{t}^{-}$ (28)
with obvious meaning of the symbols. The advantage of adopting Iverson’s brackets is that of ensuring that each quantity is, now, computed on a matrix with positive entries, i.e. is positive as well.
II.2 Concordant and discordant temporal motifs
Let us, now, define the second-order properties known as dyadic motifs. In the binary case,
$B_{ij}^{++}$ $\equiv\sum_{t=1}^{T}b_{it}^{+}b_{jt}^{+}$ (29)
counts the number of time steps in correspondence of which both series $i$ and series $j$ are characterised by positive values and
$B_{ij}^{--}$ $\equiv\sum_{t=1}^{T}b_{it}^{-}b_{jt}^{-}$ (30)
counts the number of time steps in correspondence of which both series $i$ and series $j$ are characterised by negative values. It is quite intuitive to ascribe them to the class of concordant motifs, i.e. patterns capturing the ‘agreement’ between any two series. Analogously, it is quite intuitive to ascribe
$B_{ij}^{+-}$ $\equiv\sum_{t=1}^{T}b_{it}^{+}b_{jt}^{-}$ (31)
and
$B_{ij}^{-+}$ $\equiv\sum_{t=1}^{T}b_{it}^{-}b_{jt}^{+},$ (32)
counting the number of time steps in correspondence of which series $i$ is characterised by a positive value and series $j$ is characterised by a negative value or viceversa, to the class of discordant motifs, i.e. patterns capturing the ‘disagreement’ between any two series.
Any two nodes whose associated time series exhibit a significantly large number of concordant motifs will be connected by a +1+1+ 1, while any two nodes whose associated time series exhibit a significantly large number of discordant motifs will be connected by a 1-1- 1.
Lastly, we would like to stress that our projection scheme is applicable only in the absence of null values - a condition that is naturally satisfied by BOLD time series, representing the focus of this study.
II.3 A scheme for the validated projection of multivariate time series
As depicted in fig. 1, our validation algorithm works as follows:
  • $\mathbf{A.}$
    focus on a specific pair of series, say $i$ and $j$, and measure their (binary) similarity (see section III);
  • $\mathbf{B.}$
    quantify the statistical significance of the measured similarity, with respect to a properly-defined benchmark, by computing the corresponding p-value, say $p_{ij}$ (see section IV);
  • $\mathbf{C.}$
    repeat the step above for each pair of series;
  • $\mathbf{D.}$
    apply a multiple hypothesis testing procedure and insert a (binary) link between the nodes $i$ and $j$ if and only if $p_{ij}$ is found to be less than, or equal to, the threshold value $p_{th}$ (see section V).
Let us, now, describe each step in detail.
III Computing the similarity of any
two series
The first step of our method prescribes to measure the degree of similarity of series $i$ and $j$. To this aim, let us consider the quantity named signature Gallo et al. (2024a), defined as the scalar product of the two series: upon defining $\mathbf{b}_{i}$ as the $i$-th row of the matrix $\mathbf{B}$, the signature reads
$S_{ij}\equiv\mathbf{b}_{i}\cdot\mathbf{b}_{j}$ $=\sum_{t=1}^{T}b_{it}b_{jt}$
$=\sum_{t=1}^{T}(b_{it}^{+}-b_{it}^{-})(b_{jt}^{+}-b_{jt}^{-})$
$=\sum_{t=1}^{T}[(b_{it}^{+}b_{jt}^{+}+b_{it}^{-}b_{jt}^{-})-(b_{it}^{+}b_{jt}^{-}+b_{it}^{-}b_{jt}^{+})];$ (33)
in words, the signature is the difference between two quantities, i.e. the concordance of series $i$ and $j$, reading
$C_{ij}$ $\equiv\sum_{t=1}^{T}C_{ijt}\equiv\sum_{t=1}^{T}(b_{it}^{+}b_{jt}^{+}+b_{it}^{-}b_{jt}^{-})=B_{ij}^{++}+B_{ij}^{--}$ (34)
and counting the number of concordant motifs, and the discordance of nodes $i$ and $j$, reading
$D_{ij}$ $\equiv\sum_{t=1}^{T}D_{ijt}\equiv\sum_{t=1}^{T}(b_{it}^{+}b_{jt}^{-}+b_{it}^{-}b_{jt}^{+})=B_{ij}^{+-}+B_{ij}^{-+}$ (35)
and counting the number of discordant motifs: in symbols, $S_{ij}=C_{ij}-D_{ij}$, $\forall\>i<j$. A naïve way of projecting a multivariate binary time series would prescribe to stop here and pose
[…]
i.e. apply the sign function to the signature, hence connecting series $i$ and $j$ with a positive link if $S_{ij}>0$, i.e. $C_{ij}>D_{ij}$, and with a negative link if $S_{ij}<0$, i.e. $C_{ij}<D_{ij}$.
Let us also explicitly notice that $S_{ij}$ proxies the sign of the correlation coefficient between the standardised time series $i$ and $j$ but does not necessarily coincide with it: while, in fact, the generic entry of $\mathbf{C}$ in eq. 8 is a sum of addenda, $S_{ij}$ is the sum of the signs of those addenda.
IV Testing the similarity of any
two series
The second step of our method prescribes to evaluate the statistical significance of the similarity of any two series. To this aim, let us, first, notice that
$-T\leq S_{ij}\leq T$ (37)
where $S_{ij}=-T$ if $C_{ij}=0$ (i.e. if each dyadic motif is composed by a 1-1- 1 and a +1+1+ 1) and $S_{ij}=T$ if $D_{ij}=0$ (i.e. if each dyadic motif is composed by either two 1-1- 1s or two +1+1+ 1s) and, then, clarify that we will employ ERGs defined by linear constraints, hence treat links as independent random variables.
IV.1 Homogeneous benchmark
Let us start by considering the homogeneous benchmark defined by the finite scheme [^12] A mathematical construction consisting of the list of elementary events and the associated probabilities.
$b_{it}\sim\begin{pmatrix}-1&+1\\ p^{-}&p^{+}\end{pmatrix},\quad\forall\>i,t$ (38)
which, in turn, induces the finite scheme
$b_{it}b_{jt}$ $\sim\begin{pmatrix}-1&+1\\ 2p^{+}p^{-}&(p^{+})^{2}+(p^{-})^{2}\end{pmatrix}$
$\equiv\begin{pmatrix}-1&+1\\ q^{-}&q^{+}\end{pmatrix},\quad\forall\>i,j,t$ (39)
satisfying the relationships
1 11 $=q^{+}+q^{-},$ (40)
$\langle b_{it}b_{jt}\rangle$ $=q^{+}-q^{-},$ (41)
$\text{Var}[b_{it}b_{jt}]$ $=(q^{+}+q^{-})-(q^{+}-q^{-})^{2}$
$=1-(q^{+}-q^{-})^{2}$
$=4q^{+}q^{-}.$ (42)
Let us, now, notice that $S_{ij}$ is a sum of i.i.d. Bernoulli random variables [^13] The finite scheme obeyed by $C_{ijt}-D_{ijt}$ is the same finite scheme obeyed by $b_{it}b_{jt}$., the outcomes of each elementary event being, in fact, 1-1- 1 and +1+1+ 1 instead of 0 and +1+1+ 1. Such a consideration allows us to find the probability distribution obeyed by $S_{ij}$ quite straightforwardly, by starting from the one describing $C_{ij}$ alone, i.e.
$P(C_{ij}=k)=\binom{T}{k}(q^{+})^{k}(q^{-})^{T-k};$ (43)
let us, in fact, consider the change of variable $k\equiv(T+s)/2$, leading to the binomial
$P(S_{ij}=s)=\binom{T}{\frac{T+s}{2}}(q^{+})^{\frac{T+s}{2}}(q^{-})^{\frac{T-s}{2}}$ (44)
[…]
A comment about the parity of $T$ and $s$ is needed: let us notice that, in case $T$ is odd, $s$ itself must be odd, assuming the values $s=-T,-T+2,-T+4\dots T$ (corresponding to $k=0,1,2\dots T$); in case $T$ is even, $s$ itself must be even, assuming the same values as above in correspondence of the same values of $k$.
IV.2 Heterogeneous benchmark
Let us, now, consider the heterogeneous benchmark
$b_{it}\sim\begin{pmatrix}-1&+1\\ p_{it}^{-}&p_{it}^{+}\end{pmatrix},\quad\forall\>i,t$ (45)
which, in turn, induces the finite scheme
$b_{it}b_{jt}$ $\sim\begin{pmatrix}-1&+1\\ p_{it}^{+}p_{jt}^{-}+p_{it}^{-}p_{jt}^{+}&p_{it}^{+}p_{jt}^{+}+p_{it}^{-}p_{jt}^{-}\end{pmatrix}$
$\equiv\begin{pmatrix}-1&+1\\ q_{ijt}^{-}&q_{ijt}^{+}\end{pmatrix},\quad\forall\>i,j,t$ (46)
satisfying the relationships
1 11 $=q_{ijt}^{+}+q_{ijt}^{-},$ (47)
$\langle b_{it}b_{jt}\rangle$ $=q_{ijt}^{+}-q_{ijt}^{-},$ (48)
$\text{Var}[b_{it}b_{jt}]$ $=(q_{ijt}^{+}+q_{ijt}^{-})-(q_{ijt}^{+}-q_{ijt}^{-})^{2}$
$=1-(q_{ijt}^{+}-q_{ijt}^{-})^{2}$
$=4q_{ijt}^{+}q_{ijt}^{-}.$ (49)
Since $S_{ij}$ is, now, a sum of i.n.i.d. Bernoulli random variables, the probability distribution obeyed by it is the Poisson-binomial
$P(S_{ij}=s)$ $=\sum_{C_{k}}\left[\prod_{\mu\in C_{k}}q_{ij\mu}^{+}\prod_{\nu\notin C_{k}}q_{ij\nu}^{-}\right]$ (50)
where $C_{k}$ is the set of $k$-tuples of which $\mu$ and $\nu$ are instances. Let us provide some explicit examples:
$P(S_{ij}=-T)=\prod_{\alpha=1}^{T}q_{ij\alpha}^{-},$ (51)
$P(S_{ij}=2-T)=\sum_{\beta=1}^{T}\left[q_{ij\beta}^{+}\prod_{\begin{subarray}{c}\alpha=1\\ \alpha\neq\beta\end{subarray}}^{T}q_{ij\alpha}^{-}\right],$ (52)
$P(S_{ij}=4-T)=\sum_{\beta=1}^{T}\sum_{\begin{subarray}{c}\gamma=1\\ \gamma>\beta\end{subarray}}^{T}\left[q_{ij\beta}^{+}q_{ij\gamma}^{+}\prod_{\begin{subarray}{c}\alpha=1\\ \alpha\neq\beta,\gamma\end{subarray}}^{T}q_{ij\alpha}^{-}\right]$ (53)
and so on. In words, if nodes $i$ and $j$ establish $k=0$ concordant motifs, they establish $T$ discordant motifs and the signature reads $s=-T$; if nodes $i$ and $j$ establish $k=1$ concordant motif, they establish $T-1$ discordant motifs and the signature reads $s=2-T$; if nodes $i$ and $j$ establish $k=2$ concordant motif, they establish $T-2$ discordant motifs and the signature reads $s=4-T$. The same line of reasoning can be repeated for all the admissible values of $k$.
Figure 2: Infographics illustrating our validation procedure. Probability distribution of the signature and its Gaussian approximation: while the left panel provides a graphical answer to the question is the empirical value of the signature significantly different from the one expected under the chosen benchmark?, the right panel provides a graphical answer to the question is the deviation negative (hence, the signature is significantly smaller than expected) or positive (hence, the signature is significantly larger than expected)? The red area corresponds to the region of validation of the negative links, induced by pairs of series ‘in counterphase’ most of the time; the blue area corresponds to the region of validation of the positive links, induced by pairs of series ‘in phase’ most of the time.
IV.3 The statistical significance of similarity:
a two-sided test of hypothesis
Once we have calculated the distribution of the similarity for each pair of series, we must evaluate the statistical significance of the empirical signature: since we have a signed quantity, we need both tails of such a distribution to carry out a two-sided test of hypothesis. In fact, we need to answer the two related questions i) is the empirical value of the signature significantly different from the one expected under the chosen benchmark? and ii) if so, is the deviation negative (hence, the signature is significantly smaller than expected) or positive (hence, the signature is significantly larger than expected)?
The first question can be answered upon calculating the two-sided p-value reading
$p_{ij}=2\cdot\min\left\{F(S_{ij}^{*}),1-F(S_{ij}^{*})\right\}$ (54)
with $F$ being the cumulative distribution function, defined as $F(S_{ij}^{*})=\sum_{x\leq S_{ij}^{*}}P(S_{ij}=x)$: such a number evaluates the probability of observing a deviation from the expected value in either direction.
The second question can be answered upon calculating the sign of such a deviation, by determining if either
$F(S_{ij}^{*})<1-F(S_{ij}^{*})$ (55)
or
$F(S_{ij}^{*})>1-F(S_{ij}^{*})$ (56)
holds true. In the first case, $F(S_{ij}^{*})<1/2$, the empirical value of the signature is smaller than the median of the distribution and the deviation is negative; in the second case, $F(S_{ij}^{*})>1/2$, the empirical value of the signature is larger than the median of the distribution and the deviation is positive. Intuitively [^14] The value $p_{th}$ will be determined later, at the third step of our algorithm: hereby, a merely discursive introduction to the proper validation step is provided.,
  • $\bullet$
    the two conditions $F(S_{ij}^{*})<1/2$ and $p_{ij}\leq p_{th}$ indicate that nodes $i$ and $j$ have established a number of discordant motifs that is so large to induce a significantly negative signature - and, potentially, a negative link in the projection ($a_{ij}=-1$). In a sense, the two series are ‘in counterphase’ most of the time;
  • $\bullet$
    the two conditions $F(S_{ij}^{*})>1/2$ and $p_{ij}\leq p_{th}$ indicate that nodes $i$ and $j$ have established a number of concordant motifs that is so large to induce a significantly positive signature - and, potentially, a positive link in the projection ($a_{ij}=+1$). In a sense, the two series are ‘in phase’ most of the time;
  • $\bullet$
    the condition $p_{ij}>p_{th}$ individuates a value of the signature (i.e. of concordant/discordant motifs) that is compatible with the one predicted by the chosen benchmark: stated otherwise, the empirical number of motifs could have been observed in configurations generated by the benchmark itself - hence, inducing a null link in the projection ($a_{ij}=0$).
For a graphical representation of our validation procedure, see fig. 2.
V Validating the projection of a multivariate time series
The second step of our method returns a symmetric matrix of p-values. Individuating the ones associated with the hypotheses to be actually rejected requires a procedure to deal with the comparison of multiple hypotheses at the same time. In very general terms, a threshold must be set: if the specific p-value is smaller than it, the associated event is interpreted as statistically significant and the corresponding nodes are connected in the projection.
[…]
and, then, identify the largest integer $\hat{i}$ satisfying the condition
[…]
with $t$ representing the usual single-test significance level, hereby set to the value $t=0.05$: since $H_{i}$ is the hypothesis that the distribution of dyadic motifs established by the $i$-th pair of nodes follows a binomial/Poisson-binomial, rejecting it amounts at connecting the corresponding nodes in the projection. Notice that $|H|=N(N-1)/2$.
Naturally, deciding which test is the best suited one for the problem at hand depends on the importance assigned to false positives and false negatives: as a rule of thumb, the Bonferroni correction can be deemed as appropriate when few tests are expected to be significant (i.e. when even a single false positive would be problematic); when, on the contrary, many tests are expected to be significant, using the Bonferroni correction may, in turn, produce a too large number of false negatives.
VI Benchmarks for the analysis of multivariate time series
In order to evaluate the statistical significance of the similarity of any two series $i$ and $j$, a benchmark is needed: a natural choice leads to adopt the class of null models named Exponential Random Graphs (ERGs). Within such a framework, the constrained maximisation of Shannon entropy
$S=-\sum_{\mathbf{B}\in\mathbb{B}}P(\mathbf{B})\ln P(\mathbf{B})$ (59)
leads to assign the generic bipartite network $\mathbf{B}$ the probability
$P(\mathbf{B})=\frac{e^{-H(\bm{\theta},\bm{C}(\mathbf{B}))}}{Z(\bm{\theta})}$ (60)
whose value is determined by the vector $\bm{C}(\mathbf{B})$ of topological constraints via the Hamiltonian $H(\bm{\theta},\bm{C}(\mathbf{B}))=\bm{\theta}\cdot\bm{C}(\mathbf{B})=\sum_{i}\theta_{i}C_{i}(\mathbf{B})$. The use of linear constraints allows $P(\mathbf{B})$ to be written in a factorised form, i.e. as the product of pair-specific probability coefficients Gallo et al. (2025).
In order to determine the unknown parameters $\bm{\theta}$, the likelihood maximisation recipe can be adopted: given an observed biadjacency matrix $\mathbf{B}^{*}$, it translates into solving the system of equations
$\langle C_{i}\rangle(\bm{\theta})=\sum_{\mathbf{B}\in\mathbb{B}}P(\mathbf{B})C_{i}(\mathbf{B})=C_{i}(\mathbf{B}^{*}),\quad\forall\>i$ (61)
which prescribes to equate each ensemble average to its observed counterpart.
Let us, now, generalise such a framework to accommodate models for studying multivariate time series: to this aim, we will follow Marcaccioli and Livan (2020); Gallo et al. (2024a) which, in turn, builds upon the analytical approach introduced in Park and Newman (2004) and further developed in Squartini and Garlaschelli (2011).
Let us, now, make a general observation: as each of the $N$ time series is defined by $T$ time steps, in corresponding of which either the value 1-1- 1 or the value +1+1+ 1 can be assumed, the ensemble is constituted by $|\mathbb{B}|=2^{N\cdot T}$ (bi)adjacency matrices, representing all possible, binary, signed, multivariate time series.
VI.1 Signed Random Graph Model
The first null model we consider is induced by the Hamiltonian
$H(\mathbf{B})=\alpha\>B^{+}(\mathbf{B})+\gamma\>B^{-}(\mathbf{B})$ (62)
i.e. by the two global constraints $B^{+}(\mathbf{B})$ and $B^{-}(\mathbf{B})$. Although being formally analogous to the BiSRGM Gallo et al. (2024a), let us name it bSRGM.
According to the bSRGM, each entry of a bipartite signed network is a random variable whose behaviour is described by the coefficients
$P(b_{it}=+1)$ $=\frac{e^{-\alpha}}{e^{-\alpha}+e^{-\gamma}}\equiv p^{+},$ (63)
$P(b_{it}=-1)$ $=\frac{e^{-\gamma}}{e^{-\alpha}+e^{-\gamma}}\equiv p^{-};$ (64)
in other words, $b_{it}$ obeys a generalised Bernoulli distribution whose probability coefficients are determined by the (Lagrange multipliers of the) imposed constraints: each positive link appears with probability $p^{+}$ and each negative link appears with probability $p^{-}$.
In order to employ the bSRGM for studying real-world multivariate binary time series, the parameters that define it need to be properly tuned: more specifically, one needs to ensure
$\langle B^{+}\rangle_{\text{bSRGM}}$ $=B^{+}(\mathbf{B}^{*}),$ (65)
$\langle B^{-}\rangle_{\text{bSRGM}}$ $=B^{-}(\mathbf{B}^{*})$ (66)
with the symbol $\mathbf{B}^{*}$ indicating the empirical multivariate time series under analysis. To this aim, one can maximise the likelihood function $\mathcal{L}_{\text{bSRGM}}(\alpha,\gamma)\equiv\ln P_{\text{bSRGM}}(\mathbf{B}^{*}|\alpha,\gamma)$, where $P_{\text{bSRGM}}(\mathbf{B}^{*}|\alpha,\gamma)$ is defined as in eq. 111, with respect to the unknown parameters that define it Garlaschelli and Loffredo (2008). Such a recipe leads us to find
$p^{+}$ $=B^{+}(\mathbf{B}^{*})/(N\cdot T),$ (67)
$p^{-}$ $=B^{-}(\mathbf{B}^{*})/(N\cdot T)$ (68)
with obvious meaning of the symbols (see also Appendix A).
VI.2 Signed Configuration Model
The second null model we consider is induced by the Hamiltonian
$H(\mathbf{B})=$ $\sum_{i=1}^{N}[\alpha_{i}k_{i}^{+}(\mathbf{B})+\gamma_{i}k_{i}^{-}(\mathbf{B})]$
$+\sum_{t=1}^{T}[\delta_{t}\kappa_{t}^{+}(\mathbf{B})+\eta_{t}\kappa_{t}^{-}(\mathbf{B})];$ (69)
i.e. by the $2(N+T)$ local constraints $\{k_{i}^{+}(\mathbf{B})\}_{i=1}^{N}$, $\{k_{i}^{-}(\mathbf{B})\}_{i=1}^{N}$, $\{\kappa_{t}^{+}(\mathbf{B})\}_{t=1}^{T}$ and $\{\kappa_{t}^{-}(\mathbf{B})\}_{t=1}^{T}$. Although being formally analogous to the BiSCM Gallo et al. (2024a), let us name it bSCM.
According to the bSCM, each entry of a bipartite signed network is a random variable whose behaviour is described by the coefficients
$P(b_{it}=+1)$ $=\frac{e^{-(\alpha_{i}+\delta_{t})}}{e^{-(\alpha_{i}+\delta_{t})}+e^{-(\gamma_{i}+\eta_{t})}}\equiv p_{it}^{+},$ (70)
$P(b_{it}=-1)$ $=\frac{e^{-(\gamma_{i}+\eta_{t})}}{e^{-(\alpha_{i}+\delta_{t})}+e^{-(\gamma_{i}+\eta_{t})}}\equiv p_{it}^{-};$ (71)
in other words, $b_{it}$ obeys a generalised Bernoulli distribution whose probability coefficients are determined by the (Lagrange multipliers of the) imposed constraints: the series $i$, at time $t$ assumes the value +1+1+ 1 with probability $p_{it}^{+}$ and the value 1-1- 1 with probability $p_{it}^{-}$.
[…]
In order to tune the parameters defining the bSCM to ensure
$\langle k_{i}^{+}\rangle_{\text{bSCM}}$ $=k_{i}^{+}(\mathbf{B}^{*}),\quad\forall\>i,$ (72)
$\langle k_{i}^{-}\rangle_{\text{bSCM}}$ $=k_{i}^{-}(\mathbf{B}^{*}),\quad\forall\>i,$ (73)
$\langle\kappa_{t}^{+}\rangle_{\text{bSCM}}$ $=\kappa_{t}^{+}(\mathbf{B}^{*}),\quad\forall\>t,$ (74)
$\langle\kappa_{t}^{-}\rangle_{\text{bSCM}}$ $=\kappa_{t}^{-}(\mathbf{B}^{*}),\quad\forall\>t$ (75)
let us maximise the likelihood function $\mathcal{L}_{\text{bSCM}}(\bm{\alpha},\bm{\gamma},\bm{\delta},\bm{\eta})\equiv\ln P_{\text{bSCM}}(\mathbf{B}^{*}|\bm{\alpha},\bm{\gamma},\bm{\delta},\bm{\eta})$, where $P_{\text{bSCM}}(\mathbf{B}^{*}|\bm{\alpha},\bm{\gamma},\bm{\delta},\bm{\eta})$ is defined as in eq. A, with respect to the unknown parameters that define it Garlaschelli and Loffredo (2008) (see also Appendix A). Such a recipe leads us to find
$k_{i}^{+}(\mathbf{B}^{*})$ $=\sum_{t=1}^{T}\frac{e^{-(\alpha_{i}+\delta_{t})}}{e^{-(\alpha_{i}+\delta_{t})}+e^{-(\gamma_{i}+\eta_{t})}}=\langle k_{i}^{+}\rangle,\quad\forall\>i,$ (76)
$k_{i}^{-}(\mathbf{B}^{*})$ $=\sum_{t=1}^{T}\frac{e^{-(\gamma_{i}+\eta_{t})}}{e^{-(\alpha_{i}+\delta_{t})}+e^{-(\gamma_{i}+\eta_{t})}}=\langle k_{i}^{-}\rangle,\quad\forall\>i,$ (77)
$\kappa_{t}^{+}(\mathbf{B}^{*})$ $=\sum_{i=1}^{N}\frac{e^{-(\alpha_{i}+\delta_{t})}}{e^{-(\alpha_{i}+\delta_{t})}+e^{-(\gamma_{i}+\eta_{t})}}=\langle\kappa_{t}^{+}\rangle,\quad\forall\>t,$ (78)
$\kappa_{t}^{-}(\mathbf{B}^{*})$ $=\sum_{i=1}^{N}\frac{e^{-(\gamma_{i}+\eta_{t})}}{e^{-(\alpha_{i}+\delta_{t})}+e^{-(\gamma_{i}+\eta_{t})}}=\langle\kappa_{t}^{-}\rangle,\quad\forall\>t$ (79)
a system that can be solved only numerically, along the guidelines provided in Vallarano et al. (2021) (see also Appendix B).
VII Dataset description
[…]
Figure 4: Distributions of the signed, bipartite and monopartite, connectance. Left panel: distributions of the positive and negative bipartite connectance, respectively defined as $c^{+}=B^{+}/(N\cdot T)$ (in blue) and $c^{-}=B^{-}/(N\cdot T)$ (in red), across our 100100100 subjects. Middle and right panel: distributions of the positive and negative monopartite connectance, respectively defined as $\rho^{+}=2L^{+}/N(N-1)$ (in blue) and $\rho^{-}=2L^{-}/N(N-1)$ (in red), across the three, different projections of our 100100100 subjects. Kernel density estimation has been employed for better visualisation; bin widths are computed according to the Freedman-Diaconis rule. How can a large number of bipartite, negative links co-exist with the large number of positive links characterising the naïve projection? Bipartite, negative links are evidently arranged into concordant motifs that, in turn, induce a naïve projection populated by (a majority of) positive links. Notice that the naïve and bSCM-induced distributions overlap, in the right panel.
VIII Results
Let us, now, employ the proposed validation procedure to carry out a ‘binary-to-binary’ analysis, i.e. obtain statistically validated binary projections of multivariate binary time series, defined by the ‘rule’ that any two nodes sharing a significantly large number of concordant (discordant) motifs are connected by a positive (negative) edge (see also fig. 2).
VIII.1 Consistency checks
Let us start our analysis by verifying that the ensemble distribution of each, pair-specific signature is consistent with its analytical counterpart, for all benchmarks considered here.
[…]
VIII.2 Patterns of brain (im)balance at the microscale
[…]
VIII.2.1 Connectance
When considering the unsigned case, the connectance of the projection returned by the naïve approach is typically large Saracco et al. (2017): this holds true here as well, for all the considered subjects (see fig. 4; for a graphical example, see fig. 10). To be noticed that many validated links are positive, an evidence letting us conclude that the time series associated with the related pairs of areas are concordant for the vast majority of the snapshots constituting the considered time window.
[…]
Figure 5: Distributions of CV+, CV- and signed degrees. Top panels: box-plots summing up the distributions of the coefficient of variation for positive degrees (CV+, left) and negative degrees (CV-, right) across the three, different projections of our 100100100 subjects. Each box-plot illustrates the distribution of the corresponding metric, indicating its central tendency and dispersion, for each model: the positive variant of the CV suggests that the corresponding subgraphs are characterised by a quite homogeneous structure while the negative one suggests that the opposite holds true, being the corresponding subgraphs characterised by fat-tailed degree distributions. Middle and bottom panels: distributions of the positive (left) and negative (right) degrees populating our projections, for subjects 106016 (middle panels) and 857263 (bottom panels). Kernel density estimation has been employed for better visualisation; bin widths are computed according to the Freedman-Diaconis rule. While positive distributions appear (approximately) bell-shaped around the average, negative distributions are characterised by fat tails, an evidence suggesting the presence of hubs. Notice that the naïve and bSCM-induced distributions overlap, in the right panels.
[…]
The local, bSCM-based filter, instead, accounts for nodes heterogeneity: whenever the positive degree of a node is (much) larger than its negative degree, the distribution of its similarity with any other node moves to the (far) right, hence validating negative links to a much larger extent; the local filter is, thus, stricter with positive than with negative links - and their number is, somehow, re-balanced towards the naïve value.
Figure 6: Distribution of the Jaccard similarity between pairs of projections. The numerator of such an index represents the number of positive (negative) negative links occupying the same position in the adjacency matrix of both subjects and its denominator represents the cardinality of the union of the two sets of positive (negative) links populating the adjacency matrices to be compared. Kernel density estimation has been employed for better visualisation; bin widths are computed according to the Freedman-Diaconis rule. Since a value of 111 indicates perfect overlap while a value of 0 indicates no overlap, positive links yield a larger Jaccard similarity than negative links, a result indicating the former ones as more spatially coherent than the latter ones.
VIII.2.2 Tendency to make hubs
A related index is the one named tendency to make hubs (TMH), introduced in Saberi et al. (2021) and quantifying the tendency of a given degree distribution - when dealing with signed graphs, both the positive and the negative one - to host hubs. Hereby, we adopt a slightly different index, rooted in statistical theory and reducing the functional form of a given degree distribution to just one number. Named coefficient of variation (CV), it is defined as the ratio between the standard deviation and the expected value of the reference distribution: in symbols,
$\text{CV}^{+}=\frac{\text{std}[k^{+}]}{\overline{k^{+}}}=\frac{\text{std}[k^{+}]}{2L^{+}/N}$ (80)
and
$\text{CV}^{-}=\frac{\text{std}[k^{-}]}{\overline{k^{-}}}=\frac{\text{std}[k^{-}]}{2L^{-}/N};$ (81)
[…]
As evident from fig. 5, none of our projections seems to be characterised by the presence of positive hubs; when considering the negative degrees, instead, all projections display a right-skewed distribution that suggests the presence of nodes whose number of connections is (much) larger than that of the others: in a sense, the subgraph induced by the positive links tends to be more homogeneous than the subgraph induced by the negative ones [^17] As highlighted in Luppi et al. (2024), ‘[…] whereas cooperative (positive-valued) interactions in the effective connectivity tend to be strong, modular and relatively short-range, competitive interactions are weaker but more long-range, more diffuse and less clustered […]’..
[…]
A related result is the one concerning the degree distributions of cortical and subcortical areas: as evident from the figures in Appendix C, subcortical areas tend to have a larger negative degree than cortical areas.
VIII.2.3 Spatial coherence of signed connections
From what we have learned so far, the positive subgraphs are (overall) structurally homogeneous - positive degree distributions are bell-shaped and well-represented by the average - while the negative ones are (overall) structurally heterogeneous - negative degree distributions are fat-tailed, hence over-dispersed around the average. Let us, now, ask ourselves how similar positive and negative subgraphs are across subjects.
Figure 7: Distributions of the Strong and Weak Degree of Frustration. Box-plots summing up the distributions of the Strong Degree of Frustration (SDoF) and the Weak Degree of Frustration (WDoF) across the three, different projections of our 100100100 subjects. Each box-plot illustrates the distribution of the corresponding metric, indicating its central tendency and dispersion, for each model: in both cases, the bSRGM-induced projections are characterised by the smallest median and largest concentration while the naïve and bSCM-induced projections are peaked in correspondence of largest values but show more variability. Since $T^{(---)}\simeq 0$, the SDoF and the WDoF are practically identical and both positive, a result confirming the non null level of frustration characterising our brain networks.
In order to answer this question, we calculated the Jaccard similarity between pairs of (adjacency matrices representing) individual projections, per benchmark. As fig. 6 shows, the positive variant of such an index (i.e. the percentage of positive links located in the same position) ranges between 0.60.60.6 and 111 while its negative variant (i.e. the percentage of negative links located in the same position) ranges between 0 and 0.40.40.4: in words, the spatial coherence of positive links is larger than the spatial coherence of negative links, which appear as those (better) characterising single subjects [^19] As highlighted in Luppi et al. (2024), ‘[…] Additionally, we find that the model with competitive interactions is not only better able to reflect the empirical FC that it is intended to model but is also better at distinguishing between the FC patterns of different individuals […]’ and ‘[…] Thus, we conclude that the model with competitive interactions is significantly more individual-specific […]’.
Figure 8: Heatmaps of signed connections within and between brain regions. Values are averaged across subjects, for each of our benchmarks. The intensity of each colour is proportional to the magnitude of the corresponding entry. Network labels stand for control (CONT), dorsal attention (DA), limbic (LIM), somatomotor (SM), salience/ventral attention (SVA), visual (VIS) networks and default mode network (DMN). While the largest number of positive links is found within the DMN, the largest number of negative links is found between the DMN and the DA network, although a large number of the same kind of connections is found between the DMN and the SM, SVA and VIS networks as well - consistently across all benchmarks.
VIII.2.4 Assessing the brain (im)balance at the microscale
Let us, now, inspect the way negative links are arranged within projections: to this aim, one may consider the index named Strong Degree of Balance (SDoB) Gallo et al. (2024a) and quantifying the percentage of triads characterised by an even number of negative links, i.e. either zero or two - in other words, the ones that are balanced according to the Strong Balance Theory (SBT). Here, instead, we consider its complementary to 111 and name it [^20] A natural choice would have been that of naming it Strong Degree of Frustration (SDoF); since, however, the triad ()(---)( - - - ) is heavily under-expressed (see also fig. 7), the strong and weak versions practically coincide: we have, thus, avoided to explicitly distinguish between the two. Such an evidence has been also highlighted in Luppi et al. (2024), where it is reported that ‘[…] it is less likely that competitively-interacting neighbours of a node will themselves be interacting competitively […]’ Index of Brain Imbalance (IBI): more quantitatively,
$\text{IBI}=\frac{T^{(+-+)}+T^{(---)}}{T^{(+++)}+T^{(-+-)}+T^{(+-+)}+T^{(---)}}$ (82)
with
$T^{(+++)}$ $=\frac{\text{Tr}[(\mathbf{A}^{+})^{3}]}{6},$ (83)
$T^{(-+-)}$ $=\frac{\text{Tr}[\mathbf{A}^{-}\mathbf{A}^{+}\mathbf{A}^{-}]}{2},$ (84)
$T^{(+-+)}$ $=\frac{\text{Tr}[\mathbf{A}^{+}\mathbf{A}^{-}\mathbf{A}^{+}]}{2},$ (85)
$T^{(---)}$ $=\frac{\text{Tr}[(\mathbf{A}^{-})^{3}]}{6};$ (86)
naturally, the closer the IBI value to 111, the larger the amount of unbalanced triads.
[…]
Let us, now, consider the brain parcellation carried out in Thomas Yeo et al. (2011), where the authors propose a functional segmentation of brain areas by distinguishing regions (hereby, Yeo’s regions) involved in sensory tasks (e.g. the visual and the somatomotor networks) from those involved in higher-order, cognitive tasks (e.g. the DMN): they are named control (CONT), dorsal attention (DA), limbic (LIM), somatomotor (SM), salience/ventral attention (SVA), visual (VIS) networks and default mode network (DMN). While the analysis carried out in Demertzi et al. (2022) highlights the presence of negative correlations between the DMN and the so-called task-positive regions, the one carried out in Saberi et al. (2022) shows that, in healthy individuals, unbalanced triads predominantly occur between different Yeo’s regions.
Figure 9: Tensor of the IBI values associated with the triplets of Yeo’s regions. Values are averaged across subjects, for each of our benchmarks. The size of each point is proportional to the standard deviation of the corresponding value. Network labels stand for control (CONT), dorsal attention (DA), limbic (LIM), somatomotor (SM), salience/ventral attention (SVA), visual (VIS) networks and default mode network (DMN). While the largest IBI values characterise the triplets where the DMN is present once or, at most, twice, the smallest values characterise those belonging to the same Yeo’s region (especially the SM, SVA and VIS networks) - consistently across all benchmarks.
In order to verify the first of the aforementioned findings, we have constructed two $\times 7$ 7 matrices: each entry of the first (second) one encodes the number of positive (negative) links between the corresponding Yeo’s regions, aggregated across individuals. As fig. 8 shows, i) the largest number of positive links is found within the DMN; ii) the largest number of negative links is found between the DMN and the DA network, although a large number of the same kind of connections is found between the DMN and the SM, SVA and VIS networks as well - consistently across all benchmarks.
In order to verify the second of the aforementioned findings, we have constructed a $\times 7\times 7$ 7 tensor, each entry of which encodes the IBI associated with a given triplet of Yeo’s regions, averaged across individuals (see fig. 9). Still, as the triplets with a smaller IBI exhibit a larger inter-subject variability, we have also considered the distribution of such values: according to the MW test, i) the triads belonging to the same Yeo’s region are those exhibiting the lowest level of imbalance, the smallest values characterising the triplets involving nodes from the SM, SVA and VIS networks; ii) the triads where the DMN is present once or, at most, twice, exhibit the highest level of imbalance - consistently across all benchmarks.
VIII.3 Patterns of brain (im)balance at the mesoscale
Following Gallo et al. (2024b), one can adopt an ‘agnostic’ attitude and explore the mesoscale organisation of a projection without aligning with any specific, conceptual framework: a principled approach to achieve such a goal is that of minimising the Bayesian Information Criterion (BIC), i.e.
BIC $=\kappa\ln V-2\ln\mathcal{L};$ (87)
the first addendum proxies the complexity of a model with the number of its parameters, $\kappa$, the second addendum proxies the accuracy of a model with its log-likelihood, $\ln\mathcal{L}$, and $V=N(N-1)/2$ accounts for the system dimensions.
Figure 10: Community detection for subject $\bm{106016}$. Top panels: heatmaps illustrating the communities partitioning the projections of subject 106016106016106016, detected by minimising the Bayesian Information Criterion (BIC). Although the number of detected communities varies from 555 to 888, the overall message conveyed by BIC minimisation does not change: the brain seems to be constituted by groups of areas working either ‘in phase’ or ‘in counterphase’ most of the time, each one being constituted by areas working ‘in phase’ most of the time. Bottom panels: heatmaps illustrating the communities partitioning the projections of subject 106016106016106016, detected by minimising the frustration ($F$). The number of detected communities varies from 222 to 444 but speaking of modules is, now, a bit imprecise as the $F$-induced partition is dictated by the sole request of minimising the internal number of positive links and the external number of negative links.
Since we aim at describing a projection mesoscale organisation, a natural choice is that of instantiating BIC with the Signed Stochastic Block Model (SSBM), defined by the likelihood function
$\mathcal{L}_{\text{SSBM}}=$ $\prod_{r=1}^{k}(p_{rr}^{+})^{L_{rr}^{+}}(p_{rr}^{-})^{L_{rr}^{-}}(1-p_{rr}^{+}-p_{rr}^{-})^{\binom{N_{r}}{2}-L_{rr}}$
$\prod_{r=1}^{k}\prod_{\begin{subarray}{c}s=1\\ s>r\end{subarray}}^{k}(p_{rs}^{+})^{L^{+}_{rs}}(p_{rs}^{-})^{L^{-}_{rs}}(1-p_{rs}^{+}-p_{rs}^{-})^{N_{r}N_{s}-L_{rs}}$ (88)
and a number of parameters $\kappa_{\text{SSBM}}=k(k+1)$, amounting at twice the number of modules, $k$, into which the projection can be partitioned. Naturally, $N_{r}$ is the number of nodes constituting block $r$, $p_{rr}^{+}$ is induced by the empirical number of positive links within block $r$, i.e.
$p_{rr}^{+}=\frac{2L_{rr}^{+}}{N_{r}(N_{r}-1)},$ (89)
$p_{rr}^{-}$ is induced by the empirical number of negative links within block $r$, i.e.
$p_{rr}^{-}=\frac{2L_{rr}^{-}}{N_{r}(N_{r}-1)},$ (90)
$p_{rs}^{+}$ is induced by the empirical number of positive links between blocks $r$ and $s$, i.e.
$p_{rs}^{+}=\frac{L_{rs}^{+}}{N_{r}N_{s}},$ (91)
and $p_{rs}^{-}$ is induced by the empirical number of negative links between blocks $r$ and $s$, i.e.
$p_{rs}^{-}=\frac{L_{rs}^{-}}{N_{r}N_{s}}.$ (92)
Figure 11: Community detection for subject $\bm{857263}$. Top panels: heatmaps illustrating the communities partitioning the projections of subject 857263857263857263, detected by minimising the Bayesian Information Criterion (BIC). Although the number of detected communities varies from 666 to 888, the overall message conveyed by BIC minimisation does not change: the brain seems to be constituted by groups of areas working either ‘in phase’ or ‘in counterphase’ most of the time, each one being constituted by areas working ‘in phase’ most of the time. Bottom panels: heatmaps illustrating the communities partitioning the projections of subject 857263857263857263, detected by minimising the frustration ($F$). The number of detected communities is at most 222 but speaking of modules is, now, a bit imprecise as the $F$-induced partition is dictated by the sole request of minimising the internal number of positive links and the external number of negative links.
VIII.3.1 A survey on individual projections
Let us start by considering a pair of individual partitions, stressing that a large link density does not prevent BIC minimisation from detecting statistically significant mesoscopic structures: its sensitivity to both the number and the sign of links, in fact, makes it capable of partitioning both denser and sparser configurations.
Naïve projections. As figs. 10 and 11 show, BIC minimisation reveals the presence of less than 101010 (more precisely, from 555 to 666) communities: the vast majority of links within communities - if not all, as evident upon looking at some of the modules - is positive, a result implying that the series describing the activity of these areas are concordant for most of the time. For what concerns the links between communities, instead, the picture becomes much more interesting: the third and fourth community in fig. 10 are, in fact, connected by mostly negative links, an evidence suggesting that the two blocks of areas are ‘in counterphase’ for most of the time; while this holds true for the first and third community as well, the first and fourth community are connected by mostly positive links, i.e. even if these blocks of areas are separated, they still seem to work ‘in phase’.
Validated projections. Filtering our projections with a homogeneous benchmark makes the picture even cleaner, as each block is, now, constituted by links of practically the same sign: more precisely, while positive links constitute the majority of links, negative links are found only between modules. Filtering our projections with a heterogeneous benchmark, instead, returns a picture which is halfway between the naïve one and the bSRGM-induced one, as the blocks are sparser than the naïve ones but less homogeneous than the ones induced by the bSRGM. The overall message conveyed by such an analysis, however, does not change: the brain seems to be constituted by groups of areas working either ‘in phase’ or ‘in counterphase’ most of the time, each one being constituted by areas working ‘in phase’ most of the time.
Figure 12: Heatmaps of the communities partitioning the bSCM-induced projections of four subjects. The communities partitioning the bSCM-induced projections of subjects 106016106016106016, 857263857263857263, 115320115320115320 and 654754654754654754 are detected by minimising the Bayesian Information Criterion (BIC). Matrices are coloured according to eq. 96 and eq. 97, i.e. blocks are blue (red) if the majority of links populating them is positive (negative); white blocks on the diagonal indicate singletons - specifically, a cortical region for subject 106016106016106016 and a subcortical region for subject 654754654754654754. All subjects satisfy the statistical variant of the Relaxed Balance Theory, being characterised by blocks with a majority of positive links both on and off the main diagonal - subjects 115320115320115320 and 654754654754654754 are even characterised by one, diagonal block (i.e. a community) populated by a majority of negative links.
VIII.3.2 Traditionally or relaxedly balanced?
Motivated by the last observation, we now employ our benchmarks to probe the patterns of structural (im)balance at the mesoscopic level. We also explicitly notice that speaking of balance theory for brain networks may appear a bit improper: still, this is the only existing framework to interpret our results.
Traditional theory of balance. According to the strong variant of the traditional theory of balance (TBT), the optimal partition of the set of nodes of a network consists of two groups, whose internal (external) connections are dominated by positive (negative) signs; by contrast, the weak variant of the TBT allows for any number of groups. More formally, both versions of the TBT can be probed by minimising the frustration, defined as
$F(\bm{\sigma})=L_{\bullet}^{-}+L_{\circ}^{+},$ (93)
i.e. as the number of negative links within modules (indicated with a filled dot) plus the number of positive links between modules (indicated with an empty dot).
Figure 13: Basic statistics of the partitions across subjects. Top panels: histograms of the number of communities detected on the projections returned by each benchmark and optimisation criterion. While $F$ avoids creating frustrated patterns, by individuating a smaller number of larger communities, BIC singles out a larger number of smaller communities. Middle panels: histograms of the size of communities detected on the projections returned by each benchmark and optimisation criterion. These panels refine the picture provided by the previous ones, clarifying that $F$ tends to separate few singletons from the set of remaining nodes. BIC tends to individuate singletons as well, although the size of communities into which the set of remaining nodes is partitioned varies over a broader range (approximately, from 111 to 404040). Bottom panels: number of times each brain region is individuated as a singleton. As evident from the plot, this is mainly observed for subcortical regions, acting as ‘integrators’ of cortical ones. More quantitatively, the fraction of singletons that are identifiable as subcortical regions amounts at i) 0.9$\simeq$ 0.9≃ 0.9 according to $F$ and 0.9$\simeq$ 0.9≃ 0.9 according to BIC, on the naïve projections; ii) 0.9$\simeq$ 0.9≃ 0.9 according to $F$ and 0.7$\simeq$ 0.7≃ 0.7 according to BIC, on the bSRGM-induced projections; iii) 0.8$\simeq$ 0.8≃ 0.8 according to $F$ and 0.5$\simeq$ 0.5≃ 0.5 according to BIC, on the bSCM-induced projections.
As figs. 10 and 11 show, minimising $F$ always leads to a smaller number of communities with respect to the one revealed by minimising BIC: in fact, from 222 to 444, depending on the projection. The naïve and bSCM-induced ones are the most similar, as 222 mostly positive blocks (one being much larger than the others), separated by a vast majority of negative links, are individuated. The naïve, $F$-induced partition also illustrates the way the optimisation of $F$ works: the large number of positive links leads $F$ to collect them together, being more convenient to accommodate a small number of negative links within a larger, mostly positive, block than letting a large number of positive links be placed between blocks. In any case, the minimisation of $F$ further confirms that none of our projections obeys the traditional formulation of the balance theory.
Statistical theory of balance. Such a result motivates us to inspect the theory of balance best accommodating our observations. To answer such a question, let us, first, notice that the TBT can be formalised upon writing
$p_{rr}^{-}=0,\quad\forall\>r$ (94)
and
$p_{rs}^{+}=0,\quad\forall\>r<s$ (95)
with $p_{rr}^{+}$ indicating the probability that any two nodes belonging to the same block $r$ are connected by a positive link, $p_{rs}^{+}$ indicating the probability that any two nodes belonging to the different blocks $r$ and $s$ are connected by a positive link and analogously for their negative counterparts. As noticed in Gallo et al. (2024b), however, the positions defining the TBT can be replaced with the milder ones reading
$\text{sgn}[p_{rr}^{+}-p_{rr}^{-}]=+1,\quad\forall\>r$ (96)
which amounts at requiring $p_{rr}^{+}>p_{rr}^{-}$, $\forall\>r$ and
$\text{sgn}[p_{rs}^{+}-p_{rs}^{-}]=-1,\quad\forall\>r<s$ (97)
which amounts at requiring $p_{rs}^{+}<p_{rs}^{-}$, $\forall\>r<s$: in words, the deterministic rules firstly defined by Cartwright, Harary and Davis are replaced by a set of probabilistic criteria individuating ‘a tendency’ to obey, or not to obey, the TBT. More formally, a configuration satisfying eqs. 96 and 97 will be claimed to support the statistical variant of the TBT - specifically, its strong variant if $k=2$ and its weak variant if $k>2$; if, instead, $p_{rr}^{+}\leq p_{rr}^{-}$ for some diagonal blocks or $p_{rs}^{+}\geq p_{rs}^{-}$ for some off-diagonal blocks, it will be claimed to support the statistical variant of the relaxed balance theory (RBT).
Let us, again, take a look at figs. 10 and 11: although $p_{rs}^{+}\geq p_{rs}^{-}$ for all diagonal blocks (i.e. with $r=s$), $p_{rs}^{+}\geq p_{rs}^{-}$ for several, off-diagonal blocks (i.e. with $r\neq s$) as well: we are, thus, led to the conclusion that all projections obey the statistical variant of the RBT Gallo et al. (2024b) as a non-negligible number of positive links is found between - and not only within - clusters.
To evaluate the alignment of our projections with the statistical variant of the TBT, let us count the number of subjects obeying eqs. 96 and 97. When considering the BIC-induced partitions, we find that no subject obeys the statistical variant of the TBT, irrespectively from the benchmark employed: while a vast majority of diagonal blocks is, in fact, populated by a number of positive links that is larger than the number of negative ones - although subjects admitting diagonal blocks populated by a majority of negative links exist (see also fig. 12) - also several, off-diagonal blocks are populated by a number of positive links that is larger than the number of negative ones - as the two examples in figs. 10 and 11 clearly show, $p_{rs}^{+}\geq p_{rs}^{-}$ for many of them.
When considering the $F$-induced partitions, instead, 65 naïve projections, 47 bSRGM-induced projections and 69 bSCM-induced projections obey the statistical variant of the TBT.
Figure 14: Distributions of BIC and $F$ values across the partitions induced by their minimisation. Top panels: histograms of the BIC values characterising the partitions determined by minimising it and Yeo’s partitions. As expected, the former are characterised by a slightly smaller value of such a score than Yeo’s partitions. Middle panels: histograms of the $F$ values characterising the partitions determined by minimising it and Yeo’s partitions. In this second case, the former are characterised by a much smaller value of such a score than Yeo’s partitions. Bottom panels: histograms of the $F$ values characterising the partitions determined by minimising BIC and Yeo’s partitions. In this third case, both are characterised by a quite large value of $F$. Taken together, our results indicate that brain networks do not seek to minimise frustration although BIC minimisation seems (to embody a principle that is) capable of capturing a structure quite similar to the one individuated by Yeo’s task-positive regions. Kernel density estimation has been employed for better visualisation; bin widths are computed according to the Freedman-Diaconis rule.
VIII.4 Towards a representative brain?
Let us, now, move from individual analyses to the aggregated one, the aim being that of understanding the extent to which the previous results can be generalised.
VIII.4.1 Distribution of the number of communities
First, let us plot the distribution of the number of communities per benchmark and optimisation criterion. The results are shown in fig. 13: the number of communities detected by minimising the frustration is practically always smaller than the number of communities detected by minimising BIC - the $F$-induced mode being 222 and the BIC-induced mode varying from 777 to 888; moreover, the $F$-induced distributions are narrower than the BIC-induced ones, a result hinting at a more pronounced inter-subject variability, in the second case.
VIII.4.2 The role of singletons
Figure 13 illustrates the histograms of the number of communities detected on the projections returned by each benchmark and optimisation criterion, revealing that $F$ individuates a smaller number of larger communities while BIC singles out a larger number of smaller communities. The histograms of the size of communities detected on the same projections refine the aforementioned picture, clarifying that $F$ tends to individuate (overall, too) simple structures such as few singletons detached from the rest of the nodes; although BIC individuates singletons as well, the size of communities into which the rest of the nodes is partitioned ranges from 111 to 404040.
Interestingly, singletons often coincide with the so-called subcortical regions, i.e. areas serving as hubs ‘integrating’ the cortical regions, among the other things: more quantitatively, the fraction of subcortical singletons amounts at i) 0.9$\simeq$ 0.9≃ 0.9 according to $F$ and 0.9$\simeq$ 0.9≃ 0.9 according to BIC, on the naïve projections; ii) 0.9$\simeq$ 0.9≃ 0.9 according to $F$ and 0.7$\simeq$ 0.7≃ 0.7 according to BIC, on the bSRGM-induced projections; iii) 0.8$\simeq$ 0.8≃ 0.8 according to $F$ and 0.5$\simeq$ 0.5≃ 0.5 according to BIC, on the bSCM-induced projections.
VIII.4.3 Distribution of BIC and $F$ values
A related analysis is the one concerning the distributions of the BIC and $F$ values across the partitions determined by minimising each of them. As fig. 14 confirms, the partitions retrieved by minimising BIC are characterised by a smaller value of such a score than Yeo’s partitions: while this is expected, it should be also noticed that these differences are not very large. Besides, both kinds of partitions are characterised by a similar, and quite large, value of $F$. Inspecting the distribution of the values of $F$, instead, reveals that the partitions retrieved by minimising it are much less frustrated than Yeo’s partitions. Taken together, our results suggest that while brain networks do not seek to minimise frustration, BIC minimisation represents a recipe that is capable of capturing a structure de facto very close to the one captured by Yeo’s task-positive regions.
VIII.4.4 The barycentrical partition
Let us, now, ask ourselves if a criterion to condensate the information embodied by each, individual partition can be devised; naturally, the aforementioned results let us opt for solely focusing on the BIC-induced ones.
A simple recipe could be that of averaging the individual projections in a benchmark-specific fashion; still, the presence of a vast majority of positive links and the spatial incoherence of the negative ones would cause most of the latter to be deleted.
Alternatively, one could try to individuate the ‘most representative’ brain by individuating the ‘most representative’ mesoscale structure: finding an answer to this question ultimately amounts at finding a method to compare any two partitions, say $X\equiv\{x_{1}\dots x_{n}\}$ and $Y\equiv\{y_{1}\dots y_{m}\}$.
Rand index (RI). One of the simplest metric is the RI, reading
$\text{RI}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{FP}+\text{TN}+\text{FN}}=\frac{\text{TP}+\text{TN}}{N(N-1)/2}$ (98)
[…]
Figure 15: Brain centroids individuated by our measures of distance between partitions. Top panels: brain centroids individuated by adopting the Rand Index (RI), the Normalised Mutual Information (NMI), the Rajski distance (RD), the Wallace Index (WI), the Jaccard Index (JI) and the Variation of Information (VI). Generally speaking, while RI, NMI and RD individuate projections with a number of communities lying on the far-right side of the spectrum (i.e. from 999 to 111111), WI, JI and VI individuate projections with a number of communities lying on the far-left side of the spectrum (i.e. 444). Such a result may be explained by noticing that indices like WI and JI are solely defined in terms of true positives: as a consequence, identifying a smaller number of larger communities may enhance the detection of such events. Indices like RI, on the other hand, account for true negatives as well, whence the larger number of smaller communities. Bottom panels: different metrics induces a different overlap between our structure-based partition and Yeo’sfunction-based one: while the larger communities individuated by WI, JI and VI tend to merge large portions of different Yeo’s areas, the smaller ones individuated by RI, NMI and RD tend to be populated by (even a large portion of) single Yeo’s areas.
Normalised mutual information (NMI). A second metric is the NMI, reading
$\text{NMI}(X,Y)$ $=\frac{2\text{MI}(X,Y)}{S(X)+S(Y)};$ (99)
here,
$\text{MI}(X,Y)=\sum_{i=1}^{N}\sum_{j=1}^{N}h_{ij}\ln\left(\frac{h_{ij}}{f_{i}g_{j}}\right),$ (100)
[…]
Rajski distance (RD). A third metric is the variation of information (VI), reading
$\text{VI}(X,Y)$ $=S(X,Y)-\text{MI}(X,Y)$
$=2S(X,Y)-S(X)-S(Y)$
$=S(X)+S(Y)-2\text{MI}(X,Y)$ (101)
or, more explicitly,
$\text{VI}(X,Y)=-\sum_{i=1}^{N}\sum_{j=1}^{N}h_{ij}\left[\ln\left(\frac{h_{ij}}{f_{i}}\right)+\ln\left(\frac{h_{ij}}{g_{j}}\right)\right];$ (102)
such an expression can be normalised by dividing it by the joint entropy $S(X,Y)=-\sum_{i=1}^{N}\sum_{j=1}^{N}h_{ij}\ln h_{ij}$, i.e.
$\text{RD}(X,Y)=\frac{\text{VI}(X,Y)}{S(X,Y)}=1-\frac{\text{MI}(X,Y)}{S(X,Y)};$ (103)
[…]
Finding the barycentre. Any of the metrics above can be employed to assess how different the partitions characterising the brains of any two subjects $u$ and $u^{\prime}$ are. Once such a matrix of distances, say $\mathbf{D}\equiv\{D_{uu^{\prime}}\}_{u,u^{\prime}=1}^{U}$, with $U=100$, has been obtained, a second method is needed to condensate such a wealth of information. Our proposal is that of calculating the vector $\underline{\mathbf{D}}\equiv\{\underline{D}_{u}\}$, where
$\underline{D}_{u}=\frac{\sum_{u^{\prime}=1}^{U}D_{uu^{\prime}}}{U},$ (104)
and take the partition lying at distance
$d=\min\left\{\underline{D}_{u}\right\}_{u=1}^{U}$ (105)
from the others, i.e. at the minimum, average one. Such a partition will be refereed to as to the barycentrical partition or brain centroid. Let us explicitly notice that while VI and RD are proper distances over the space of possible partitions of a system, hence inducing the positions $D_{uu^{\prime}}^{\text{VI}}=\text{VI}_{uu^{\prime}}$ and $D_{uu^{\prime}}^{\text{RD}}=\text{RD}_{uu^{\prime}}$, RI and NMI quantify the similarity of any two partitions [^21] In fact, both RI and NMI could be directly maximised., hence inducing the positions $D_{uu^{\prime}}^{\text{RI}}=1-\text{RI}_{uu^{\prime}}$ and $D_{uu^{\prime}}^{\text{NMI}}=1-\text{NMI}_{uu^{\prime}}$.
The observations about the importance of considering filtered projections and the prominent role played by heterogeneous benchmarks motivate us to focus on the partitions induced by the bSCM.
Upon looking at fig. 15, we realise that i) RI individuates a barycentrical projection with 111111 communities; ii) NMI and RD individuate the same projection, in turn very close to the one individuated by RI - the only difference being the number of communities, now reading 999; iii) WI, JI and VI individuate a barycentrical projection with 444, major communities (plus a number of singletons). To sum up, while RI, NMI and RD individuate projections with a larger number of communities, WI, JI and VI individuate projections with a smaller number of communities. The explanation of such a result may lie in the evidence that indices like WI and JI are solely defined in terms of true positives: as a consequence, identifying a smaller number of larger communities may enhance the detection of such events - although negative connections become under-represented; indices like RI, on the other hand, account for true negatives as well, whence the larger number of smaller communities - connected by a vast majority of negative links.
As a last comment, let us notice that RI, NMI and RD retrieve partitions with a larger BIC than those retrieved by WI, JI and VI.
Let us, now, compare the community structure of the different brain centroids with the partition defined by the Yeo’s regions introduced before. As fig. 15 shows, each metrics induces a different overlap between our structure-based partition and Yeo’s function-based one: still, while the larger communities individuated by WI, JI and VI tend to merge (even) large portions of different Yeo’s areas, some of the smaller ones individuated by RI, NMI and RD tend to be populated by (even a large portion of) single Yeo’s areas. Moreover, WI, JI and VI tend to return singletons, represented by practically empty rows.
IX Discussion
The study of the role played by negative connections within and between brain areas is a topic that has recently gained attention within the neuroscientific literature. The method that we have proposed here produces connections connotated by a negative sign as the result of a validation procedure that counts the number of times the corresponding series are found to be discordant [^22] Aggregating positive and negative contributions can, in fact, obscure the presence of anti-correlations. For instance, while the DMN and the DA network are connected by, approximately, $3\cdot 10^{5}$ positive links, they are also connected by, roughly, $9\cdot 10^{4}$ negative links: summing these numbers together would yield a net positive count, misleadingly hiding the presence of anti-correlated activity. and compare it with the prediction of a properly defined benchmark. Such a validation scheme allows us to answer a number of questions of neuroscientific interest.
The first one concerns the frustration of the brain, at both the microscopic and the mesoscopic level. For what concerns the first one, the evidence that the median IBI value is positive - irrespectively from the benchmark employed to project our set of multivariate time series - lets us conclude that brain networks are indeed populated by a vast majority of (++)(+-+)( + - + ) triples - the ones reading ()(---)( - - - ) being almost absent. Interestingly, although the negative subgraph is more prone to inter-subject variability and less homogeneous than the positive one, the negative connections constituting the frustrated triples are mostly found between the groups of nodes constituting the DMN and the groups of nodes constituting task-positive regions such as the DA, the SM, the SVM and the VIS. As already mentioned, these results match the ones in Fox et al. (2005) as well as those discussed in Demertzi et al. (2022); Greicius et al. (2003); Hampson et al. (2010).
Some contributions observe that the empirical level of (im)balance is smaller than the one characterising randomised configurations [^23] It should be, however, noticed that the results obtained in Saberi et al. (2021) depend on the chosen threshold, as also explicitly recognised by the authors. and deduce that the brain tends to stabilise itself Saberi et al. (2021). Our results, however, lead to a different conclusion: although it is true that the bSCM returns (slightly) less balanced configurations than the naïve ones, brain networks do not seem to minimise frustration: as confirmed by the distribution of $F$, the values assumed by it across the partitions characterising our and Yeo’s projections are quite larger than zero.
It is, thus, of interest asking which principle is obeyed by the brain in order to self-organise its functioning. An answer seems to be provided by fig. 14: task-based partitions align with those recovered by minimising BIC; in a sense, thus, some sort of optimisation principle seems to be at work at the mesoscopic scale. Still, individuating the partition ‘best representing’ a signed brain network remains a challenging task, as different metrics provide different answers; nevertheless, a message whose validity appear as quite general seems to emerge: brain networks align with the (statistical variant) of the Relaxed Balanced Theory, as a vast majority of positive signs populates off-diagonal blocks as well.
X Conclusions
A critical, yet often overlooked, aspect of functional brain data is the role played by negative links: frequently ignored due to the interpretative challenges brought along, emerging evidence indicates that negative links play a fundamental role in shaping the (structural and dynamical) properties of resting-state networks, especially those related to self-stabilisation. As excluding negative links, thus, means underestimating the brain ability to function properly, we have proposed a validation technique allowing statistically significant signed links (be they positive or negative) to emerge out of the noise accompanying raw data. To this aim, two, different benchmarks have been proposed, filtering multivariate time series to a different extent and letting diverse patterns emerge out of the wealth of empirical observations: in particular, while the bSRGM emphasizes structural balance by levelling out the differences between time series, the bSCM preserves their heterogeneity, thus inducing sparser projections that retain more information about the original, bipartite structure. In particular, our analysis points out that brain networks are, indeed, frustrated, the patterns induced by signed connections overlapping to a quite large extent with those individuated by Yeo’s task-oriented partitions and returning the picture of a system aligning (better) with the statistical variant of the Relaxed Balance Theory.
A future research direction is that of employing our algorithm to carry out both a ‘weighted-to-binary’ and a ‘weighted-to-weighted’ analysis, i.e. obtain statistically validated binary and weighted projections of multivariate weighted time series, defined by the ‘rule’ that any two nodes sharing a number of (either concordant or discordant) relationships characterised by a significantly large weight are connected by an edge ‘dressed’ with the validated weight.
XI Acknowledgments
This work has been supported by the projects: ‘SoBigData RI PPP - SoBigData RI Preparatory Phase Project’, funded by the European Union under the scheme HORIZON-INFRA-2021-DEV-02-01, preparatory phase of new ESFRI research infrastructure projects, G.A. 101079043; ‘Reconstruction, Resilience and Recovery of Socio-Economic Networks’ RECON-NET EP_FAIR_005 - PE0000013 ‘FAIR’, funded by the European Commission under the Next Generation EU program, PNRR Mission 4 Component 2 Investment 3.1; ‘FAIR - Future Artificial Intelligence Research’ - Spoke 1 ‘Human-centered AI’, funded by the European Commission under the Next Generation EU program, PNRR Mission 4 Component 2 Investment 3.1, G.A. PE00000013; ‘XAI: Science and technology for the eXplanation of AI decision making’, ERC-2018-ADG, G.A. 834756.
XII Data availability
XIII Code availability
The Python package named SIBERIA (SIgned BEnchmarks foR tIme series Analysis), implementing our benchmarks for the analysis of multivariate time series, is available on PyPI and at the URL https://github.com/MarsMDK/SIBERIA.
XIV Author contributions
Study conception and design: MIB, MDV, TG, TS. Literature review: EA, MDV, ST. Data collection: EA. Analysis and interpretation of results: EA, MDV, TS. Draft manuscript preparation: EA, MDV, TS. Draft manuscript revision: EA, MIB, MDV, ST, TG, TS.
XV Competing Interests
The authors declare no competing interests.
Appendix A Probabilistic models for binary undirected bipartite signed networks
The generalisation of the ERG formalism for the analysis of binary undirected bipartite signed graphs rests upon the constrained maximisation of Shannon entropy, i.e.
$\mathscr{L}=S[P]-\sum_{i=0}^{M}\theta_{i}[P(\mathbf{B})C_{i}(\mathbf{B})-\langle C_{i}\rangle]$ (106)
where
$S=-\sum_{\mathbf{B}\in\mathbb{B}}P(\mathbf{B})\ln P(\mathbf{B}),$ (107)
$C_{0}\equiv\langle C_{0}\rangle\equiv 1$ sums up the normalisation condition and the remaining $M-1$ constraints represent proper, topological properties. Such an optimisation procedure defines the expression
$P(\mathbf{B})=\frac{e^{-H(\mathbf{B})}}{Z}=\frac{e^{-H(\mathbf{B})}}{\sum_{\mathbf{B}\in\mathbb{B}}e^{-H(\mathbf{B})}}=\frac{e^{-\sum_{i=1}^{M}\theta_{i}C_{i}(\mathbf{B})}}{\sum_{\mathbf{B}\in\mathbb{B}}e^{-\sum_{i=1}^{M}\theta_{i}C_{i}(\mathbf{B})}}$ (108)
that can be made explicit only after having chosen a specific set of constraints.
1. Signed Random Graph Model
Let us consider the properties $L^{+}(\mathbf{B})$ and $L^{-}(\mathbf{B})$. The Hamiltonian describing such a problem reads
$H(\mathbf{B})=\alpha\>B^{+}(\mathbf{B})+\gamma\>B^{-}(\mathbf{B})=\sum_{i=1}^{N}\sum_{t=1}^{T}(\alpha\>b_{it}^{+}+\gamma\>b_{it}^{-})$ (109)
and induces a partition function reading
$Z$ $=\sum_{\mathbf{B}\in\mathbb{B}}e^{-H(\mathbf{B})}$
$=\sum_{\mathbf{B}\in\mathbb{B}}e^{-\alpha\>B^{+}(\mathbf{B})-\gamma\>B^{-}(\mathbf{B})}$
$=\sum_{\mathbf{B}\in\mathbb{B}}e^{-\sum_{i=1}^{N}\sum_{t=1}^{T}(\alpha\>b_{it}^{+}+\gamma\>b_{it}^{-})}$
$=\sum_{\mathbf{B}\in\mathbb{B}}\prod_{i=1}^{N}\prod_{t=1}^{T}e^{-(\alpha\>b_{it}^{+}+\gamma\>b_{it}^{-})}$
$=\prod_{i=1}^{N}\prod_{t=1}^{T}\left[\sum_{b_{it}=-1,1}e^{-(\alpha\>b_{it}^{+}+\gamma\>b_{it}^{-})}\right]$
$=\prod_{i=1}^{N}\prod_{t=1}^{T}(e^{-\alpha}+e^{-\gamma})$
$=(e^{-\alpha}+e^{-\gamma})^{N\cdot T}.$ (110)
The expression above leads us to find
$P_{\text{bSRGM}}(\mathbf{B})=\frac{e^{-\alpha\>B^{+}(\mathbf{B})-\gamma\>B^{-}(\mathbf{B})}}{(e^{-\alpha}+e^{-\gamma})^{N\cdot T}}\equiv\frac{x^{B^{+}(\mathbf{B})}y^{B^{-}(\mathbf{B})}}{(x+y)^{N\cdot T}}\equiv(p^{-})^{B^{-}(\mathbf{B})}(p^{+})^{B^{+}(\mathbf{B})}$ (111)
having posed $p^{-}\equiv\frac{e^{-\gamma}}{e^{-\alpha}+e^{-\gamma}}\equiv\frac{y}{x+y}$ and $p^{+}\equiv\frac{e^{-\alpha}}{e^{-\beta}+e^{-\gamma}}\equiv\frac{x}{x+y}$, where $p^{+}$ is the probability that a generic series assumes a positive value at time $t$ and $p^{-}$ is the probability that a generic series assumes a negative value at time $t$.
In order to determine the parameters that define the bSRGM, let us maximise the likelihood function
$\mathcal{L}_{\text{bSRGM}}(x,y)\equiv\ln P_{\text{bSRGM}}(\mathbf{B}^{*}|x,y)=B^{+}(\mathbf{B}^{*})\ln x+B^{-}(\mathbf{B}^{*})\ln y-(N\cdot T)\ln(x+y)$ (112)
with respect to $x$ and $y$. Upon doing so, we obtain the pair of equations
$\frac{\partial\mathcal{L}_{\text{bSRGM}}(x,y)}{\partial x}$ $=\frac{B^{+}(\mathbf{B}^{*})}{x}-\frac{N\cdot T}{x+y},$ (113)
$\frac{\partial\mathcal{L}_{\text{bSRGM}}(x,y)}{\partial y}$ $=\frac{B^{-}(\mathbf{B}^{*})}{y}-\frac{N\cdot T}{x+y};$ (114)
equating them to zero leads us to find
$p^{+}$ $=B^{+}(\mathbf{B}^{*})/(N\cdot T),$ (115)
$p^{-}$ $=B^{-}(\mathbf{B}^{*})/(N\cdot T).$ (116)
2. Signed Configuration Model
Let us consider the properties $\{k_{i}^{+}(\mathbf{B})\}_{i=1}^{N}$, $\{k_{i}^{-}(\mathbf{B})\}_{i=1}^{N}$, $\{\kappa_{t}^{+}(\mathbf{B})\}_{t=1}^{T}$ and $\{\kappa_{t}^{-}(\mathbf{B})\}_{t=1}^{T}$. The Hamiltonian describing such a problem reads
$H(\mathbf{B})=\sum_{i=1}^{N}[\alpha_{i}k_{i}^{+}(\mathbf{B})+\gamma_{i}k_{i}^{-}(\mathbf{B})]+\sum_{t=1}^{T}[\delta_{t}\kappa_{t}^{+}(\mathbf{B})+\eta_{t}\kappa_{t}^{-}(\mathbf{B})]$ (117)
and induces a partition function reading
$Z$ $=\sum_{\mathbf{B}\in\mathbb{B}}e^{-H(\mathbf{B})}$
$=\sum_{\mathbf{B}\in\mathbb{B}}e^{-\sum_{i=1}^{N}[\alpha_{i}k_{i}^{+}(\mathbf{B})+\gamma_{i}k_{i}^{-}(\mathbf{B})]-\sum_{t=1}^{T}[\delta_{t}\kappa_{t}^{+}(\mathbf{B})+\eta_{t}\kappa_{t}^{-}(\mathbf{B})]}$
$=\sum_{\mathbf{B}\in\mathbb{B}}e^{-\sum_{i=1}^{N}\sum_{t=1}^{T}[(\alpha_{i}+\delta_{t})b_{it}^{+}+(\gamma_{i}+\eta_{t})b_{it}^{-}]}$
$=\sum_{\mathbf{B}\in\mathbb{B}}\prod_{i=1}^{N}\prod_{t=1}^{T}e^{-(\alpha_{i}+\delta_{t})b_{it}^{+}-(\gamma_{i}+\eta_{t})b_{it}^{-}}$
$=\prod_{i=1}^{N}\prod_{t=1}^{T}\left[\sum_{b_{it}=-1,1}e^{-(\alpha_{i}+\delta_{t})b_{it}^{+}-(\gamma_{i}+\eta_{t})b_{it}^{-}}\right]$
$=\prod_{i=1}^{N}\prod_{t=1}^{T}\left[e^{-(\alpha_{i}+\delta_{t})}+e^{-(\gamma_{i}+\eta_{t})}\right].$ (118)
The expression above leads us to find
$P_{\text{bSCM}}(\mathbf{B})$ $=\frac{e^{-\sum_{i=1}^{N}[\alpha_{i}k_{i}^{+}(\mathbf{B})+\gamma_{i}k_{i}^{-}(\mathbf{B})]-\sum_{t=1}^{T}[\delta_{t}\kappa_{t}^{+}(\mathbf{B})+\eta_{t}\kappa_{t}^{-}(\mathbf{B})]}}{\prod_{i=1}^{N}\prod_{t=1}^{T}\left[e^{-(\alpha_{i}+\delta_{t})}+e^{-(\gamma_{i}+\eta_{t})}\right]}$
$=\prod_{i=1}^{N}\prod_{t=1}^{T}\frac{x_{i}^{k_{i}^{+}(\mathbf{B})}y_{i}^{k_{i}^{-}(\mathbf{B})}z_{t}^{\kappa_{t}^{+}(\mathbf{B})}w_{t}^{\kappa_{t}^{-}(\mathbf{B})}}{x_{i}z_{t}+y_{i}w_{t}}$
$=\prod_{i=1}^{N}\prod_{t=1}^{T}(p_{it}^{-})^{b_{it}^{-}}(p_{it}^{+})^{b_{it}^{+}}$ (119)
having posed $p_{it}^{-}\equiv\frac{e^{-(\gamma_{i}+\eta_{t})}}{e^{-(\alpha_{i}+\delta_{t})}+e^{-(\gamma_{i}+\eta_{t})}}\equiv\frac{y_{i}w_{t}}{x_{i}z_{t}+y_{i}w_{t}}$ and $p_{it}^{+}\equiv\frac{e^{-(\alpha_{i}+\eta_{t})}}{e^{-(\alpha_{i}+\delta_{t})}+e^{-(\gamma_{i}+\eta_{t})}}\equiv\frac{x_{i}z_{t}}{x_{i}z_{t}+y_{i}w_{t}}$, where $p_{it}^{+}$ is the probability that series $i$ assumes a positive value at time $t$ and $p_{it}^{-}$ is the probability that series $i$ assumes a negative value at time $t$.
In order to determine the parameters that define the bSCM, let us maximise the likelihood function
$\mathcal{L}_{\text{bSCM}}(\bm{x},\bm{y},\bm{z},\bm{w})\equiv\ln P_{\text{bSCM}}(\mathbf{B}^{*}|\bm{x},\bm{y},\bm{z},\bm{w})$
$=\sum_{i=1}^{N}k_{i}^{+}(\mathbf{B}^{*})\ln x_{i}+\sum_{i=1}^{N}k_{i}^{-}(\mathbf{B}^{*})\ln y_{i}+\sum_{t=1}^{T}\kappa_{t}^{+}(\mathbf{B}^{*})\ln z_{t}+\sum_{t=1}^{T}\kappa_{t}^{-}(\mathbf{B}^{*})\ln w_{t}-\sum_{i=1}^{N}\sum_{t=1}^{T}\ln(x_{i}z_{t}+y_{i}w_{t})$ (120)
with respect to $x_{i}$, $y_{i}$, $z_{t}$ and $w_{t}$, $\forall\>i,t$. Upon doing so, we obtain the system of equations
$\frac{\partial\mathcal{L}_{\text{bSCM}}(\bm{x},\bm{y},\bm{z},\bm{w})}{\partial x_{i}}$ $=\frac{k_{i}^{+}(\mathbf{B}^{*})}{x_{i}}-\sum_{t=1}^{T}\frac{z_{t}}{x_{i}z_{t}+y_{i}w_{t}},\quad\forall\>i,$ (121)
$\frac{\partial\mathcal{L}_{\text{bSCM}}(\bm{x},\bm{y},\bm{z},\bm{w})}{\partial y_{i}}$ $=\frac{k_{i}^{-}(\mathbf{B}^{*})}{y_{i}}-\sum_{t=1}^{T}\frac{w_{t}}{x_{i}z_{t}+y_{i}w_{t}},\quad\forall\>i,$ (122)
$\frac{\partial\mathcal{L}_{\text{bSCM}}(\bm{x},\bm{y},\bm{z},\bm{w})}{\partial z_{\alpha}}$ $=\frac{\kappa_{t}^{+}(\mathbf{B}^{*})}{z_{t}}-\sum_{i=1}^{N}\frac{x_{i}}{x_{i}z_{t}+y_{i}w_{t}},\quad\forall\>t,$ (123)
$\frac{\partial\mathcal{L}_{\text{bSCM}}(\bm{x},\bm{y},\bm{z},\bm{w})}{\partial w_{\alpha}}$ $=\frac{\kappa_{t}^{-}(\mathbf{B}^{*})}{w_{t}}-\sum_{i=1}^{N}\frac{y_{i}}{x_{i}z_{t}+y_{i}w_{t}},\quad\forall\>t;$ (124)
equating them to zero leads us to find
$k_{i}^{+}(\mathbf{B}^{*})$ $=\sum_{t=1}^{T}\frac{x_{i}z_{t}}{x_{i}z_{t}+y_{i}w_{t}}=\sum_{t=1}^{T}p_{it}^{+}=\langle k_{i}^{+}\rangle,\quad\forall\>i,$ (125)
$k_{i}^{-}(\mathbf{B}^{*})$ $=\sum_{t=1}^{T}\frac{y_{i}w_{t}}{x_{i}z_{t}+y_{i}w_{t}}=\sum_{t=1}^{T}p_{it}^{-}=\langle k_{i}^{-}\rangle,\quad\forall\>i,$ (126)
$\kappa_{t}^{+}(\mathbf{B}^{*})$ $=\sum_{i=1}^{N}\frac{x_{i}z_{t}}{x_{i}z_{t}+y_{i}w_{t}}=\sum_{i=1}^{N}p_{it}^{+}=\langle\kappa_{t}^{+}\rangle,\quad\forall\>t,$ (127)
$\kappa_{t}^{-}(\mathbf{B}^{*})$ $=\sum_{i=1}^{N}\frac{y_{i}w_{t}}{x_{i}z_{t}+y_{i}w_{t}}=\sum_{i=1}^{N}p_{it}^{-}=\langle\kappa_{t}^{-}\rangle,\quad\forall\>t.$ (128)
The system above can be solved only numerically (see Appendix B).
Appendix B Numerical optimisation of likelihood functions
In order to numerically solve the systems of equations defining the bSCM, we can follow the guidelines provided in Vallarano et al. (2021): more specifically, we will adapt the iterative recipe provided there to our setting. First, let us notice that, upon posing $x_{i}\equiv e^{-\alpha_{i}}$, $y_{i}\equiv e^{-\gamma_{i}}$, $\forall\>i$ and $z_{t}\equiv e^{-\delta_{t}}$, $w_{t}\equiv e^{-\eta_{t}}$, $\forall\>t$, the system of equations defining the bSCM can be re-written as
$x_{i}$ $=\frac{k_{i}^{+}(\mathbf{B^{*}})}{\sum_{t=1}^{T}\dfrac{z_{t}}{x_{i}z_{t}+y_{i}w_{t}}}\>\Longrightarrow\>x_{i}^{(n)}=\frac{k_{i}^{+}(\mathbf{B^{*}})}{\sum_{t=1}^{T}\dfrac{z_{t}^{(n-1)}}{x_{i}^{(n-1)}z_{t}^{(n-1)}+y_{i}^{(n-1)}w_{t}^{(n-1)}}},\quad\forall\>i,$ (129)
$y_{i}$ $=\frac{k_{i}^{-}(\mathbf{B^{*}})}{\sum_{t=1}^{T}\dfrac{w_{t}}{x_{i}z_{t}+y_{i}w_{t}}}\>\Longrightarrow\>y_{i}^{(n)}=\frac{k_{i}^{-}(\mathbf{B^{*}})}{\sum_{t=1}^{T}\dfrac{w_{t}^{(n-1)}}{x_{i}^{(n-1)}z_{t}^{(n-1)}+y_{i}^{(n-1)}w_{t}^{(n-1)}}},\quad\forall\>i,$ (130)
$z_{t}$ $=\frac{\kappa_{t}^{+}(\mathbf{B^{*}})}{\sum_{i=1}^{N}\dfrac{x_{i}}{x_{i}z_{t}+y_{i}w_{t}}}\>\Longrightarrow\>z_{t}^{(n)}=\frac{\kappa_{t}^{+}(\mathbf{B^{*}})}{\sum_{i=1}^{N}\dfrac{x_{i}^{(n-1)}}{x_{i}^{(n-1)}z_{t}^{(n-1)}+y_{i}^{(n-1)}w_{t}^{(n-1)}}},\quad\forall\>t,$ (131)
$w_{t}$ $=\frac{\kappa_{t}^{-}(\mathbf{B^{*}})}{\sum_{i=1}^{N}\dfrac{y_{i}}{x_{i}z_{t}+y_{i}w_{t}}}\>\Longrightarrow\>w_{t}^{(n)}=\frac{\kappa_{t}^{-}(\mathbf{B^{*}})}{\sum_{i=1}^{N}\dfrac{y_{i}^{(n-1)}}{x_{i}^{(n-1)}z_{t}^{(n-1)}+y_{i}^{(n-1)}w_{t}^{(n-1)}}},\quad\forall\>t;$ (132)
in order for our iterative recipe to converge, an appropriate vector of initial conditions needs to be chosen; here, we have opted the following one: $x_{i}=k_{i}^{+}(\mathbf{B}^{*})/\sqrt{L^{+}(\mathbf{B}^{*})}$, $\forall\>i$, $y_{i}=k_{i}^{-}(\mathbf{B}^{*})/\sqrt{L^{-}(\mathbf{B}^{*})}$, $\forall\>i$, $z_{t}=\kappa_{t}^{+}(\mathbf{B}^{*})/\sqrt{L^{+}(\mathbf{B}^{*})}$, $\forall\>t$ and $w_{t}=\kappa_{t}^{-}(\mathbf{B}^{*})/\sqrt{L^{-}(\mathbf{B}^{*})}$, $\forall\>t$.
Moreover, we have adopted two, different stopping criteria: the first one is a condition on the Euclidean norm of the vector of differences between the values of the parameters at subsequent iterations, i.e. $||\Delta\vec{\theta}||_{2}=\sqrt{\sum_{i=1}^{N}(\Delta\theta_{i})^{2}}\leq 10^{-8}$; the second one is a condition on the maximum number of iterations of our iterative algorithm, set to $10^{3}$.
The accuracy of our method in estimating the constraints has been evaluated by computing the maximum absolute error (MAE), defined as
bMAE $=\max_{i,t}\left\{|k_{i}^{+}(\mathbf{B^{*}})-\langle k_{i}^{+}\rangle|,\>|k_{i}^{-}(\mathbf{B^{*}})-\langle k_{i}^{-}\rangle|,\>|\kappa_{t}^{+}(\mathbf{B^{*}})-\langle\kappa_{t}^{+}\rangle|,\>|\kappa_{t}^{-}(\mathbf{B^{*}})-\langle\kappa_{t}^{-}\rangle|\right\}$ (133)
(i.e. as the infinite norm of the difference between the vector of the empirical values of the constraints and the vector of their expected values) and the maximum relative error (MRE), defined as
bMRE $=\max_{i,t}\left\{\frac{|k_{i}^{+}(\mathbf{B^{*}})-\langle k_{i}^{+}\rangle|}{k_{i}^{+}(\mathbf{B^{*}})},\>\frac{|k_{i}^{-}(\mathbf{B^{*}})-\langle k_{i}^{-}\rangle|}{k_{i}^{-}(\mathbf{B^{*}})},\>\frac{|\kappa_{t}^{+}(\mathbf{B^{*}})-\langle\kappa_{t}^{+}\rangle|}{\kappa_{t}^{+}(\mathbf{B^{*}})},\>\frac{|\kappa_{t}^{-}(\mathbf{B^{*}})-\langle\kappa_{t}^{-}\rangle|}{\kappa_{t}^{-}(\mathbf{B^{*}})}\right\}$ (134)
(i.e. as the infinite norm of the relative difference between the vector of the empirical values of the constraints and the vector of their expected values).
Table 1, providing information on the basic statistics of 555, illustrative time series coming from the HCP-Young Adult dataset shows the time employed by our algorithm to converge as well as its accuracy in reproducing the constraints defining the bSCM: our method is, overall, fast and accurate, as the absolute errors are around $10^{-3}$ and the relative errors are around $10^{-6}$, the time employed to achieve such an accuracy being always less than a minute.
Basic statistics of time series bSCM $N$ $T$ $B^{+}$ $B^{-}$ $c^{+}$ $c^{-}$ MAE MRE Time (s) HCP106016 116 2.400 142.407 135.993 0.512$\simeq$ 0.512≃ 0.512 0.488$\simeq$ 0.488≃ 0.488 $\simeq 1.10\cdot 10^{-3}$ $\simeq 9.94\cdot 10^{-7}$ 28.8$\simeq$ 28.8≃ 28.8 HCP857263 116 2.400 136.376 142.024 0.490$\simeq$ 0.490≃ 0.490 0.510$\simeq$ 0.510≃ 0.510 $\simeq 1.10\cdot 10^{-3}$ $\simeq 9.90\cdot 10^{-7}$ 33.3$\simeq$ 33.3≃ 33.3 HCP125525 116 2.400 139.054 139.346 0.499$\simeq$ 0.499≃ 0.499 0.501$\simeq$ 0.501≃ 0.501 $\simeq 1.26\cdot 10^{-3}$ $\simeq 1.17\cdot 10^{-6}$ 37.8$\simeq$ 37.8≃ 37.8 HCP100408 116 2.400 137.804 140.596 0.495$\simeq$ 0.495≃ 0.495 0.505$\simeq$ 0.505≃ 0.505 $\simeq 1.30\cdot 10^{-3}$ $\simeq 1.15\cdot 10^{-6}$ 37.8$\simeq$ 37.8≃ 37.8 HCP130316 116 2.400 137.196 141.204 0.493$\simeq$ 0.493≃ 0.493 0.507$\simeq$ 0.507≃ 0.507 $\simeq 1.97\cdot 10^{-3}$ $\simeq 1.89\cdot 10^{-6}$ 37.7$\simeq$ 37.7≃ 37.7
Table 1: Performance of our fixed-point algorithm to solve the system of equations defining the bSCM on 555, illustrative time series coming from the HCP-Young Adult dataset (each pedex refers to a subject). $N$ is the number of brain regions, $T$ is the duration of the related time series, $B^{+}$ is the overall number of positive values, $B^{-}$ is the overall number of negative values, $c^{+}=B^{+}/(N\cdot T)$ is the density of positive values and $c^{-}=B^{-}/(N\cdot T)$ is the density of negative values. Our fixed-point algorithm leads to MAE values around $10^{-3}$ and MRE values of the order of $10^{-6}$ in less than forty seconds.
Appendix C Degree distributions of cortical and subcortical areas
As anticipated in the main text, subcortical areas tend to have a larger negative degree than cortical areas. This is true in both a subject-wise (see fig. 16) and a pooled fashion (see fig. 17)
Figure 16: Complementary cumulative distribution functions (CCDFs) of positive ($k^{+}$) and negative ($k^{-}$) degrees for cortical and subcortical areas, per benchmark. Analyses have been carried out in a subject-wise fashion (i.e. for subjects 106016106016106016 and 857263857263857263).
Figure 17: Complementary cumulative distribution functions (CCDFs) of positive ($k^{+}$) and negative ($k^{-}$) degrees for cortical and subcortical areas, per benchmark. Analyses have been carried out in a pooled fashion (i.e. across the 100100100 subjects considered for the present analysis).

### v3

@@
 Abstract
-Many complex systems - be they financial, natural or social - are composed by units - such as stocks, neurons or agents - whose joint activity can be represented as a multivariate time series.
-An issue of both practical and theoretical importance concerns the possibility of inferring the presence of a static relationships between any two units solely from their dynamic state.
-The present contribution aims at providing an answer within the frame of traditional hypothesis testing.
-Briefly speaking, our suggestion is that of linking any two units if behaving in a sufficiently similar way.
-To achieve such a goal, we project a multivariate time series onto a signed graph, by i) comparing the empirical properties of the former with those expected under a suitable benchmark and ii) linking any two units with a positive (negative) edge in case the corresponding series share a significantly large number of concordant (discordant) values.
-To define our benchmarks, we adopt an information-theoretic approach that is rooted into the constrained maximisation of Shannon entropy, a procedure inducing an ensemble of multivariate time series that preserves some of the empirical properties on average while randomising everything else.
+Many complex systems - be they financial, natural, or social - are composed of units - such as stocks, neurons, or agents - whose joint activity can be represented as a multivariate time series.
+An issue of both practical and theoretical importance concerns the possibility of inferring the presence of a static relationship between any two units solely from their dynamic state.
+The present contribution aims at tackling such an issue within the frame of traditional hypothesis testing: briefly speaking, our suggestion is that of linking any two units if behaving in a sufficiently similar way.
+To achieve such a goal, we project a multivariate time series onto a signed graph by i) comparing the empirical properties of the former with those expected under a suitable benchmark and ii) linking any two units with a positive (negative) edge in case the corresponding series shares a significantly large number of concordant (discordant) values.
+To define our benchmarks, we adopt an information-theoretic approach that is rooted into the constrained maximisation of Shannon entropy, a procedure inducing an ensemble of multivariate time series that preserves some of the empirical properties on average, while randomising everything else.
 We showcase the possible applications of our method by addressing one of the most timely issues in the domain of neurosciences, i.e.
-that of determining if brain networks are frustrated or not - and, in case, to what extent.
-As our results suggest, this is indeed the case, the structure of the negative subgraph being more prone to inter-subject variability than the complementary, positive subgraph.
-At the mesoscopic level, instead, the minimisation of the Bayesian Information Criterion instantiated with the Signed Stochastic Block Model reveals that brain areas gather into modules aligning with the statistical variant of the Relaxed Balance Theory.
+that of determining if brain networks are frustrated or not, and, if so, to what extent.
+As our results suggest, this is indeed the case, with the major contribution to the underlying negative subgraph coming from the subcortical structures (and, to a lesser extent, from the limbic regions).
+At the mesoscopic level, the minimisation of the Bayesian Information Criterion, instantiated with the Signed Stochastic Block Model, reveals that brain areas gather into modules aligning with the statistical variant of the Relaxed Balance Theory.
 pacs:
@@
 I Introduction
-Time series are ubiquitous and versatile representations of dynamical complex systems, be they financial, natural or social.
+Time series are ubiquitous and versatile representations of dynamical complex systems, be they financial, natural, or social.
 A major challenge is that of inferring the underlying dependencies between units from the (often non-linear) dynamics characterising them.
-Being capable of transforming a time series into a symbolic sequence, while preserving essential information about the ‘original’ generative process, is of practical relevance, as the dynamical properties of the former can be revealed by focusing on the structural properties of the latter Onnela et al.
-(2004); Sinha et al.
-(2010); Bouchaud and Potters (2003).
-To this aim, sequential data are often represented as traditional graphs, with nodes replacing individual units (such as stocks, neurons or agents) and edges reflecting their activity-driven interactions.
-This choice plays a major role in the neuroscientific domain, where network-based approaches have revolutionised our understanding of the brain functioning, especially through the study of functional connectivity networks derived from Blood Oxygen Level Dependent (BOLD) fMRI signals.
-Rooted in graph theory and statistical physics, this perspective, often referred to as network neuroscience, aims at revealing the organisational principles of cognition and behaviour Kulkarni and Bassett (2025) by modelling the brain as an integrated system of interacting regions Bassett and Sporns (2017).
-A review of the approaches that have been proposed so far to extract relationships between nodes directly from the time series of the original units is provided below.
+Being capable of transforming a time series into a symbolic sequence, while preserving essential information about the ‘original’ generative process, is of practical relevance, as the dynamical properties of the former can be revealed by focusing on the structural properties of the latter Onnela2004; Sinha2010; Bouchaud2003.
+To this aim, sequential data are often represented as traditional graphs, with nodes replacing individual units (such as stocks, neurons, or agents) and edges reflecting their activity-driven interactions.
+This choice plays a major role in the neuroscientific domain, where network-based approaches have revolutionised our understanding of brain functioning, especially through the study of functional connectivity networks derived from Blood Oxygen Level-Dependent (BOLD) fMRI signals.
+Rooted in graph theory and statistical physics, this perspective, often referred to as network neuroscience, aims at revealing the organisational principles of cognition and behaviour kulkarni2024principlesbrainnetworkorganization by modelling the brain as an integrated system of interacting regions bassettnetwork2017.
+In the following subsection, a review of the approaches proposed so far to extract relationships between nodes directly from the time series of the original units is provided.
 I.1 From time series to proximity and from proximity to graphs
-Given a set of time series, the most common approach is that of extracting a proximity matrix from it and, then, a graph from the proximity matrix.
-Although several definitions of proximity have been proposed [^1] We redirect the interested reader to the reviews Cliff et al.
-(2023); Liu et al.
-(2024), where hundreds of proximity measures (such as the dynamic time warping, the mutual information and the transfer entropy) are discussed and compared., the commonest choice is identifying it with the notion of pairwise correlation, as returned by the calculation of the Pearson correlation coefficient
+Given a set of time series, the most common approach is that of extracting a proximity matrix that is subsequently used for defining a graph.
+Although several definitions of proximity have been proposed [^1] We redirect the interested reader to the reviews Cliff2023; Liu2024, where hundreds of proximity measures (such as the dynamic time warping, the mutual information and the transfer entropy) are discussed and compared., the commonest choice is identifying it with the notion of pairwise correlation, as returned by the calculation of the Pearson correlation coefficient
 $C_{ij}=\frac{\sum_{t=1}^{T}(X_{it}-\underline{X}_{i})\cdot(X_{jt}-\underline{X}_{j})}{\sqrt{\sum_{t=1}^{T}(X_{it}-\underline{X}_{i})^{2}}\cdot\sqrt{\sum_{t=1}^{T}(X_{jt}-\underline{X}_{j})^{2}}},$ (1)
@@
 represent the average value of series $\mathbf{X}_{i}=\{X_{it}\}_{t=1}^{T}$ and $\mathbf{X}_{j}=\{X_{jt}\}_{t=1}^{T}$, respectively.
-A less popular, but still worth to be mentioned, definition of proximity is the one resting upon the notion of partial correlation Hampson et al.
-(2002); Fransson and Marrelec (2008).
-If we indicate the matrix of pairwise correlations with $\mathbf{C}$, the generic entry of the matrix of partial correlations, then, reads $\hat{J}_{ik}=J_{ik}/\sqrt{J_{ii}\cdot J_{kk}}$, with $\mathbf{J}=\mathbf{C}^{-1}$ being the so-called precision matrix.
-Estimations of the functional connectivity (FC) based on partial correlations are affected by a smaller inter-subject variability Brier et al.
-(2015); Rahim et al.
-(2019), as they have been shown to resemble more closely the so-called structural connectivity (SC) patterns Huang et al.
-(2010); Salvador et al.
-(2005); Lee et al.
-(2011); Santucci et al.
-(2025).
+A less popular, but still worth mentioning, definition of proximity is based on the notion of partial correlation Hampson2002; Fransson2008.
+If we keep indicating the matrix of pairwise correlations with $\mathbf{C}$, the generic entry of the matrix of partial correlations, then, reads $\hat{J}_{ik}=J_{ik}/\sqrt{J_{ii}\cdot J_{kk}}$, with $\mathbf{J}=\mathbf{C}^{-1}$ being the so-called precision matrix.
+Estimations of the functional connectivity (FC) based on partial correlations are affected by a smaller inter-subject variability Brier2015; Rahim2019, as they have been shown to resemble more closely the so-called structural connectivity (SC) patterns Huang2010; Salvador2005; Lee2011; Santucci2025.
 I.1.1 Fixed thresholding algorithms
-The simplest approach of this kind is based on the interpretation of a correlation matrix (be it partial or not) as a fully connected weighted matrix.
-In order to obtain non-trivial structures within such a context, sparsification procedures become, thus, necessary, the simplest one being that of thresholding correlations Bullmore and Sporns (2009).
-This procedure ultimately boils down to remove the weaker connections - usually the most affected by experimental noise Rubinov and Sporns (2010) - while retaining the stronger ones.
-In the neuroscientific literature, one of the earliest attempts of this kind was the one pursued in Stam (2004), where measures of synchronisation between brain regions, derived from magnetoencephalographic (MEG) recordings, were considered to extract frequency band-specific networks by fixing their average degree; a similar attempt was pursued in Micheloyannis et al.
-(2009), where electroencephalographic (EEG) recordings were, instead, considered.
-In the financial literature, the network resulting from the application of a threshold is also known with the name of asset graph Sinha et al.
-(2010); Onnela et al.
-(2004); Heimo et al.
-(2009).
-In Onnela et al.
-(2004), the authors derived an asset graph from the correlation matrix of a set of stocks by filling an empty graph, edge by edge, starting from the one corresponding to the largest correlation and stopping at the $N$-th largest correlation (with $N$ being the number of nodes).
+The simplest approach of this kind is based on interpreting a correlation matrix (be it partial or not) as a fully-connected, weighted matrix.
+Non-trivial structures can thus be obtained by implementing sparsification procedures, the simplest of which is thresholding correlations Bullmore2009.
+This procedure ultimately boils down to remove the weaker connections - usually the most affected by experimental noise Rubinov2010 - while retaining the stronger ones; still, as reported by the authors of MacMahon2015, thresholding may prevent the detection of a modular organisation, as the correlations within modules may be cut out even if significantly stronger than those between modules.
+In the neuroscientific literature, one of the earliest algorithms of this kind was the one pursued in Stam2004, where measures of synchronisation between brain regions, derived from magnetoencephalographic (MEG) recordings, were considered and frequency band-specific networks extracted by fixing the average degree; a similar attempt was pursued in Micheloyannis2007, where electroencephalographic (EEG) recordings were, instead, considered.
+In the financial literature, the network resulting from the application of a threshold is also known as the asset graph Sinha2010; Onnela2004; Heimo2009.
+In Onnela2004, the authors derived an asset graph from the correlation matrix of a set of stocks by filling an empty graph, edge by edge, starting from the one corresponding to the largest correlation and stopping at the $N$-th largest correlation (with $N$ being the number of nodes).
 Upon doing so, the considered asset graph is comparable in size with the most popular asset tree, although being less constrained by requests such as the one of forbidding cycles.
-As a last comment, we report the concern expressed by the authors of MacMahon and Garlaschelli (2015), according to whom thresholding may prevent the detection of a modular organisation, as the correlations within modules may be cut out even if significantly stronger than those between modules.
-Since letting the network density drive the thresholding procedure appears as arbitrary as setting a threshold directly on correlations, it has been proposed to determine it by requiring that the resulting network remains connected.
-Such a data-driven procedure removes the weaker correlations until the percolation threshold is reached, i.e.
-the value in correspondence of which the resulting network disconnects Alexander-Bloch et al.
-(2010); Gallos et al.
-(2012).
+Since letting the network density drive the thresholding procedure appears as arbitrary as setting a threshold directly on correlations, the proposal to determine the resulting network by requiring it to remain connected has been advanced.
+Such a data-driven procedure removes the weaker correlations until the value in correspondence of which the resulting network disconnects is reached Alexander-Bloch2010; Gallos2012.
 Although this kind of analysis has been popularised with the name of percolation analysis, the aforementioned value is, technically speaking, the connectivity threshold.
 A different, yet related, criterion prescribes to adopt the threshold that maximises the value $\text{APL}_{\text{ER}}=\ln N/\ln\overline{k}$, with $\overline{k}=\sum_{i=1}^{N}k_{i}/N$, i.e.
-the (expected) average path length of an Erdös-Rényi network whose connectance matches the empirical one Hayasaka and Laurienti (2010); Marangolo et al.
-(2016); alternatively, one may adopt the threshold [^2] As the authors themselves explicitly state, their criteria individuate a whole range of link densities, more precisely from 0.4$\simeq$ 0.4≃ 0.4 to 0.5$\simeq$ 0.5≃ 0.5: below 0.4$\simeq$ 0.4≃ 0.4 graphs fragment and above 0.5$\simeq$ 0.5≃ 0.5 their topology becomes increasingly random.
-induced by the request of having a small-world network with no isolated nodes Bassett et al.
-(2008); Lynall et al.
-(2010).
-If the previous approaches prescribe to stop pruning correlations once the percolation threshold is reached, a second approach focuses on the mesoscopic organisation of the network and prescribes to consider a correlation-induced distance333 Examples are provided the one reading $d_{ij}=\sqrt{2(1-C_{ij})}$, mapping distances between time series into the interval [0,2][0,2][ 0 , 2 ], and the one reading $d_{ij}=\sqrt{(1-C_{ij})/2}$, mapping distances between time series into the interval [0,1][0,1][ 0 , 1 ].
+the (expected) average path length of an Erdös-Rényi network whose connectance matches the empirical one Hayasaka2010; Marangolo2016; alternatively, one may adopt the threshold [^2] As the authors themselves explicitly state, their criterion individuate a whole range of link densities, more precisely from $\simeq 0.4$ to $\simeq 0.5$: below $\simeq 0.4$ graphs fragment and above $\simeq 0.5$ their topology becomes increasingly random.
+induced by the request of having a small-world network with no isolated nodes Bassett2008; Lynall2010.
+If the previous approaches prescribe to stop pruning correlations once the connectivity threshold is reached, a second approach focuses on the mesoscopic organisation of the network and prescribes to consider a correlation-induced distance [^3] Examples are provided the one reading $d_{ij}=\sqrt{2(1-C_{ij})}$, mapping distances between time series into the interval $[0,2]$, and the one reading $d_{ij}=\sqrt{(1-C_{ij})/2}$, mapping distances between time series into the interval $[0,1]$.
 In the first case, 0 corresponds to linearly correlated time series, 2 corresponds to anti-correlated time series and $\sqrt{2}$ corresponds to the case where a linear correlation is not found; in the second case, 0 corresponds to linearly correlated time series, 1 corresponds to anti-correlated time series and $\sqrt{2}/2$ corresponds to the case where a linear correlation is not found., draw the corresponding dendrogram [^4] A dendrogram is constructed by clustering objects in a hierarchical fashion: to this aim, a distance between clusters is needed.
-Some of these ‘linkages’ algorithms induce an ultra-metric distance on the data Rammal et al.
-(1986) that has been argued to induce distances progressively less similar to the original ones as higher levels of the taxonomic tree are climbed: according to this intuition, such a method is more reliable to determine the low-level structure of the tree (induced by the strongest correlations) than the high-level one (induced by the weakest correlations) MacMahon and Garlaschelli (2015); Milligan (1979).
-and cut it in correspondence of a properly-defined threshold Mantegna and Stanley (1999); Tola et al.
-(2008).
-A possible criterion to individuate it is that of maximising the modular character of the resulting network structure, analysed upon running one of the many algorithms that have been proposed to this aim (Louvain, Infomap, Surprise, etc.) Newman (2006); Bordier et al.
-(2017); Marchese et al.
-(2022).
-Examples from the financial literature are provided by Di Matteo et al.
-(2004), whose authors cut the dendrogram induced by 343434 weekly time series of interest rates at a distance which is halfway between the one induced by $C_{ij}=0$ and the one induced by $C_{ij}=1$, thus recovering 666 clusters throughout the whole period considered there, i.e.
-198219821982-199719971997, and Tola et al.
-(2008), whose authors compare the performance of a similar clustering algorithm with that of an algorithm based on Random Matrix Theory in singling out groups of assets with comparable returns.
+Some of these ‘linkages’ algorithms induce an ultra-metric distance on the data Rammal1986 that has been argued to induce distances progressively less similar to the original ones as higher levels of the taxonomic tree are climbed.
+According to this intuition, such a method is more reliable to determine the low-level structure of the tree (induced by the strongest correlations) than the high-level one (induced by the weakest correlations) MacMahon2015; Milligan1979.
+and cut it in correspondence of a properly-defined threshold Mantegna1999; Tola2008.
+A possible criterion to individuate it is that of maximising the modular character of the resulting network structure, analysed upon running one of the many algorithms that have been proposed to this aim (Louvain, Infomap, Surprise, etc.) Newman2006; Bifone2017; Marchese2022.
+Examples from the financial literature are provided by DiMatteo2004, whose authors cut the dendrogram induced by $34$ weekly time series of interest rates at a distance which is halfway between the one induced by $C_{ij}=0$ and the one induced by $C_{ij}=1$, thus recovering $6$ clusters throughout the whole period considered there, i.e.
+19821982-$1997$, and Tola2008, whose authors compare the performance of a similar clustering algorithm with that of an algorithm based on Random Matrix Theory in singling out groups of assets with comparable returns.
 I.1.2 Progressive thresholding algorithms
-Since individuating a unique threshold rises doubts about the arbitrariness of such a choice - which may significantly impact the results - alternatives have been proposed to make this approach a bit more principled, such as monitoring the changes in the resulting network structure upon rising the threshold itself.
-The earliest attempt, pursued in Eguíluz et al.
-(2005), consisted in identifying three threshold values, i.e.
-0.60.60.6, 0.70.70.7 and 0.80.80.8: topological traits (averaged over the three sets of threshold-specific configurations) such as the scale-freeness of the distribution of functional connections and the small-worldness of the resulting network structure were, however, found to be robust against progressive thresholding.
-In Honey et al.
-(2009), the authors investigated how the relationships between FC and SC evolve upon progressively removing all FC values, the aim being that of quantifying the extent to which the SC values can be inferred from the ‘surviving’ FC ones.
-Studying the progressive engagement of brain regions has been proven useful to reveal the hierarchical character of the mouse brain Bardella et al.
-(2016) as well as distinguish healthy subjects from schizophrenic patients Mastrandrea et al.
-(2021).
+Since individuating a unique threshold raises doubts about the arbitrariness of such a choice - which may significantly impact the results - alternatives have been proposed to make this approach a bit more principled, such as monitoring the changes in the resulting network structure upon raising the threshold itself.
+The earliest attempt, pursued in Eguiluz2005, consisted in identifying three threshold values, i.e.
+$0.6$, $0.7$, and $0.8$: topological traits (averaged over the three sets of threshold-specific configurations) such as the scale-freeness of the distribution of functional connections and the small-worldness of the resulting network structure were, however, found to be robust against progressive thresholding.
+In Honey2009, the authors investigated how the relationships between FC and SC evolve upon progressively removing all FC values, the aim being that of quantifying the extent to which the SC values can be inferred from the ‘surviving’ FC ones.
+Studying the progressive engagement of brain regions has been proven useful to reveal the hierarchical character of the mouse brain Bardella2016, as well as to distinguish healthy subjects from schizophrenic patients Mastrandrea2021.
 I.1.3 Filtering algorithms
@@
 Filtering correlations via the MST.
-The most popular filtering approach returns the so-called Maximum Spanning Tree (MST), a technique retaining the $N-1$ largest correlations and ensuring that i) loops are discarded; ii) each node is reachable from any other node via a path that is the widest between its endpoints; iii) the total weight is greater than, or equal to, the total weight of any other spanning tree.
-It can be obtained upon running the Kruskal’s algorithm and is often employed to distinguish the connector areas (i.e.
+The most popular filtering approach returns the so-called Maximum Spanning Tree (MST), a technique retaining the N1N-1 largest correlations while ensuring that i) loops are discarded; ii) each node is reachable from any other node via a path that is the widest between its endpoints; iii) the total weight is greater than, or equal to, the total weight of any other spanning tree.
+This method can be applied upon running the Kruskal’s algorithm and is often employed to distinguish the connector areas (i.e.
 the most central ones) from the provincial areas (i.e.
-the most peripheral ones) Rammal et al.
-(1986); Mantegna and Stanley (1999); Caldarelli et al.
-(2012); Mastrandrea et al.
-(2017, 2021).
+the most peripheral ones) Rammal1986; Mantegna1999; Caldarelli2012; Mastrandrea2017; Mastrandrea2021.
 The Kruskal’s algorithm can be slightly modified by adding the extra condition that no edge between any two nodes that are already connected is allowed.
@@
 a set of disconnected trees usually interpreted as the backbone of a set of communities [^6] Notice that such a technique differs from the one prescribing to run the Kruskal’s algorithm on a disconnected graph, as it discards the edges that would join the different components of the MST, beside those generating loops..
-Both the MST and the MSF are rather different from their random counterparts Bardella et al.
-(2016), their structural organisation providing a topological signature of the arising of diseases such as schizophrenia Mastrandrea et al.
-(2021).
-Coming to the financial literature, the authors of Bonanno et al.
-(2004) filter the correlation matrix induced by a set of daily time series of stock returns from the NYSE throughout the whole period considered there, i.e.
-198719871987-199819981998, by calculating the Minimum Spanning Tree defined by the $N-1$ smallest correlation-induced distances (see footnote 3).
+Both the MST and the MSF are rather different from their random counterparts Bardella2016, their structural organisation providing a topological signature of the emergence of diseases such as schizophrenia Mastrandrea2021.
+Coming to the financial literature, the authors of Bonanno2004 filter the correlation matrix induced by a set of daily time series of stock returns from the NYSE throughout the whole period considered there, i.e.
+19871987-$1998$, by calculating the Minimum Spanning Tree defined by the N1N-1 smallest correlation-induced distances (see footnote 3).
 Filtering correlations via the PMFG.
-An alternative approach, discarding less information, returns the so-called Planar Maximally Filtered Graph (PMFG) Aste et al.
-(2005); Tumminello et al.
-(2005).
+An alternative approach, discarding less information, returns the so-called Planar Maximally Filtered Graph (PMFG) Aste2005; Tumminello2005.
 The criterion informing such a method prescribes to retain the correlations ensuring that the resulting network is a planar graph, i.e.
-a graph that can be drawn with no links crossing each other (i.e.
-intersecting each other not in correspondence of a node).
+a graph that can be drawn with no links crossing each other - more precisely, one in which edges may intersect only at common endpoints (nodes).
 Since the PMFG always contains the MST, the former complements the latter by providing additional (and not just different) information.
-The PMFG has also been described as the simplest instance of a more general procedure embedding high-dimensional data into lower-dimensional manifolds having a controllable number of ‘holes’ called ‘genus’.
-The PMFG corresponds to the case in which the genus is zero Tumminello et al.
-(2005).
+The PMFG has also been described as the simplest instance of a more general procedure embedding high-dimensional data into lower-dimensional manifolds with a controllable number of ‘holes’ called ‘genus’.
+The PMFG corresponds to the case in which the genus is zero Tumminello2005.
 Although versatile - the method has been extended in a variety of ways, e.g.
-to produce a nested hierarchy of time series Song et al.
-(2011, 2012); Morales et al.
-(2014) - it suffers from some degree of arbitrariness too, affecting the properties of the postulated, approximating structure: the value of the genus must, in fact, be fixed a priori; besides, there is no obvious reason why time series should find a natural embedding into a bidimensional plane.
+to produce a nested hierarchy of time series Song2011; Song2012; Morales2014 - it suffers from some degree of arbitrariness too, affecting the properties of the postulated, approximating structure: the value of the genus must, in fact, be fixed a priori; besides, there is no obvious reason why time series should find a natural embedding into a bidimensional plane.
 I.1.4 Testing correlations
 A more principled approach to extract information from a correlation matrix is that of comparing (some of) its properties with the ones expected under a properly defined benchmark, the aim of this class of techniques being that of identifying non-random collective information about the set of time series induced by $\mathbf{C}$.
-Quite interestingly, these techniques do not discard negative correlations, known to be a consequence of spatial and temporal heterogeneities induced by non-trivial, neural-mediated, hemodynamic mechanisms and contributing to the stability of resting-state brain networks, by affecting the synchronisation of neural activity Saberi et al.
-(2021).
-Although isolated attempts exist - such as the one proposed in Salvador et al.
-(2005), whose authors test the null hypothesis that the average value of the distribution of partial correlations, for each pair of areas and across subjects, is zero - methods of this kind can be split into those based on Random Matrix Theory (RMT) and those based on Entropy Maximization (EM).
+Interestingly, these techniques do not discard negative correlations, which are known to be a consequence of spatial and temporal heterogeneities induced by non-trivial, neural-mediated hemodynamic mechanisms affecting the synchronisation of neural activity, thus contributing to the stability of resting-state brain networks Saberi2021.
+Although isolated attempts exist - such as the one proposed in Salvador2005, whose authors test the null hypothesis that the average value of the distribution of partial correlations, for each pair of areas and across subjects, is zero - methods of this kind can be split into those based on Random Matrix Theory (RMT) and those based on Entropy Maximization (EM).
 Null models based on RMT.
-One of the most popular models studied in the context of RMT is the Wishart ensemble Wishart (1948); Livan et al.
-(2018) which, in its simplest form, leads to the Marcenko-Pastur distribution Marčenko and Pastur (1967), i.e.
+One of the most popular models studied in the context of RMT is the Wishart ensemble Wishart1948; Livan2018 which, in its simplest form, leads to the Marcenko-Pastur distribution MP, i.e.
 the distribution of the eigenvalues of a correlation matrix induced by $N$ completely random time series of duration $T$ in the limits $N\rightarrow+\infty$ and $T\rightarrow+\infty$ with $1<T/N<+\infty$, and reads
@@
 with $\sqrt{\lambda_{\pm}}=1\pm\sqrt{N/T}$ and $\rho(\lambda)=0$ otherwise.
-As a consequence, while the eigenvalues of an empirical correlation matrix falling within the range $[\lambda_{-},\lambda_{+}]$ can be imputable to random noise, any other eigenvalue can be deemed as representing meaningful structure in the data.
-As a consequence, the random component of any empirical correlation matrix can be identified with the tensor reading
-$|\mathbf{C}^{(r)}=\sum_{i:\lambda_{i}\leq\lambda_{+}}\lambda_{i}\ket{v_{i}}\bra{v_{i}}$ | (4)
-where |$\ket{v_{i}}$ is the eigenvector associated with the eigenvalue [^7] The eigenvalues smaller than $\lambda_{-}$ are usually also included. $\lambda_{i}$.
-A typical feature of the spectrum of empirical correlation matrices is that of having an eigenvalue $\lambda_{m}$ that is (much) larger than any other eigenvalue and whose corresponding eigenvector |$\ket{v_{m}}$ has all positive components: such a ‘market mode’ contributes to the random component of an empirical correlation matrix $\mathbf{C}$ with a tensor reading $|\mathbf{C}^{(m)}=\lambda_{m}\ket{v_{m}}\bra{v_{m}}$ |.
+As a consequence, while the eigenvalues of an empirical correlation matrix falling within the range $\lambda_{-},\lambda_{+}]$ can be imputable to random noise, any other eigenvalue can be deemed as representing meaningful structure in the data.
+Thus, the random component of any empirical correlation matrix can be identified with the tensor reading
+$|\mathbf{C}^{(r)}=\sum_{i:\lambda_{i}\leq\lambda_{+}}\lambda_{i}\ket{v_{i}}\bra{v_{i}}$ (4)
+[…]
+[…]
 The structured part of $\mathbf{C}$, reading
-$\mathbf{C}^{(g)}=\mathbf{C}-\mathbf{C}^{(r)}-\mathbf{C}^{(m)}=\sum_{i:\lambda_{+}<\lambda_{i}<\lambda_{m}}\lambda_{i}\ket{v_{i}}\bra{v_{i}},$ (5)
+$|,\mathbf{C}^{(g)}=\mathbf{C}-\mathbf{C}^{(r)}-\mathbf{C}^{(m)}=\sum_{i:\lambda_{+}<\lambda_{i}<\lambda_{m}}\lambda_{i}\ket{v_{i}}\bra{v_{i}}$, (5)
 is, then, inserted into the novel definition of modularity
 $Q(\mathbf{\sigma})=\frac{1}{|\mathbf{C}|}\sum_{i=1}^{N}\sum_{j=1}^{N}C^{(g)}_{ij}\delta_{\sigma_{i}\sigma_{j}}$ (6)
-with $|\mathbf{C}|=\sum_{i=1}^{N}\sum_{j=1}^{N}C_{ij}$ and whose maximisation can be proven to lead to clusters of nodes that are positively correlated while nodes belonging to different clusters are negatively correlated MacMahon and Garlaschelli (2015).
-In Plerou et al.
-(2002), the authors employ models of this kind to analyse correlations of time series of stock market returns: by comparing them with null models based on RMT, the authors find that most eigenvalues are compatible with statistical noise, those capturing meaningful correlations (and being linked to economic factors such as market-wide movements or sectoral influences) being the few, large, deviating ones.
-In Utsugi et al.
-(2004), the authors find that the distribution of the eigenvalues of the correlation matrix induced by a set of daily time series of stock prices from the TSE is stable, throughout the whole period of time considered there, i.e.
-199319931993-200120012001.
-In Potters et al.
-(2005), the authors explore how RMT can be applied to better estimate financial risk.
-In Zema et al.
-(2025), the authors build upon such a scheme by developing a novel approach for portfolio management.
-Null models based on EM. A different, yet related, approach considers null models induced by entropy maximisation Masuda et al. (2018): in this case, one seeks to maximise the differential entropy
-[…]
+with $|\mathbf{C}|=\sum_{i=1}^{N}\sum_{j=1}^{N}C_{ij}$ and whose maximisation can be proven to lead to clusters of nodes that are positively correlated while nodes belonging to different clusters are negatively correlated MacMahon2015.
+In Plerou2002, the authors employ models of this kind to analyse correlations of time series of stock market returns: by comparing them with null models based on RMT, the authors find that most eigenvalues are compatible with statistical noise, those capturing meaningful correlations (and being linked to economic factors such as market-wide movements or sectoral influences) being the few, large, deviating ones.
+In Utsugi2004, the authors find that the distribution of the eigenvalues of the correlation matrix induced by a set of daily time series of stock prices from the TSE is stable throughout the whole period of time considered there, i.e.
+19931993-$2001$.
+In Potters2005, the authors explore how RMT can be applied to better estimate financial risk.
+In Zema2024, the authors build upon such a scheme by developing a novel approach for portfolio management.
+Null models based on EM. A different, yet related, approach considers null models induced by entropy maximisation Masuda2018: in this case, one seeks to maximise the differential entropy
+$S=-\int P(\mathbf{X})\ln P(\mathbf{X})d\mathbf{X},$ (7)
 where $\mathbf{X}=\{X_{it}\}$ is an $N\times T$ (i.e.
@@
 The resulting distribution $P(\mathbf{X})$ is a multivariate normal distribution defined by a covariance matrix $\mathbf{C}^{(r)}$ obeying a Wishart distribution with $T$ degrees of freedom.
-In Kojaku and Masuda (2019), the authors employ the same null model to filter covariance matrices, inducing networks in which any two nodes are linked if their correlation is found to be significantly larger than expected: to this aim, the matrix $\bm{\Delta}=\mathbf{C}-\mathbf{C}^{(r)}$, defined as the difference between the empirical and the random covariance matrix, is numerically determined by maximising the log-likelihood function [^8] To avoid overfitting, the log-likelihood is penalised by an adaptive-Lasso term whose parameters are determined by minimising the Bayesian Information Criterion (BIC).
-A package named ‘SCOLA’ (an acronym stating for ‘Sparse Networks Construction from Correlational Data with Lasso’) has been also released.
+In Kojaku2019, the authors employ the same null model to filter covariance matrices, inducing networks in which any two nodes are linked if their correlation is found to be significantly larger than expected: to this aim, the matrix $\bm{\Delta}=\mathbf{C}-\mathbf{C}^{(r)}$, defined as the difference between the empirical and the random covariance matrix, is numerically determined by maximising the log-likelihood function [^8] To avoid overfitting, the log-likelihood is penalised by an adaptive-Lasso term whose parameters are determined by minimising the Bayesian Information Criterion (BIC).
+A package named ‘SCOLA’ (an acronym standing for ‘Sparse Networks Construction from Correlational Data with Lasso’) has also been released.
 $\mathcal{L}=\ln P(\mathbf{X}|\mathbf{C})$.
 Bayesian inference.
-Another way for inferring a sparse graph from a set of time series, while avoiding thresholding, is that of inferring a precision matrix with $\ell_{1}$ penalization, where the value of the lasso regularisation parameter is usually set by cross-validation.
-This is equivalent to choosing the precision matrix that maximises the posterior distribution given the data, using a Gaussian likelihood and a prior encoding the $\ell_{1}$ constraint Ibáñez Berganza et al.
-(2023); Santucci et al.
-(2025); Varoquaux and Craddock (2013).
-Such an approach leads to a sparse, positive-definite precision matrix, interpretable as an adjacency matrix.
+A third way for inferring a sparse graph from a set of time series, while avoiding thresholding, is that of inferring a precision matrix with $\ell_{1}$ penalisation, where the value of the lasso regularisation parameter is usually set by cross-validation.
+This is equivalent to choosing the precision matrix that maximises the posterior distribution given the data, using a Gaussian likelihood and a prior encoding the $\ell_{1}$ constraint ibanez2023; Santucci2025; VAROQUAUX2013.
+Such an approach leads to a sparse, positive-definite precision matrix, which is interpretable as an adjacency matrix.
 I.2 From univariate time series to graphs
-Different (classes of) approaches exist that seek to extract a network from one or more time series while ignoring (the intermediate step represented by) the proximity matrix [^9] We redirect the interested reader to the reviews Silva et al.
-(2021); Zou et al.
-(2019), discussing the differences between the existing (classes of) algorithms and comparing their performances..
+Different (classes of) approaches that seek to extract a network from one or more time series while ignoring (the intermediate step represented by) the proximity matrix exist [^9] We redirect the interested reader to the reviews Silva2021; Zou2019, discussing the differences between the existing (classes of) algorithms and comparing their performances..
 I.2.1 Visibility-induced mapping
@@
 Different types of visibility algorithms exist, designed to obtain undirected/directed, binary/weighted graphs.
-[…]
-A variant of the aforementioned procedure has been proposed in Luque et al.
-(2009), where the concept of horizontal visibility - as opposed to the one of natural visibility - is defined.
+The first algorithm of this kind was proposed in Lacasa2008 and is based on the idea of representing each observation as a vertical bar whose height equals the numerical value of the observation itself: if the tops of any two bars can be directly connected via a ‘visibility line’, the corresponding nodes are, then, connected in the resulting network.
+The authors of Lacasa2008 have shown that visibility graphs (VGs) inherit several properties from the original time series, e.g.
+VGs are always connected because each observation (node) is visible by, at least, its temporal (structural) neighbours; hubs are induced by (local) maxima, since the corresponding bars see more nodes than other points; the presence of hubs usually induces a modular structure; a periodic time series induces a network whose degree distribution is characterised by a number of peaks that is smaller than its period $T$, as the degree of node $i$ will be the same of the nodes $i\pm T$, $i\pm 2T$, $i\pm 3T$ and so on; a white noise process induces a (random) network whose degree distribution is exponential while a fractal time series induces a (scale-free) network whose degree distribution is power-law.
+Visibility algorithms are easy to implement, computationally fast, and parameter-free; besides, they have been employed to diagnose Alzheimer’s disease automatically, achieving an accuracy of $\simeq 98\%$ Ahmadlou2010.
+A variant of the aforementioned procedure has been proposed in Luque2009, where the concept of horizontal visibility - as opposed to the one of natural visibility - is defined.
 A horizontal visibility graph (HVG) is constructed by considering (only) horizontal visibility lines: as a consequence, the HVG is always a subgraph of the VG, hence containing less information, although its construction is faster.
-HVGs have been employed to analyse seismic signals Telesca and Lovallo (2012), the volatility of returns of financial time series Zhang et al.
-(2015), heartbeat rates Xie et al.
-(2019) and for classification tasks Li et al.
-(2018).
-A third variant is defined by implementing the concept of directed visibility, whose introduction has been justified by noticing that time has a natural direction: in Lacasa et al.
-(2012), the authors adopted the rule according to which any pair of time steps $t_{i}<t_{j}$ populates a directed horizontal visibility graph (DHVG) with an edge $i\rightarrow j$ and employed DHVGs to quantify time series irreversibility.
-A fourth variant is defined by implementing the concept of limited penetrable visibility, proposed in Zhou et al.
-(2012) to reduce the level of noise accompanying data to a larger extent than that achievable by using VGs and HVGs.
-In this case, any two nodes are connected if the visibility line joining the tops of the two associated bars crosses a number of ‘obstacles’ that is smaller than, or equal to, a fixed value $L$ naturally, the limited penetrable visibility graph (LPVG) reduces to the VG when $L=0$, i.e.
+HVGs have been employed to analyse seismic signals Telesca2012, the volatility of returns of financial time series Zhang2015, heartbeat rates Xie2019, and for classification tasks Li2018.
+A third variant is defined by implementing the concept of directed visibility, whose introduction has been justified by noticing that time has a natural direction: in Lacasa2012, the authors adopted the rule according to which any pair of time steps $t_{i}<t_{j}$ populates a directed horizontal visibility graph (DHVG) with an edge $i\rightarrow j$ and employed DHVGs to quantify time series irreversibility.
+A fourth variant is defined by implementing the concept of limited penetrable visibility, proposed in Zhou2012 to reduce the level of noise accompanying data to a larger extent than that achievable by using VGs and HVGs.
+In this case, any two nodes are connected if the visibility line joining the tops of the two associated bars crosses a number of ‘obstacles’ that is smaller than, or equal to, a fixed value $L$ - naturally, the limited penetrable visibility graph (LPVG) reduces to the VG when $L=0$, i.e.
 when there are no ‘obstacles’.
-LPVGs have been employed to analyse EEG signals of people in different conditions (affected by Alzheimer’s disease Wang et al.
-(2016) or under manual acupuncture Pei et al.
-(2014)).
+LPVGs have been employed to analyse EEG signals of people in different conditions (affected by Alzheimer’s disease Wang2016 or under manual acupuncture Pei2014).
 A fifth variant is defined by implementing the concept of weighted visibility.
-The method, proposed in Supriya et al.
-(2016), prescribes to enrich each edge of a DHVG with a weight matching the ‘view angle’ of the observing node, its sign with respect to the horizontal axis indicating if the values of the time series increase (if positive) or decrease (if negative); since such a method allows for negative weights, the absolute value of the angle is often considered.
+The method, proposed in Supriya2016, prescribes to enrich each edge of a DHVG with a weight matching the ‘view angle’ of the observing node, its sign with respect to the horizontal axis indicating if the values of the time series increase (if positive) or decrease (if negative); since such a method allows for negative weights, the absolute value of the angle is often considered.
 The authors employed the method to identify EEG signals associated with epilepsy.
-A last variant is defined by implementing the concept of differential visibility, realised by removing the edge set of the HVG from the edge set of the VG Zhu et al. (2014).
+A last variant is defined by implementing the concept of differential visibility, which is realised by removing the edge set of the HVG from the edge set of the VG Zhu2014.
 I.2.2 Transition-induced mapping
 According to such a mapping, connections are established between (symbols assigned to) different portions of univariate time series, intended to represent states between which transitions occur: methods of this kind, in fact, represent a time series as a directed, weighted graph.
-The mapping introduced in Campanharo et al.
-(2011) is based on a partition of the distribution of the values assumed by a time series in quantiles, to each of which a symbol is assigned - the number of quantiles is an arbitrary parameter whose choice should depend on the length of the series itself Shirazi et al.
-(2009).
-The variant of this method introduced in Weng et al.
-(2017) allows ‘memory networks’ to be constructed by assigning the attribute $t$ to the edge directed from $i$ to $j$ if the transition between the corresponding symbols occurs at time step $t$.
-A second variant is the one introduced in Campanharo and Ramos (2016), where the concept of renewal intervals, i.e.
+The mapping introduced in Campanharo2011 is based on a partition of the distribution of the values assumed by a time series in quantiles, to each of which a symbol is assigned - the number of quantiles is an arbitrary parameter whose choice should depend on the length of the series itself Shirazi2009.
+The variant of this method introduced in Weng2017 allows ‘memory networks’ to be constructed by assigning the attribute $t$ to the edge directed from $i$ to $j$ if the transition between the corresponding symbols occurs at time step $t$.
+A second variant is the one introduced in Campanharo2016, where the concept of renewal intervals, i.e.
 transitions between symbols posed at various temporal distances, is introduced.
-The mapping introduced in Gao and Jin (2009) is named Coarse-Grained Phase Space Graphs and requires the definition of a time window $\omega$ to determine the number of values of the series to be represented as an $\omega$-dimensional vector, whose components are numerically determined by adopting a rule similar to the quantiles-based one.
-A different, yet related, mapping is the one introduced in Small (2013), named Ordinal Partition Transition Networks.
+The mapping introduced in Gao2009 is named Coarse-Grained Phase Space Graphs and requires the definition of a time window $\omega$ to determine the number of values of the series to be represented as an $\omega$-dimensional vector, whose components are numerically determined by adopting a rule similar to the quantiles-based one.
+A different, yet related, mapping is the one introduced in Small2013, named Ordinal Partition Transition Networks.
 Here, the time series is embedded in a $\omega$-dimensional space where, starting from each point of the series, the current and the following $\omega-1$ numbers, interspaced by a time delay $\tau$, are considered (for instance, setting $\omega=3$ and $\tau=2$, the $i$-th embedding vector will contain the values at $i$, $i+2$, and $i+4$).
-The embedding vectors are, then, associated with patterns induced by the rank of their values and each ordinal pattern is represented as a node of the network: nodes $i$ and $j$ are connected if the corresponding patterns occur at subsequent time steps (i.e.
+The embedding vectors are, then, associated with patterns induced by the rank of their values, and each ordinal pattern is represented as a node of the network: nodes $i$ and $j$ are connected if the corresponding patterns occur at subsequent time steps (i.e.
 if $t_{j}=t_{i}+1$).
-A third variant of the same kind of mapping is the one introduced in Stephen et al.
-(2015) and named Visibility Graphlets Networks.
+A third variant of the same kind of mapping is the one introduced in Mutua2015 and named Visibility Graphlets Networks.
 This method requires the definition of a time window $\omega$ to determine the number of values to be considered: the time steps from $t$ to $t+\omega$ are, then, mapped into a directed visibility graph (DVG).
-Such a procedure is repeated for each time step and the DVGs having the same adjacency matrix are associated with the same symbol, i.e.
+Such a procedure is repeated for each time step, and the DVGs having the same adjacency matrix are associated with the same symbol, i.e.
 one of the nodes of the resulting network.
@@
 According to such a mapping, connections are established between (groups of) time steps of a univariate time series, according to measures of distance/similarity between them.
-The first method of this kind was introduced in Zhang et al.
-(2006) to represent pseudo-periodic time series.
+The first method of this kind was introduced in Zhang2006 to represent pseudo-periodic time series.
 First, it segments a time series into cycles (e.g.
-dictated by a periodic behaviour) but without requiring them to have the same length; then, correlations between pairs of cycles are computed - by shifting the shortest one, if needed - and the largest correlation value is compared against a threshold: if found to be larger, an edge between the two cycles, represented as nodes in the resulting network, is established.
-An alternative way of employing correlations to implement a proximity-induced mapping is introduced in Yang and Yang (2008) and prescribes to compute the correlations between the (either overlapping or non-overlapping) portions into which a time series can be partitioned: the nodes representing them are linked if the correlation between the corresponding portions exceeds a certain threshold.
-A different class of methods is the one gathering the so-called Recurrence Networks (RNs) and is based on the idea that the points of a time series are recurrent if their distance, in a properly defined embedding space, is ‘short enough’.
-Similarly to the prescription characterising the Ordinal Partition Transition Networks algorithm, each time step is mapped into a point of an embedding space.
+dictated by a periodic behaviour) without requiring them to have the same length; then, correlations between pairs of cycles are computed - by shifting the shortest one, if needed - and the largest correlation value is compared against a threshold: if found to be larger, an edge between the two cycles, represented as nodes in the resulting network, is established.
+An alternative method for employing correlations to implement a proximity-induced mapping is introduced in Yang2008 and prescribes computing the correlations between the (either overlapping or non-overlapping) portions into which a time series can be partitioned: the nodes representing them are then linked if the correlation between the corresponding portions exceeds a certain threshold.
+A different class of methods is the one gathering the so-called Recurrence Networks (RNs), which is based on the idea that the points of a time series are recurrent if their distance, in a properly defined embedding space, is ‘short enough’.
+Similarly to the prescription characterising the Ordinal Partition Transition Networks algorithm, each time step is mapped to a point of an embedding space.
 Distances are, then, computed by using one of the many norms that can be defined.
-The way the embedding space is analysed distinguishes RNs in three sub-categories: i) the first one, introduced in Small et al.
-(2009), is composed by the so-called k-Nearest Neighbour Networks and requires to connect each node with the $k$ nodes whose corresponding points in the embedding space are closest to its corresponding point - to be noticed that such a variant has been defined for directed networks; ii) the second one, introduced in Xu et al.
-(2008), is composed by the so-called Adaptive Nearest Neighbour Networks and implements the undirected version of the aforementioned recipe; iii) the third one, introduced in Donner et al.
-(2010), is composed by the so-called $\epsilon$-Recurrence Networks ($\epsilon$-RNs) and prescribes to link the nodes whose corresponding points lie at a distance smaller than $\epsilon$: in symbols, the generic element of an $\epsilon$-RN matrix is defined as
+The way the embedding space is analysed distinguishes RNs in three sub-categories: i) the first one, introduced in Small2009, is composed by the so-called k-Nearest Neighbour Networks and requires to connect each node with the $k$ nodes whose corresponding points in the embedding space are closest to its corresponding point - to be noticed that such a variant has been defined for directed networks; ii) the second one, introduced in Xu2008, is composed by the so-called Adaptive Nearest Neighbour Networks and implements the undirected version of the aforementioned recipe; iii) the third one, introduced in Donner2010, is composed by the so-called $\epsilon$-Recurrence Networks ($\epsilon$-RNs) and prescribes to link the nodes whose corresponding points lie at a distance smaller than $\epsilon$: in symbols, the generic element of an $\epsilon$-RN matrix is defined as
 $\mathbf{RN}_{tt^{\prime}}^{(x,x)}=\begin{cases}1,&\text{if\>}\left\|\bm{x}_{t}-\bm{x}_{t^{\prime}}\right\|\leq\epsilon\\ 0,&\text{otherwise}\end{cases}$ (11)
@@
 I.3.1 Single-layer mapping
-This first kind of mapping collects generalisations of the aforementioned RNs, hereby named Cross Recurrence Networks (CRNs), Inter-System Recurrence Networks (ISRNs) and Joint Recurrence Networks (JRNs) Feldhoff et al.
-(2012, 2013); Romano et al.
-(2004).
+This first kind of mapping collects generalisations of the aforementioned RNs, hereby named Cross Recurrence Networks (CRNs), Inter-System Recurrence Networks (ISRNs), and Joint Recurrence Networks (JRNs) Feldhoff2012; Feldhoff2013; Romano2004.
 CRNs are defined for pairs of time series: after mapping the two time series $\mathbf{X}$ and $\mathbf{Y}$ into the same space, a biadjacency matrix is induced as follows:
 $\mathbf{CRN}_{tt^{\prime}}^{(x,y)}=\begin{cases}1,&\text{if\>}\left\|\bm{x}_{t}-\bm{y}_{t^{\prime}}\right\|\leq\epsilon\\ 0,&\text{otherwise}\end{cases}$ (12)
-with $\bm{x}_{t}$ and $\bm{y}_{t^{\prime}}$ indicating the points of the chosen embedding space representing the values of the time series $X_{t}$ and $Y_{t^{\prime}}$ notice that the two time series do not need to be of the same length, hence inducing a rectangular table Marwan et al.
-(2002); Zbilut et al.
-(1998).
+with $\bm{x}_{t}$ and $\bm{y}_{t^{\prime}}$ indicating the points of the chosen embedding space representing the values of the time series $X_{t}$ and $Y_{t^{\prime}}$ - notice that the two time series do not need to be of the same length, hence inducing a rectangular table Marwan2002; Zbilut1998.
 Such a recipe can be further extended to deal with multivariate time series: in this case, we have the so-called ISRNs, inducing adjacency matrices like
 $\mathbf{ISRN}=\begin{pmatrix}\mathbf{RN}^{(1,1)}&\mathbf{CRN}^{(1,2)}&\dots&\mathbf{CRN}^{(1,m)}\\ \mathbf{CRN}^{(2,1)}&\mathbf{RN}^{(2,2)}&\dots&\mathbf{CRN}^{(2,m)}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{CRN}^{(m,1)}&\mathbf{CRN}^{(m,2)}&\dots&\mathbf{RN}^{(m,m)}\end{pmatrix}$, (13)
-constituted by (matrices representing the) RNs along the diagonal and by (matrices representing the) CRNs off the diagonal Feldhoff et al. (2012).
-An alternative recipe is represented by the JRNs Romano et al. (2004); Feldhoff et al. (2013), inducing adjacency matrices like the one reading
+constituted by (matrices representing the) RNs along the diagonal and by (matrices representing the) CRNs off the diagonal Feldhoff2012.
+An alternative recipe is represented by the JRNs Romano2004; Feldhoff2013, inducing adjacency matrices like the one reading
 $\mathbf{JRN}_{tt^{\prime}}^{(x,y)}=\begin{cases}1,&\text{if\>}\left\|\bm{x}_{t}-\bm{x}_{t^{\prime}}\right\|\leq\epsilon_{x}\cap\left\|\bm{y}_{t}-\bm{y}_{t^{\prime}}\right\|\leq\epsilon_{y}\\ 0,&\text{otherwise}\end{cases}$ (14)
 that, similarly to what has been observed for the ISRNs, is block-wise, each diagonal block referring to a single time series and the off-diagonal blocks referring to pairs of time series.
-A generalisation of the Ordinal Partition Transition Networks for multivariate time series was introduced in Zhang et al.
-(2017).
-The method requires a binarisation of the univariate time series, letting an increasing trend [^10] Defined as ‘the value of the series at time $t$ is larger than the value of the series at time $t-1$’.
-being represented with a 111 and a decreasing trend being represented with a 0: upon doing so, $2^{N}$ combinations, with $N$ being the number of time series, are possible for each time step, $t$; finally, each combination is assigned a node in the resulting network, any two nodes connected by a directed edge representing combinations found at subsequent steps.
-For deterministic and periodic time series, some of the nodes remain isolated; random processes are, instead, associated with densely (if not completely) connected graphs.
+A generalisation of the Ordinal Partition Transition Networks for multivariate time series was introduced in Zhang2017.
+The method requires a binarisation of the univariate time series, letting an increasing trend [^10] Defined as ‘the value of the series at time $t$ is larger than the value of the series at time t1t-1’.
+being represented with a $1$ and a decreasing trend being represented with a 0: upon doing so, $2^{N}$ combinations, with $N$ being the number of time series, are possible for each time step, $t$; finally, each combination is assigned a node in the resulting network, any two nodes connected by a directed edge representing combinations found at subsequent steps.
+For deterministic and periodic time series, some of the nodes remain isolated; in contrast, random processes are associated with densely (if not completely) connected graphs.
 I.3.2 Multiple-layer mapping
-This second kind of mapping, instead, associates a multivariate time series with a multi-layer network.
-A first example of such a mapping is provided by Lacasa et al.
-(2015): here, a multiplex visibility graph (MVG) is obtained by mapping each univariate time series into a graph by employing a VG; afterwards, the MVG is aggregated into a single-layer weighted graph, each node representing a layer-specific VG and each weight proxying the similarity between the corresponding layers.
-Such a method has been employed to analyse fMRI data Sannino et al.
-(2017).
-The authors of Eroglu et al. (2018) follow the same pipeline but replace VGs with RNs, thus employing Multiplex Recurrence Networks (MRNs) to study paleoclimatic time series.
+This second kind of mapping, on the other hand, associates a multivariate time series with a multi-layer network.
+A first example of such a mapping is provided by Lacasa2015: here, a multiplex visibility graph (MVG) is obtained by mapping each univariate time series into a graph by employing a VG; afterwards, the MVG is aggregated into a single-layer weighted graph, each node representing a layer-specific VG and each weight proxying the similarity between the corresponding layers.
+Such a method has been employed to analyse fMRI data Sannino2017.
+The authors of Eroglu2018 follow the same pipeline but replace VGs with RNs, thus employing Multiplex Recurrence Networks (MRNs) to study paleoclimatic time series.
 I.4 From multivariate time series to
 validated graphs
-The most recent approaches emphasise the importance of probabilistic (null, as well as generative) models to benchmark network features and disentangle fundamental, structural properties from statistical artifacts Varoquaux and Craddock (2013); Váša and Mišić (2022).
+The most recent approaches emphasise the importance of probabilistic (null, as well as generative) models to benchmark network features and disentangle fundamental, structural properties from statistical artifacts VAROQUAUX2013; vasanull2022.
 These models play a crucial role in understanding whether empirical features (e.g.
-a modular structure, a small-world architecture, a rich-club organisation) are indicators of a genuine degree of self-organisation or simple by-products of lower-order constraints [^11] More explicitly, the practice of employing graph-theoretical estimations for inter-subject analyses without comparing them with a proper benchmark gives ‘[…] a fairly unspecific characterisation of the brain, being fragile to noise […] Another caveat is that […] correlation matrices display small-world properties […] by construction […] This observation highlights the need for well-defined null hypotheses […] but also for controlled recovery of brain functional connectivity going beyond empirical correlation matrices […]’ Varoquaux and Craddock (2013)..
-In the present contribution, we focus on the issue of representing sequential data as relational data, by approaching the question how can a set of time series be represented as a traditional graph?
+a modular structure, a small-world architecture, a rich-club organisation) are indicators of a genuine degree of self-organisation or simple by-products of lower-order constraints [^11] More explicitly, the practice of employing graph-theoretical estimations for inter-subject analyses without comparing them with a proper benchmark gives ‘[…] a fairly unspecific characterisation of the brain, being fragile to noise […] Another caveat is that […] correlation matrices display small-world properties […] by construction […] This observation highlights the need for well-defined null hypotheses […] but also for controlled recovery of brain functional connectivity going beyond empirical correlation matrices […]’ VAROQUAUX2013..
+In the present contribution, we focus on the issue of representing sequential data as relational data, approaching the question how can a set of time series be represented as a traditional graph?
 from the perspective of traditional hypothesis testing.
-After comparing the statistical properties of the former against those expected under a suitable benchmark, our suggestion is that of projecting the set of time series onto a signed graph, by linking any two series with a positive (negative) edge in case they share a statistically significant number of concordant (discordant) values.
-More formally, we adopt an information-theoretic approach that is rooted into the constrained maximisation of Shannon entropy, a procedure inducing an ensemble of multivariate time series that preserves some of their empirical properties on average while randomising everything else Park and Newman (2004); Garlaschelli and Loffredo (2008); Squartini and Garlaschelli (2011).
-Our method deals with signed data Gallo et al.
-(2024a, b, 2025): one may, thus, wonder how relevant is the information provided by signs within the neuroscientific context.
-Several studies have recently investigated the role played by the negative correlations characterising resting-state fMRI data (rs-fMRI): the authors of Saberi et al.
-(2021) show that negative correlations are essential for the stability of resting-state brain networks against perturbations or changes; the authors of Saberi et al.
-(2022) notice that the interplay between positive and negative correlations induces frustrated patterns preventing the brain from reaching a state of minimum energy and enhancing its adaptability to dynamical changes; finally, the authors of Demertzi et al.
-(2022) show that the default mode network (DMN) works ‘in opposition’ with the various task-positive networks, a supposedly crucial feature to maintain cognitive functions and consciousness.
-Let us, however, stress that pre-processing procedures like the Global Signal Regression may artificially induce negative correlations, an evidence further stressing the importance of such a preliminary phase when cleaning neuroscientific data Murphy et al.
-(2009).
+After comparing the statistical properties of the former against those expected under a suitable benchmark, our suggestion is to project the set of time series onto a signed graph by linking any two series with a positive (negative) edge in case they share a statistically significant number of concordant (discordant) values.
+More formally, we adopt an information-theoretic approach that is rooted in the constrained maximisation of Shannon entropy, a procedure inducing an ensemble of multivariate time series that preserves some of their empirical properties on average while randomising everything else Park2004; Garlaschelli2008; Squartini2011.
+As our method deals with signed data gallo2024testing; gallo2024assessing; gallo2025statistically, one may wonder how relevant the information provided by signs is within the neuroscientific framework.
+Several studies have recently investigated the role played by the negative correlations characterising resting-state fMRI data (rs-fMRI): the authors of Saberi2021 show that negative correlations are essential for the stability of resting-state brain networks against perturbations or changes; the authors of Saberi2022 notice that the interplay between positive and negative correlations induces frustrated patterns preventing the brain from reaching a state of minimum energy and enhancing its adaptability to dynamical changes; finally, the authors of Demertzi2022 show that the default mode network (DMN) works ‘in opposition’ with the various task-based networks, a supposedly crucial feature to maintain cognitive functions and consciousness.
+Let us, however, stress that pre-processing procedures, such as Global Signal Regression, may artificially induce negative correlations, further emphasizing the importance of such a preliminary phase when cleaning neuroscientific data Murphy2009.
 II Setting up the formalism
@@
 are the average value and standard deviation of series $i$.
-Since $w_{it}$ can be either positive or negative, to ease mathematical manipulations let us employ Iverson’s brackets - a notation ensuring all quantities of interest to be non-negative Gallo et al.
-(2024a) - and define
+Since $w_{it}$ can be either positive or negative, to ease mathematical manipulations, let us employ Iverson’s brackets - a notation ensuring all quantities of interest are non-negative gallo2024testing - and define
 $w_{it}^{-}$ $\equiv|w_{it}|\cdot[w_{it}<0],\quad\forall\>i,t$ (17)
@@
 in an entry-wise fashion - or, more compactly, $\mathbf{B}\equiv\mathbf{B}^{+}-\mathbf{B}^{-}$ and $|\mathbf{B}|\equiv\mathbf{B}^{+}+\mathbf{B}^{-}$.
-The present analysis solely focuses on binarised time series: beside easing the mathematical tractability of the latter, binarisation has been also shown to lead to better clustering results Thomas Yeo et al.
-(2011).
-Figure 1: Infographics illustrating our analysis pipeline.
+The present analysis solely focuses on binarised time series: besides easing the mathematical tractability of the latter, binarisation has also been shown to lead to better clustering results yeo2011organization.
+Adopting the terminology introduced for bipartite networks, we will also refer to the rectangular matrix $\mathbf{B}$ as the biadjacency matrix associated with the considered set of time series.
+Figure 1: Infographics illustrating the pipeline of our analysis.
 Pictorial representation of the pipeline we follow in the present contribution, from the registration of brain activity to the identification of communities in statistically validated signed projections: (a) rs-fMRI signals are recorded; (b) pairwise interactions between brain regions are considered; (c) the scalar product of each pair of (standardised and binarised) time series is calculated to quantify their signature; (d) its empirical value is validated against a null model to remove statistical noise; (e) a significantly large positive (negative) signature induces a positive (negative) link between the two involved brain regions; (f) a BIC-based community detection reveals the modules partitioning our statistically validated signed projections.
+The first brain image was adapted from Canva (additional elements were added by the authors), while the last one was obtained through the Enigma Toolbox Lariviere2021.
 II.1 Local properties and dependencies of time series
-In analogy with the quantities considered for the study of bipartite signed networks, let us define the total number of positive values as
-$B^{+}\equiv\sum_{i=1}^{N}\sum_{t=1}^{T}b_{it}^{+}=\sum_{i=1}^{N}k_{i}^{+}=\sum_{t=1}^{T}\kappa_{t}^{+};$ (25)
-the expression above suggests the definition of two (sets of) local positive quantities, i.e.
+Let us start by defining a number of quantities of interest for our study.
+In analogy with the properties traditionally considered to analyse bipartite signed networks, the total number of positive values reads
+$B^{+}\equiv\sum_{i=1}^{N}\sum_{t=1}^{T}b_{it}^{+}=\sum_{i=1}^{N}k_{i}^{+}=\sum_{t=1}^{T}\kappa_{t}^{+},$ (25)
+an expression further suggesting the definition of two (sets of) local positive quantities, i.e.
 $k_{i}^{+}\equiv\sum_{t=1}^{T}b_{it}^{+},$ (26)
-the positive degree of series $i$, capturing the total number of positive entries characterising the $i$-th series and providing information about its auto-correlation, and
+or positive degree of series $i$, capturing the total number of positive entries characterising the $i$-th series and providing information about its auto-correlation, and
 $\kappa_{t}^{+}\equiv\sum_{i=1}^{N}b_{it}^{+},$ (27)
-the positive degree at time $t$, capturing the total number of positive entries characterising the $t$-th time step and providing information about the correlations between the members of the whole set of time series; analogously,
+or positive degree at time $t$, capturing the total number of positive entries characterising the $t$-th time step and providing information about the correlations between the members of the whole set of time series.
+Analogously, the total number of negative values reads
 $B^{-}\equiv\sum_{i=1}^{N}\sum_{t=1}^{T}b_{it}^{-}=\sum_{i=1}^{N}k_{i}^{-}=\sum_{t=1}^{T}\kappa_{t}^{-}$ (28)
-with obvious meaning of the symbols. The advantage of adopting Iverson’s brackets is that of ensuring that each quantity is, now, computed on a matrix with positive entries, i.e. is positive as well.
+with obvious meaning of the symbols.
+The density of positive and negative values, reading $c^{+}\equiv B^{+}/(N\cdot T)$ and $c^{-}\equiv B^{-}/(N\cdot T)$ and referred to as positive and negative bipartite connectance, remain naturally defined.
 II.2 Concordant and discordant temporal motifs
@@
 patterns capturing the ‘disagreement’ between any two series.
-Any two nodes whose associated time series exhibit a significantly large number of concordant motifs will be connected by a +1+1+ 1, while any two nodes whose associated time series exhibit a significantly large number of discordant motifs will be connected by a 1-1- 1.
-Lastly, we would like to stress that our projection scheme is applicable only in the absence of null values - a condition that is naturally satisfied by BOLD time series, representing the focus of this study.
+Any two nodes whose associated time series exhibit a significantly large number of concordant motifs will be connected by a +1+1, while any two nodes whose associated time series exhibit a significantly large number of discordant motifs will be connected by a 1-1.
+We would also like to stress that our projection scheme is applicable only in the absence of null values - a condition that is naturally satisfied by BOLD time series, representing the focus of this study.
 II.3 A scheme for the validated projection of multivariate time series
@@
     apply a multiple hypothesis testing procedure and insert a (binary) link between the nodes $i$ and $j$ if and only if $p_{ij}$ is found to be less than, or equal to, the threshold value $p_{th}$ (see section V).
-Let us, now, describe each step in detail.
+Let us now describe each step in detail.
 III Computing the similarity of any
 two series
-The first step of our method prescribes to measure the degree of similarity of series $i$ and $j$.
-To this aim, let us consider the quantity named signature Gallo et al.
-(2024a), defined as the scalar product of the two series: upon defining $\mathbf{b}_{i}$ as the $i$-th row of the matrix $\mathbf{B}$, the signature reads
+The first step of our method prescribes measuring the degree of similarity of series $i$ and $j$.
+To this aim, let us consider the quantity named signature gallo2024testing, defined as the scalar product of the two series: upon defining $\mathbf{b}_{i}$ as the $i$-th row of the matrix $\mathbf{B}$, the signature reads
 $S_{ij}\equiv\mathbf{b}_{i}\cdot\mathbf{b}_{j}$ $=\sum_{t=1}^{T}b_{it}b_{jt}$
@@
 two series
-The second step of our method prescribes to evaluate the statistical significance of the similarity of any two series. To this aim, let us, first, notice that
-$-T\leq S_{ij}\leq T$ (37)
+The second step of our method prescribes evaluating the statistical significance of the similarity of any two series. To this aim, let us first notice that
+$T\leq S_{ij}\leq T$ (37)
 where $S_{ij}=-T$ if $C_{ij}=0$ (i.e.
-if each dyadic motif is composed by a 1-1- 1 and a +1+1+ 1) and $S_{ij}=T$ if $D_{ij}=0$ (i.e.
-if each dyadic motif is composed by either two 1-1- 1s or two +1+1+ 1s) and, then, clarify that we will employ ERGs defined by linear constraints, hence treat links as independent random variables.
+if each dyadic motif is composed by a 1-1 and a +1+1) and $S_{ij}=T$ if $D_{ij}=0$ (i.e.
+if each dyadic motif is composed by either two 1-1s or two +1+1s) and, then, clarify that we will employ ERGs defined by linear constraints, hence treat links as independent random variables.
 IV.1 Homogeneous benchmark
@@
 satisfying the relationships
-1 11 $=q^{+}+q^{-},$ (40)
+$1$ $=q^{+}+q^{-},$ (40)
 $\langle b_{it}b_{jt}\rangle$ $=q^{+}-q^{-},$ (41)
@@
 Let us, now, notice that $S_{ij}$ is a sum of i.i.d.
-Bernoulli random variables [^13] The finite scheme obeyed by $C_{ijt}-D_{ijt}$ is the same finite scheme obeyed by $b_{it}b_{jt}$., the outcomes of each elementary event being, in fact, 1-1- 1 and +1+1+ 1 instead of 0 and +1+1+ 1.
+Bernoulli random variables [^13] The finite scheme obeyed by $C_{ijt}-D_{ijt}$ is the same finite scheme obeyed by $b_{it}b_{jt}$., the outcomes of each elementary event being, in fact, 1-1 and +1+1 instead of 0 and +1+1.
 Such a consideration allows us to find the probability distribution obeyed by $S_{ij}$ quite straightforwardly, by starting from the one describing $C_{ij}$ alone, i.e.
@@
 $P(S_{ij}=s)=\binom{T}{\frac{T+s}{2}}(q^{+})^{\frac{T+s}{2}}(q^{-})^{\frac{T-s}{2}}$ (44)
-[…]
+that ranges from T-T to $T$.
+Let us, now, convince ourselves that this transformation is indeed the correct one, by providing some explicit examples: if nodes $i$ and $j$ establish $k=0$ concordant motifs, they establish $T$ discordant motifs and the signature reads s=Ts=-T; if nodes $i$ and $j$ establish $k=T/2$ concordant motifs, they establish the same number of discordant motifs and the signature reads $s=0$; if nodes $i$ and $j$ establish $k=T$ concordant motifs, they establish zero discordant motifs and the signature reads $s=T$.
+In other words, the aforementioned change of variable transforms a discrete distribution whose support extends from 0 to $T$ into a discrete distribution whose support extends from T-T to $T$: more intuitively, we have moved from considering sequences of 0s and +1+1s to considering sequences of 1-1s and +1+1s, where each +1+1 can be identified with a concordant motif and each 1-1 can be identified with a discordant motif.
 A comment about the parity of $T$ and $s$ is needed: let us notice that, in case $T$ is odd, $s$ itself must be odd, assuming the values $s=-T,-T+2,-T+4\dots T$ (corresponding to $k=0,1,2\dots T$); in case $T$ is even, $s$ itself must be even, assuming the same values as above in correspondence of the same values of $k$.
@@
 satisfying the relationships
-1 11 $=q_{ijt}^{+}+q_{ijt}^{-},$ (47)
+$1$ $=q_{ijt}^{+}+q_{ijt}^{-},$ (47)
 $\langle b_{it}b_{jt}\rangle$ $=q_{ijt}^{+}-q_{ijt}^{-},$ (48)
@@
 $=4q_{ijt}^{+}q_{ijt}^{-}.$ (49)
+Figure 2: Infographics illustrating our validation procedure.
+Probability distribution of the signature and its Gaussian approximation: while the left panel provides a graphical answer to the question is the empirical value of the signature significantly different from the one expected under the chosen benchmark?, the right panel provides a graphical answer to the question is the deviation negative (hence, the signature is significantly smaller than expected) or positive (hence, the signature is significantly larger than expected)?
+The red area corresponds to the region of validation of the negative links, induced by pairs of series ‘in counterphase’ most of the time; the blue area corresponds to the region of validation of the positive links, induced by pairs of series ‘in phase’ most of the time.
 Since $S_{ij}$ is, now, a sum of i.n.i.d. Bernoulli random variables, the probability distribution obeyed by it is the Poisson-binomial
@@
 and so on.
-In words, if nodes $i$ and $j$ establish $k=0$ concordant motifs, they establish $T$ discordant motifs and the signature reads $s=-T$; if nodes $i$ and $j$ establish $k=1$ concordant motif, they establish $T-1$ discordant motifs and the signature reads $s=2-T$; if nodes $i$ and $j$ establish $k=2$ concordant motif, they establish $T-2$ discordant motifs and the signature reads $s=4-T$.
+In words, if nodes $i$ and $j$ establish $k=0$ concordant motifs, they establish $T$ discordant motifs and the signature reads s=Ts=-T; if nodes $i$ and $j$ establish $k=1$ concordant motif, they establish T1T-1 discordant motifs and the signature reads s=2Ts=2-T; if nodes $i$ and $j$ establish $k=2$ concordant motif, they establish T2T-2 discordant motifs and the signature reads s=4Ts=4-T.
 The same line of reasoning can be repeated for all the admissible values of $k$.
-Figure 2: Infographics illustrating our validation procedure.
-Probability distribution of the signature and its Gaussian approximation: while the left panel provides a graphical answer to the question is the empirical value of the signature significantly different from the one expected under the chosen benchmark?, the right panel provides a graphical answer to the question is the deviation negative (hence, the signature is significantly smaller than expected) or positive (hence, the signature is significantly larger than expected)?
-The red area corresponds to the region of validation of the negative links, induced by pairs of series ‘in counterphase’ most of the time; the blue area corresponds to the region of validation of the positive links, induced by pairs of series ‘in phase’ most of the time.
 IV.3 The statistical significance of similarity:
@@
 and ii) if so, is the deviation negative (hence, the signature is significantly smaller than expected) or positive (hence, the signature is significantly larger than expected)?
-The first question can be answered upon calculating the two-sided p-value reading
+The first question can be answered upon calculating the two-sided p-value, reading
 $p_{ij}=2\cdot\min\left\{F(S_{ij}^{*}),1-F(S_{ij}^{*})\right\}$ (54)
@@
 In very general terms, a threshold must be set: if the specific p-value is smaller than it, the associated event is interpreted as statistically significant and the corresponding nodes are connected in the projection.
-[…]
-[…]
-[…]
+One approach is that of adopting the ‘standard’ threshold $t=0.05$ and carrying out each test independently from the others but such a procedure leads to increase [^15] A simple argument can, indeed, be provided.
+The probability that, solely by chance, at least one out of $|H|$ hypotheses is rejected, and the corresponding link validated, amounts to $\text{FWER}=1-(1-t)^{|H|}$: hence, $\text{FWER}\simeq 1$ for (just) $|H|=100$ tests conducted at the significance level of $t=0.05$.
+the number of incorrectly rejected null hypotheses (i.e.
+validated links).
+An alternative approach is that of adopting the Bonferroni correction BonferroniCarloEmilio1936Tsdc; mantegna2011statistically; gualdi2016statistically but such a procedure is known to severely increase [^16] Indeed, the stricter condition $\text{FWER}=0.05$ leads to the threshold $p_{th}=t/|H|$, rapidly vanishing as $|H|$ grows: as a consequence, very sparse (if not empty) projections are obtained.
+the number of incorrectly retained null hypotheses (i.e.
+discarded links).
+In the present paper we apply the so-called False Discovery Rate (FDR) procedure benjamini1995controlling.
+FDR allows one to control for the expected number of ‘false discoveries’ (i.e.
+incorrectly rejected null hypotheses or incorrectly validated links), regardless of the independence of the hypotheses tested.
+Whenever $|H|$ different hypotheses $H_{1},\>H_{2}\dots$, characterised by $|H|$ different p-values, must be tested at the same time, FDR prescribes to first sort the $|H|$ p-values in increasing order
+$\text{p-value}_{1}\leq\dots\leq\text{p-value}_{|H|}$ (57)
 and, then, identify the largest integer $\hat{i}$ satisfying the condition
-[…]
-with $t$ representing the usual single-test significance level, hereby set to the value $t=0.05$: since $H_{i}$ is the hypothesis that the distribution of dyadic motifs established by the $i$-th pair of nodes follows a binomial/Poisson-binomial, rejecting it amounts at connecting the corresponding nodes in the projection.
-Notice that $|H|=N(N-1)/2$.
+$\text{p-value}_{\hat{i}}\leq\frac{\hat{i}t}{|H|}$ (58)
+with $t$ representing the usual single-test significance level, hereby set to the value $t=0.05$: since $H_{i}$ is the hypothesis that the distribution of dyadic motifs established by the $i$-th pair of nodes follows a binomial/Poisson-binomial, rejecting it amounts to connecting the corresponding nodes in the projection.
+Notice that |H|=N(N1)/2|H|=N(N-1)/2.
 Naturally, deciding which test is the best suited one for the problem at hand depends on the importance assigned to false positives and false negatives: as a rule of thumb, the Bonferroni correction can be deemed as appropriate when few tests are expected to be significant (i.e.
@@
 VI Benchmarks for the analysis of multivariate time series
-In order to evaluate the statistical significance of the similarity of any two series $i$ and $j$, a benchmark is needed: a natural choice leads to adopt the class of null models named Exponential Random Graphs (ERGs).
+To evaluate the statistical significance of the similarity between any two series $i$ and $j$, a benchmark is required.
+A natural choice is to adopt the class of null models known as Exponential Random Graphs (ERGs).
 Within such a framework, the constrained maximisation of Shannon entropy
 $S=-\sum_{\mathbf{B}\in\mathbb{B}}P(\mathbf{B})\ln P(\mathbf{B})$ (59)
-leads to assign the generic bipartite network $\mathbf{B}$ the probability
+leads to assign the generic biadjacency matrix $\mathbf{B}$ the probability
 $P(\mathbf{B})=\frac{e^{-H(\bm{\theta},\bm{C}(\mathbf{B}))}}{Z(\bm{\theta})}$ (60)
@@
 The use of linear constraints allows $P(\mathbf{B})$ to be written in a factorised form, i.e.
-as the product of pair-specific probability coefficients Gallo et al.
-(2025).
+as the product of pair-specific probability coefficients gallo2025statistically.
 In order to determine the unknown parameters $\bm{\theta}$, the likelihood maximisation recipe can be adopted: given an observed biadjacency matrix $\mathbf{B}^{*}$, it translates into solving the system of equations
@@
 which prescribes to equate each ensemble average to its observed counterpart.
-Let us, now, generalise such a framework to accommodate models for studying multivariate time series: to this aim, we will follow Marcaccioli and Livan (2020); Gallo et al.
-(2024a) which, in turn, builds upon the analytical approach introduced in Park and Newman (2004) and further developed in Squartini and Garlaschelli (2011).
-Let us, now, make a general observation: as each of the $N$ time series is defined by $T$ time steps, in corresponding of which either the value 1-1- 1 or the value +1+1+ 1 can be assumed, the ensemble is constituted by $|\mathbb{B}|=2^{N\cdot T}$ (bi)adjacency matrices, representing all possible, binary, signed, multivariate time series.
+Let us now generalise such a framework to accommodate models for studying binary multivariate time series.
+To this aim, we will follow Livan2020; gallo2024testing, which, in turn, builds upon the analytical approach introduced in Park2004 and further developed in Squartini2011.
+Let us, now, make a general observation: as each of the $N$ time series is defined by $T$ time steps, in corresponding of which either the value 1-1 or the value +1+1 can be assumed, the ensemble is constituted by all possible $|\mathbb{B}|=2^{N\cdot T}$ binary, signed biadjacency matrices.
 VI.1 Signed Random Graph Model
@@
 $H(\mathbf{B})=\alpha\>B^{+}(\mathbf{B})+\gamma\>B^{-}(\mathbf{B})$ (62)
-i.e. by the two global constraints $B^{+}(\mathbf{B})$ and $B^{-}(\mathbf{B})$. Although being formally analogous to the BiSRGM Gallo et al. (2024a), let us name it bSRGM.
-According to the bSRGM, each entry of a bipartite signed network is a random variable whose behaviour is described by the coefficients
+i.e. by the two global constraints $B^{+}(\mathbf{B})$ and $B^{-}(\mathbf{B})$. Although being formally analogous to the BiSRGM gallo2024testing, we will refer to it as bSRGM.
+According to the bSRGM, each element of a multivariate time series is a random variable whose behaviour is described by the coefficients
 $P(b_{it}=+1)$ $=\frac{e^{-\alpha}}{e^{-\alpha}+e^{-\gamma}}\equiv p^{+},$ (63)
@@
 in other words, $b_{it}$ obeys a generalised Bernoulli distribution whose probability coefficients are determined by the (Lagrange multipliers of the) imposed constraints: each positive link appears with probability $p^{+}$ and each negative link appears with probability $p^{-}$.
-In order to employ the bSRGM for studying real-world multivariate binary time series, the parameters that define it need to be properly tuned: more specifically, one needs to ensure
+In order to employ the bSRGM for studying real-world multivariate time series, the parameters that define it need to be properly tuned: more specifically, one needs to ensure
 $\langle B^{+}\rangle_{\text{bSRGM}}$ $=B^{+}(\mathbf{B}^{*}),$ (65)
@@
 To this aim, one can maximise the likelihood function $\mathcal{L}_{\text{bSRGM}}(\alpha,\gamma)\equiv\ln P_{\text{bSRGM}}(\mathbf{B}^{*}|\alpha,\gamma)$, where $P_{\text{bSRGM}}(\mathbf{B}^{*}|\alpha,\gamma)$ is defined as in eq.
-111, with respect to the unknown parameters that define it Garlaschelli and Loffredo (2008).
+112, with respect to the unknown parameters that define it Garlaschelli2008.
 Such a recipe leads us to find
@@
 by the $2(N+T)$ local constraints $\{k_{i}^{+}(\mathbf{B})\}_{i=1}^{N}$, $\{k_{i}^{-}(\mathbf{B})\}_{i=1}^{N}$, $\{\kappa_{t}^{+}(\mathbf{B})\}_{t=1}^{T}$ and $\{\kappa_{t}^{-}(\mathbf{B})\}_{t=1}^{T}$.
-Although being formally analogous to the BiSCM Gallo et al.
-(2024a), let us name it bSCM.
-According to the bSCM, each entry of a bipartite signed network is a random variable whose behaviour is described by the coefficients
+Although being formally analogous to the BiSCM gallo2024testing, we refer to it as bSCM.
+According to the bSCM, each element of a multivariate time series is a random variable whose behaviour is described by the coefficients
 $P(b_{it}=+1)$ $=\frac{e^{-(\alpha_{i}+\delta_{t})}}{e^{-(\alpha_{i}+\delta_{t})}+e^{-(\gamma_{i}+\eta_{t})}}\equiv p_{it}^{+},$ (70)
 $P(b_{it}=-1)$ $=\frac{e^{-(\gamma_{i}+\eta_{t})}}{e^{-(\alpha_{i}+\delta_{t})}+e^{-(\gamma_{i}+\eta_{t})}}\equiv p_{it}^{-};$ (71)
-in other words, $b_{it}$ obeys a generalised Bernoulli distribution whose probability coefficients are determined by the (Lagrange multipliers of the) imposed constraints: the series $i$, at time $t$ assumes the value +1+1+ 1 with probability $p_{it}^{+}$ and the value 1-1- 1 with probability $p_{it}^{-}$.
-[…]
+in other words, $b_{it}$ obeys a generalised Bernoulli distribution whose probability coefficients are determined by the (Lagrange multipliers of the) imposed constraints: the series $i$, at time $t$ assumes the value +1+1 with probability $p_{it}^{+}$ and the value 1-1 with probability $p_{it}^{-}$.
+Figure 3: Consistency checks regarding the distribution of the signature for subject $\bm{\#100307}$.
+Comparison between the distribution of the dyadic signature induced by the bSRGM/bSCM (left panel/middle panel) and their numerical counterparts, obtained by explicitly sampling $10^{5}$ realisations from the corresponding ensemble: in both cases, the KS test rejects the hypothesis that the two distributions coincide.
+The box-plots summing up the distributions of the KS-scores for the bSRGM and the bSCM, obtained by explicitly sampling $10^{3}$ realisations from the corresponding ensemble (right panel), show that, under both models, the median KS-score amounts to $\text{KS}_{5\%}\simeq 0.965$, the $95\%$ of the values ranging between the $2.5$-th and the $97.5$-th percentiles reading $q_{2.5},q_{97.5}]=[0.961,0.970]$.
+In words, our sampling procedure can be considered satisfactory enough to reproduce the analytical distributions defined by eq.
+44 and eq.
+50, even in case of statistical disagreement.
 In order to tune the parameters defining the bSCM to ensure
@@
 let us maximise the likelihood function $\mathcal{L}_{\text{bSCM}}(\bm{\alpha},\bm{\gamma},\bm{\delta},\bm{\eta})\equiv\ln P_{\text{bSCM}}(\mathbf{B}^{*}|\bm{\alpha},\bm{\gamma},\bm{\delta},\bm{\eta})$, where $P_{\text{bSCM}}(\mathbf{B}^{*}|\bm{\alpha},\bm{\gamma},\bm{\delta},\bm{\eta})$ is defined as in eq.
-A, with respect to the unknown parameters that define it Garlaschelli and Loffredo (2008) (see also Appendix A).
+A, with respect to the unknown parameters that define it Garlaschelli2008 (see also Appendix A).
 Such a recipe leads us to find
@@
 $\kappa_{t}^{-}(\mathbf{B}^{*})$ $=\sum_{i=1}^{N}\frac{e^{-(\gamma_{i}+\eta_{t})}}{e^{-(\alpha_{i}+\delta_{t})}+e^{-(\gamma_{i}+\eta_{t})}}=\langle\kappa_{t}^{-}\rangle,\quad\forall\>t$ (79)
-a system that can be solved only numerically, along the guidelines provided in Vallarano et al. (2021) (see also Appendix B).
+a system that can be solved only numerically, along the guidelines provided in vallarano2021 (see also Appendix B).
+Figure 4: Distributions of the signed, bipartite, and monopartite connectance.
+Left panel: distributions of the positive and negative bipartite connectance, respectively defined as $c^{+}=B^{+}/(N\cdot T)$ and $c^{-}=B^{-}/(N\cdot T)$, across our $100$ subjects.
+Middle and right panel: distributions of the positive and negative monopartite connectance, respectively defined as $\rho^{+}=2L^{+}/N(N-1)$ and $\rho^{-}=2L^{-}/N(N-1)$, across the three different projections of our $100$ subjects.
+Bin widths are computed according to the Freedman-Diaconis rule.
+Averages are plotted as vertical, dashed lines.
+How can a large number of bipartite, negative links co-exist with the large number of positive links characterising the naïve projection?
+Bipartite, negative links are evidently arranged into concordant motifs that, in turn, let the naïve and bSRGM-induced projections be populated by a majority of positive links.
+Employing the bSCM mitigates this situation, as the number of positive and negative links is now rebalanced.
 VII Dataset description
-[…]
-Figure 4: Distributions of the signed, bipartite and monopartite, connectance.
-Left panel: distributions of the positive and negative bipartite connectance, respectively defined as $c^{+}=B^{+}/(N\cdot T)$ (in blue) and $c^{-}=B^{-}/(N\cdot T)$ (in red), across our 100100100 subjects.
-Middle and right panel: distributions of the positive and negative monopartite connectance, respectively defined as $\rho^{+}=2L^{+}/N(N-1)$ (in blue) and $\rho^{-}=2L^{-}/N(N-1)$ (in red), across the three, different projections of our 100100100 subjects.
-Kernel density estimation has been employed for better visualisation; bin widths are computed according to the Freedman-Diaconis rule.
-How can a large number of bipartite, negative links co-exist with the large number of positive links characterising the naïve projection?
-Bipartite, negative links are evidently arranged into concordant motifs that, in turn, induce a naïve projection populated by (a majority of) positive links.
-Notice that the naïve and bSCM-induced distributions overlap, in the right panel.
+We have applied the methodology above to 3T rs-fMRI data from $100$ unrelated subjects of the Young Adult Human Connectome Project (HCP) VanEssen2013.
+Overall, this dataset includes $1200$ participants aged $22$ to $35$ of whom $1087$ underwent at least one rs-fMRI scan.
+Spontaneous, slowly fluctuating brain activity was measured during fMRI, using a Multi-Band Accelerated Echo-Planar-Imaging (EPI) 2D BOLD sequence with whole-brain coverage, while subjects maintained fixation on a central cross and were instructed to lie still and rest quietly - $\text{TR}=720$ ms, $\text{TE}=33.1$ ms, $\text{flip angle}=52$ deg, $\text{FOV}=208\times 180$ $m$ ($\text{RO}\times\text{PE}$), $\text{matrix}=104\times 90$ ($\text{RO}\times\text{PE}$), $\text{slice thickness}=2.0$ $m$, $\text{number of slices}=72$, $\text{resolution}=2.0$ mm isotropic voxels, $\text{multiband factor}=8$, $\text{echo spacing}=0.58$ ms, $\text{BW}=2290$ Hz/Px, $\text{frames per run}=1200$, $\text{run duration\>(min:sec)}=14:33$.
+Within each session, oblique axial acquisitions alternated between phase encoding in a right-to-left (RL) direction in one run and phase encoding in a left-to-right (LR) direction in the other run.
+fMRI recordings underwent a two-step pre-processing procedure.
+First, the data were motion-corrected, intensity-adjusted, and normalised Glasser2013.
+Afterwards, structured noise was corrected using the FIX denoising pipeline Griffanti2014.
+This clean-up included $24$ confound time series derived from the motion estimation (the $6$ rigid-body parameter time series, their backwards-looking temporal derivatives and all $12$ resulting regressors squared Satterthwaite2013).
+On the motion parameters, the temporal high-pass filtering was previously applied.
+They are, then, regressed out of the data aggressively, as they are not expected to contain variance of interest.
+To improve the length of BOLD time series, the right-to-left and left-to-right runs were concatenated after the mean across time removal.
+Finally, each brain was parcellated into $116$ regions of interest (ROIs) over which the signal was averaged.
+More specifically, the cortical regions were parcellated following the Schaefer $100$ atlas Schaefer2018 and the subcortical ones following the Tian $16$ atlas Tian2020.
+As a final result, we obtained a collection of $100$ sets of time series, each one having dimensions $N\times T=116\times 2400$.
 VIII Results
-Let us, now, employ the proposed validation procedure to carry out a ‘binary-to-binary’ analysis, i.e.
-obtain statistically validated binary projections of multivariate binary time series, defined by the ‘rule’ that any two nodes sharing a significantly large number of concordant (discordant) motifs are connected by a positive (negative) edge (see also fig.
-2).
+Let us, now, employ our validation procedure to carry out a ‘binary-to-binary’ analysis, i.e.
+obtain statistically validated binary projections of binary multivariate time series, defined by the ‘rule’ that any two nodes sharing a significantly large number of concordant (discordant) motifs are connected by a positive (negative) edge.
 VIII.1 Consistency checks
-Let us start our analysis by verifying that the ensemble distribution of each, pair-specific signature is consistent with its analytical counterpart, for all benchmarks considered here.
-[…]
+Let us start our analysis by verifying that the ensemble distribution of each pair-specific signature is consistent with its analytical counterpart, for all benchmarks considered here.
+Specifically, let us focus on each subject and test i) if the distribution defined by eq.
+44 is compatible with the ensemble distribution of $S_{ij}$, $\forall\>i<j$ induced by the bSRGM; ii) if the distribution defined by eq.
+50 is compatible with the ensemble distribution of $S_{ij}$, $\forall\>i<j$ induced by the bSCM.
+Each subject can be associated with a KS-score, i.e.
+the percentage of times the Kolmogorov-Smirnov (KS) test (at the level of $5\%$) does not reject the hypothesis of compatibility: while a value of $1$ indicates perfect compatibility, a value of 0 indicates perfect incompatibility.
+The results are summed up by the box-plots depicted in fig.
+3: the median KS-score is $\text{KS}_{5\%}\simeq 0.965$ under both the bSRGM and the bSCM, the $95\%$ of these distributions ranging between the $2.5$-th and the $97.5$-th percentiles reading $q_{2.5},q_{97.5}]=[0.961,0.970]$; in words, our sampling procedure satisfactorily reproduces the analytical distributions defined by eq.
+44 and eq.
+50 even when the KS test rejects the hypothesis of compatibility (e.g.
+for subject #100307\#100307).
+Figure 5: Distributions of CV+, CV- and signed degrees.
+Top panels: box-plots summing up the distributions of the coefficient of variation for positive degrees (CV+, left) and negative degrees (CV-, right) across the three different projections of our $100$ subjects.
+Each box-plot illustrates the distribution of the corresponding metric, indicating its central tendency and dispersion for each model: the positive variant of the CV suggests that the corresponding subgraphs are characterised by a quite homogeneous structure, while the negative one suggests that the opposite holds true; yet, these differences are levelled out when considering bSCM-induced projections.
+Bottom panels: distributions of the positive (left) and negative (right) degrees populating our projections, pooled across our subjects.
+While naïve and bSRGM-induced positive distributions appear as left-skewed, their negative counterparts appear as right-skewed; bSCM-induced distributions, instead, are much flatter in both cases, suggesting a larger heterogeneity of the degrees of this kind.
+The leftmost peak on $k^{+}$ and the rightmost peak on $k^{-}$ are due to subcortical areas.
 VIII.2 Patterns of brain (im)balance at the microscale
@@
 VIII.2.1 Connectance
-When considering the unsigned case, the connectance of the projection returned by the naïve approach is typically large Saracco et al.
-(2017): this holds true here as well, for all the considered subjects (see fig.
-4; for a graphical example, see fig.
-10).
-To be noticed that many validated links are positive, an evidence letting us conclude that the time series associated with the related pairs of areas are concordant for the vast majority of the snapshots constituting the considered time window.
+When considering the unsigned case, the connectance of the projection returned by the naïve approach is typically large saracco2017: this holds true here as well, for all the considered subjects (see fig.
+4).
+It is worth noting that many validated links are positive.
+This leads us to conclude that the time series associated with the related pairs of areas are concordant for the vast majority of the snapshots constituting the considered time window.
 […]
-Figure 5: Distributions of CV+, CV- and signed degrees.
-Top panels: box-plots summing up the distributions of the coefficient of variation for positive degrees (CV+, left) and negative degrees (CV-, right) across the three, different projections of our 100100100 subjects.
-Each box-plot illustrates the distribution of the corresponding metric, indicating its central tendency and dispersion, for each model: the positive variant of the CV suggests that the corresponding subgraphs are characterised by a quite homogeneous structure while the negative one suggests that the opposite holds true, being the corresponding subgraphs characterised by fat-tailed degree distributions.
-Middle and bottom panels: distributions of the positive (left) and negative (right) degrees populating our projections, for subjects 106016 (middle panels) and 857263 (bottom panels).
-Kernel density estimation has been employed for better visualisation; bin widths are computed according to the Freedman-Diaconis rule.
-While positive distributions appear (approximately) bell-shaped around the average, negative distributions are characterised by fat tails, an evidence suggesting the presence of hubs.
-Notice that the naïve and bSCM-induced distributions overlap, in the right panels.
 […]
-The local, bSCM-based filter, instead, accounts for nodes heterogeneity: whenever the positive degree of a node is (much) larger than its negative degree, the distribution of its similarity with any other node moves to the (far) right, hence validating negative links to a much larger extent; the local filter is, thus, stricter with positive than with negative links - and their number is, somehow, re-balanced towards the naïve value.
+The local, bSCM-based filter, instead, accounts for nodes heterogeneity: whenever the positive degree of a node is (much) larger than its negative degree, the distribution of its similarity with any other node moves to the (far) right, hence validating negative links to a much larger extent; the local filter is, thus, stricter with positive than with negative links - and their number is, somehow, re-balanced.
 Figure 6: Distribution of the Jaccard similarity between pairs of projections.
-The numerator of such an index represents the number of positive (negative) negative links occupying the same position in the adjacency matrix of both subjects and its denominator represents the cardinality of the union of the two sets of positive (negative) links populating the adjacency matrices to be compared.
-Kernel density estimation has been employed for better visualisation; bin widths are computed according to the Freedman-Diaconis rule.
-Since a value of 111 indicates perfect overlap while a value of 0 indicates no overlap, positive links yield a larger Jaccard similarity than negative links, a result indicating the former ones as more spatially coherent than the latter ones.
+The numerator of such an index represents the number of positive (negative) links occupying the same position in the adjacency matrix of the two subjects to compare and its denominator represents the cardinality of the union of the two sets of positive (negative) links.
+Bin widths are computed according to the Freedman-Diaconis rule.
+Averages are plotted as vertical, dashed lines.
+Since a value of $1$ indicates perfect overlap while a value of 0 indicates no overlap, positive links yield a larger Jaccard similarity than negative links on both the naïve and bSRGM-induced projections, a result indicating the former ones as more spatially coherent than the latter ones.
+When considering bSCM-induced projections, instead, the spatial coherence of the two kinds of links appears very similar.
 VIII.2.2 Tendency to make hubs
-A related index is the one named tendency to make hubs (TMH), introduced in Saberi et al.
-(2021) and quantifying the tendency of a given degree distribution - when dealing with signed graphs, both the positive and the negative one - to host hubs.
-Hereby, we adopt a slightly different index, rooted in statistical theory and reducing the functional form of a given degree distribution to just one number.
-Named coefficient of variation (CV), it is defined as the ratio between the standard deviation and the expected value of the reference distribution: in symbols,
+A related index is the one named tendency to make hubs (TMH), introduced in Saberi2021 and quantifying the tendency of a given degree distribution - when dealing with signed graphs, both the positive and the negative degree distributions need to be considered - to host hubs.
+Hereby, we adopt a slightly different index, rooted in statistical theory and reducing the functional form of a given degree distribution to just one number: named coefficient of variation (CV), it is defined as the ratio between the standard deviation and the expected value of the reference distribution, i.e.
 $\text{CV}^{+}=\frac{\text{std}[k^{+}]}{\overline{k^{+}}}=\frac{\text{std}[k^{+}]}{2L^{+}/N}$ (80)
@@
 $\text{CV}^{-}=\frac{\text{std}[k^{-}]}{\overline{k^{-}}}=\frac{\text{std}[k^{-}]}{2L^{-}/N};$ (81)
+while a CV smaller than $1$ indicates that the expected value is representative of the entire distribution (as its dispersion around the average is smaller than it), observing a CV larger than 1 indicates that the expected value is not representative of the entire distribution (as its dispersion around the average is larger than it) - from this perspective, a fat-tailed distribution is expected to obey $\text{CV}>1$.
 […]
-As evident from fig.
-5, none of our projections seems to be characterised by the presence of positive hubs; when considering the negative degrees, instead, all projections display a right-skewed distribution that suggests the presence of nodes whose number of connections is (much) larger than that of the others: in a sense, the subgraph induced by the positive links tends to be more homogeneous than the subgraph induced by the negative ones [^17] As highlighted in Luppi et al.
-(2024), ‘[…] whereas cooperative (positive-valued) interactions in the effective connectivity tend to be strong, modular and relatively short-range, competitive interactions are weaker but more long-range, more diffuse and less clustered […]’..
-[…]
-A related result is the one concerning the degree distributions of cortical and subcortical areas: as evident from the figures in Appendix C, subcortical areas tend to have a larger negative degree than cortical areas.
+We also explicitly show the distributions of the positive and negative degrees pooled across our subjects.
+As KS tests (at the level of $5\%$) reveal, practically all pairwise comparisons point out our three (naïve, bSRGM-induced, and bSCM-induced) positive degree distributions as significantly different; moreover, the appearance of positive hubs is enhanced when the global filter is used.
+A related result concerns the degree distributions of cortical and subcortical areas: as evident from the figure in Appendix C, subcortical areas tend to have a larger negative degree than cortical areas.
 VIII.2.3 Spatial coherence of signed connections
-From what we have learned so far, the positive subgraphs are (overall) structurally homogeneous - positive degree distributions are bell-shaped and well-represented by the average - while the negative ones are (overall) structurally heterogeneous - negative degree distributions are fat-tailed, hence over-dispersed around the average.
+From what we have learned so far, the positive subgraphs are (overall) structurally homogeneous, while the negative ones are (overall) structurally heterogeneous.
 Let us, now, ask ourselves how similar positive and negative subgraphs are across subjects.
-Figure 7: Distributions of the Strong and Weak Degree of Frustration.
-Box-plots summing up the distributions of the Strong Degree of Frustration (SDoF) and the Weak Degree of Frustration (WDoF) across the three, different projections of our 100100100 subjects.
-Each box-plot illustrates the distribution of the corresponding metric, indicating its central tendency and dispersion, for each model: in both cases, the bSRGM-induced projections are characterised by the smallest median and largest concentration while the naïve and bSCM-induced projections are peaked in correspondence of largest values but show more variability.
-Since $T^{(---)}\simeq 0$, the SDoF and the WDoF are practically identical and both positive, a result confirming the non null level of frustration characterising our brain networks.
+Figure 7: Distributions of the IBI values across subjects.
+Box-plots summing up the distributions of the values of the strong (left) and weak (right) variants of the Index of Brain Imbalance (IBI) across the three different projections of our $100$ subjects.
+Each box-plot proxies the distribution of the corresponding metric, indicating its central tendency and dispersion: in both cases, the bSRGM-induced projections are characterised by the smallest median, while the naïve projections are peaked in correspondence of larger values, although showing more variability.
+The difference between the ISBI and the IWBI values characterising the bSCM-induced projections is due to the sensitivity of this benchmark to the abundance of the triad ()(---), whose presence is enhanced by the predominantly negative relationships observed within the subcortical and the limbic regions (see also fig.
+8).
+In any case, the positivity of the IBI values confirms the non-null level of frustration characterising humans’ brain networks.
 In order to answer this question, we calculated the Jaccard similarity between pairs of (adjacency matrices representing) individual projections, per benchmark.
@@
 6 shows, the positive variant of such an index (i.e.
-the percentage of positive links located in the same position) ranges between 0.60.60.6 and 111 while its negative variant (i.e.
-the percentage of negative links located in the same position) ranges between 0 and 0.40.40.4: in words, the spatial coherence of positive links is larger than the spatial coherence of negative links, which appear as those (better) characterising single subjects [^19] As highlighted in Luppi et al.
-(2024), ‘[…] Additionally, we find that the model with competitive interactions is not only better able to reflect the empirical FC that it is intended to model but is also better at distinguishing between the FC patterns of different individuals […]’ and ‘[…] Thus, we conclude that the model with competitive interactions is significantly more individual-specific […]’.
-Figure 8: Heatmaps of signed connections within and between brain regions.
-Values are averaged across subjects, for each of our benchmarks.
+the percentage of positive links located in the same position) ranges between $0.6$ and $1$ while its negative variant (i.e.
+the percentage of negative links located in the same position) ranges between 0 and $0.4$, on both the naïve and bSRGM-induced projections: in words, the spatial coherence of positive links is larger than the spatial coherence of negative links, which appear as those (better) characterising single subjects.
+This picture changes when bSCM-induced projections are considered: in such a case, in fact, the spatial coherence of the two kinds of links appears as very similar.
+VIII.2.4 Assessing the brain (im)balance at the microscale
+Let us now inspect the arrangement of negative links within projections.
+To this aim, two choices are possible: i) the first one amounts at considering the index named Strong Degree of Balance (SDoB) gallo2024testing and counting the triads characterised by an even number of negative links, i.e.
+either zero or two - in other words, the ones that are balanced according to the Strong Balance Theory (SBT); ii) the second one amounts at considering the index named Weak Degree of Balance (WDoB) gallo2024testing and counting the triads characterised by zero, two or three negative links - in other words, the ones that are balanced according to the Weak Balance Theory (WBT).
+Figure 8: Heatmaps of the signed connections within and between Yeo’s brain regions.
+Values are averaged across subjects for each of our benchmarks.
 The intensity of each colour is proportional to the magnitude of the corresponding entry.
-Network labels stand for control (CONT), dorsal attention (DA), limbic (LIM), somatomotor (SM), salience/ventral attention (SVA), visual (VIS) networks and default mode network (DMN).
-While the largest number of positive links is found within the DMN, the largest number of negative links is found between the DMN and the DA network, although a large number of the same kind of connections is found between the DMN and the SM, SVA and VIS networks as well - consistently across all benchmarks.
-VIII.2.4 Assessing the brain (im)balance at the microscale
-Let us, now, inspect the way negative links are arranged within projections: to this aim, one may consider the index named Strong Degree of Balance (SDoB) Gallo et al.
-(2024a) and quantifying the percentage of triads characterised by an even number of negative links, i.e.
-either zero or two - in other words, the ones that are balanced according to the Strong Balance Theory (SBT).
-Here, instead, we consider its complementary to 111 and name it [^20] A natural choice would have been that of naming it Strong Degree of Frustration (SDoF); since, however, the triad ()(---)( - - - ) is heavily under-expressed (see also fig.
-7), the strong and weak versions practically coincide: we have, thus, avoided to explicitly distinguish between the two.
-Such an evidence has been also highlighted in Luppi et al.
-(2024), where it is reported that ‘[…] it is less likely that competitively-interacting neighbours of a node will themselves be interacting competitively […]’ Index of Brain Imbalance (IBI): more quantitatively,
-$\text{IBI}=\frac{T^{(+-+)}+T^{(---)}}{T^{(+++)}+T^{(-+-)}+T^{(+-+)}+T^{(---)}}$ (82)
+Network labels stand for subcortical (SUBC), control (CONT), dorsal attention (DA), limbic (LIM), somatomotor (SM), salience/ventral attention (SVA), visual (VIS) networks, and default mode network (DMN).
+The neatest signal is observed for the bSCM-induced projections: while the largest number of positive links is found within the DMN, a large number of negative links is found between the DMN and every other area, as well as between the SUBC and every other area - the largest number of negative connections being located precisely between the DMN and the SUBC.
+Notice that, concerning the blocks lying on the main diagonal, $L^{-}>L^{+}$ (only) for the SUBC and LIM regions (see also fig.
+7).
+The indices above induce two variants of the Index of Brain Imbalance (IBI - practically, their complementary to $1$), i.e. the Index of Strong Brain Imbalance
+ISBI $=\frac{T^{(+-+)}+T^{(---)}}{T^{(+++)}+T^{(-+-)}+T^{(+-+)}+T^{(---)}}$ (82)
+and the Index of Weak Brain Imbalance
+IWBI $=\frac{T^{(+-+)}}{T^{(+++)}+T^{(-+-)}+T^{(+-+)}+T^{(---)}},$ (83)
 with
@@
 $T^{(---)}$ $=\frac{\text{Tr}[(\mathbf{A}^{-})^{3}]}{6};$ (87)
-naturally, the closer the IBI value to 111, the larger the amount of unbalanced triads.
-[…]
-Let us, now, consider the brain parcellation carried out in Thomas Yeo et al.
-(2011), where the authors propose a functional segmentation of brain areas by distinguishing regions (hereby, Yeo’s regions) involved in sensory tasks (e.g.
+naturally, the closer their value to $1$, the larger the amount of unbalanced triads.
+As fig.
+7 shows, the bSRGM-induced projections are the ones characterised by the largest amount of balanced triads - a result that is not surprising once considering the small amount of negative links populating them; the naïve projections, instead, are characterised by a number of balanced triads that amounts at five-to-six times the number of unbalanced triads.
+Lastly, and most interestingly, the behaviour of the bSCM-induced projections is sensitive to the one of the triad ()(---), that can be quite abundant: such a behaviour characterises the subcortical and the limbic regions, partly reconciling our findings with those in Luppi2024181818Speaking of cortical regions, there it is reported that ‘[…] it is less likely that competitively-interacting neighbours of a node will themselves be interacting competitively […]’..
+Overall, however, each median IBI value is positive, a result confirming the non-null level of frustration characterising humans’ brain networks: if, according to Saberi2021, the value $\text{IBI}=0$ is interpreted as the one characterising the energy ground state, the aforementioned results imply that humans’ brain networks are found in some sort of excited state.
+Figure 9: Distributions of the ISBI values associated with triplets of brain areas.
+Cumulative distributions of the ISBI values across the three, different projections of our $100$ subjects: top panels refer to all areas, middle panels refer to all areas without subcortical ones, bottom panels refer to all areas without subcortical and limbic ones.
+The largest ISBI values characterise the triplets of subcortical regions: removing them leads to a different picture as the largest ISBI values, now, characterise the triplets within the LIM region; removing them as well reveals that the largest ISBI values characterise the triplets where the DMN is present once or twice - and, to a lesser extent, those within the DMN and CONT regions.
+The smallest ISBI values, instead, characterise those belonging to the same Yeo’s region (especially the DA, SM and SVA networks).
+Let us, now, consider the brain parcellation carried out in yeo2011organization, where the authors propose a functional segmentation of brain areas by distinguishing regions (hereby, Yeo’s regions) involved in sensory tasks (e.g.
 the visual and the somatomotor networks) from those involved in higher-order, cognitive tasks (e.g.
 the DMN): they are named control (CONT), dorsal attention (DA), limbic (LIM), somatomotor (SM), salience/ventral attention (SVA), visual (VIS) networks and default mode network (DMN).
-While the analysis carried out in Demertzi et al.
-(2022) highlights the presence of negative correlations between the DMN and the so-called task-positive regions, the one carried out in Saberi et al.
-(2022) shows that, in healthy individuals, unbalanced triads predominantly occur between different Yeo’s regions.
-Figure 9: Tensor of the IBI values associated with the triplets of Yeo’s regions.
-Values are averaged across subjects, for each of our benchmarks.
-The size of each point is proportional to the standard deviation of the corresponding value.
-Network labels stand for control (CONT), dorsal attention (DA), limbic (LIM), somatomotor (SM), salience/ventral attention (SVA), visual (VIS) networks and default mode network (DMN).
-While the largest IBI values characterise the triplets where the DMN is present once or, at most, twice, the smallest values characterise those belonging to the same Yeo’s region (especially the SM, SVA and VIS networks) - consistently across all benchmarks.
-In order to verify the first of the aforementioned findings, we have constructed two $\times 7$ 7 matrices: each entry of the first (second) one encodes the number of positive (negative) links between the corresponding Yeo’s regions, aggregated across individuals.
+While the analysis carried out in Demertzi2022 highlights the presence of negative correlations between the DMN and the other regions, the one carried out in Saberi2022 shows that, in healthy individuals, unbalanced triads predominantly occur between different Yeo’s regions.
+In order to verify the first of the aforementioned findings, we have constructed two $8\times 8$ matrices: each entry of the first (second) one encodes the number of positive (negative) links between the corresponding Yeo’s regions - to which the subcortical areas (SUBC) have been added - aggregated across individuals.
 As fig.
-8 shows, i) the largest number of positive links is found within the DMN; ii) the largest number of negative links is found between the DMN and the DA network, although a large number of the same kind of connections is found between the DMN and the SM, SVA and VIS networks as well - consistently across all benchmarks.
-In order to verify the second of the aforementioned findings, we have constructed a $\times 7\times 7$ 7 tensor, each entry of which encodes the IBI associated with a given triplet of Yeo’s regions, averaged across individuals (see fig.
-9).
-Still, as the triplets with a smaller IBI exhibit a larger inter-subject variability, we have also considered the distribution of such values: according to the MW test, i) the triads belonging to the same Yeo’s region are those exhibiting the lowest level of imbalance, the smallest values characterising the triplets involving nodes from the SM, SVA and VIS networks; ii) the triads where the DMN is present once or, at most, twice, exhibit the highest level of imbalance - consistently across all benchmarks.
+8 shows, i) the largest number of positive links is found within the DMN; ii) a large number of negative links is found between the DMN and every other area, as well as between the SUBC and every other area - quite consistently across benchmarks.
+In order to verify the second finding above, we have inspected the distributions of the ISBI values associated with the admissible triplets of areas: as evident from fig.
+9, i) the subcortical triplets are those exhibiting the highest level of imbalance [^19] As highlighted in Saberi2022, ‘Our results suggest that subcortical regions have a prominent role in frustration formation in the brain network at both nodal and connectional levels.
+So they can bring instability and alter properties that facilitate systemic level neural changes, providing adaptive characteristics to the brain’.; ii) the cortical triplets belonging to the LIM region exhibit the highest level of imbalance, followed by those where the DMN is present once or, at most, twice; iii) the cortical triplets belonging to the same Yeo’s region are those exhibiting the lowest level of imbalance (the smallest values characterising the triplets involving nodes from the DA, SM and SVA networks) - with the exception of the LIM region and, to a lesser extent, the DMN and CONT regions.
+Figure 10: Signed communities partitioning the brain of subject $\bm{\#106016}$.
+Top panels: heatmaps illustrating the communities partitioning the projections of subject #106016\#106016, detected by minimising the Bayesian Information Criterion (BIC).
+Bottom panels: heatmaps illustrating the communities partitioning the projections of subject #106016\#106016, detected by minimising the frustration ($F$).
+Although the number of detected communities varies, the overall message conveyed by the two methods can be summed up as follows: the brain seems to be constituted by groups of areas working ‘in phase’ most of the time and ‘in counterphase’ with those of other groups; still, BIC minimisation admits areas from the same group working ‘in counterphase’, as well as areas from different groups working ‘in phase’ - while $F$ minimisation does not.
 VIII.3 Patterns of brain (im)balance at the mesoscale
-Following Gallo et al.
-(2024b), one can adopt an ‘agnostic’ attitude and explore the mesoscale organisation of a projection without aligning with any specific, conceptual framework: a principled approach to achieve such a goal is that of minimising the Bayesian Information Criterion (BIC), i.e.
+Following gallo2024assessing, one can adopt an ‘agnostic’ attitude and explore the mesoscale organisation of a projection without aligning with any specific conceptual framework.
+A principled approach to achieve such a goal is that of minimising the Bayesian Information Criterion (BIC), i.e.
 BIC $=\kappa\ln V-2\ln\mathcal{L};$ (88)
-the first addendum proxies the complexity of a model with the number of its parameters, $\kappa$, the second addendum proxies the accuracy of a model with its log-likelihood, $\ln\mathcal{L}$, and $V=N(N-1)/2$ accounts for the system dimensions.
-Figure 10: Community detection for subject $\bm{106016}$.
-Top panels: heatmaps illustrating the communities partitioning the projections of subject 106016106016106016, detected by minimising the Bayesian Information Criterion (BIC).
-Although the number of detected communities varies from 555 to 888, the overall message conveyed by BIC minimisation does not change: the brain seems to be constituted by groups of areas working either ‘in phase’ or ‘in counterphase’ most of the time, each one being constituted by areas working ‘in phase’ most of the time.
-Bottom panels: heatmaps illustrating the communities partitioning the projections of subject 106016106016106016, detected by minimising the frustration ($F$).
-The number of detected communities varies from 222 to 444 but speaking of modules is, now, a bit imprecise as the $F$-induced partition is dictated by the sole request of minimising the internal number of positive links and the external number of negative links.
+the first addendum proxies the complexity of a model with the number of its parameters, $\kappa$, the second addendum proxies the accuracy of a model with its log-likelihood, $\ln\mathcal{L}$, and V=N(N1)/2V=N(N-1)/2 accounts for the system dimensions.
 Since we aim at describing a projection mesoscale organisation, a natural choice is that of instantiating BIC with the Signed Stochastic Block Model (SSBM), defined by the likelihood function
@@
 $\prod_{r=1}^{k}\prod_{\begin{subarray}{c}s=1\\ s>r\end{subarray}}^{k}(p_{rs}^{+})^{L^{+}_{rs}}(p_{rs}^{-})^{L^{-}_{rs}}(1-p_{rs}^{+}-p_{rs}^{-})^{N_{r}N_{s}-L_{rs}}$ (89)
-and a number of parameters $\kappa_{\text{SSBM}}=k(k+1)$, amounting at twice the number of modules, $k$, into which the projection can be partitioned.
+and a number of parameters $\kappa_{\text{SSBM}}=k(k+1)$, with $k$ indicating the number of modules into which the projection is partitioned.
 Naturally, $N_{r}$ is the number of nodes constituting block $r$, $p_{rr}^{+}$ is induced by the empirical number of positive links within block $r$, i.e.
@@
 $p_{rs}^{-}=\frac{L_{rs}^{-}}{N_{r}N_{s}}.$ (93)
-Figure 11: Community detection for subject $\bm{857263}$.
-Top panels: heatmaps illustrating the communities partitioning the projections of subject 857263857263857263, detected by minimising the Bayesian Information Criterion (BIC).
-Although the number of detected communities varies from 666 to 888, the overall message conveyed by BIC minimisation does not change: the brain seems to be constituted by groups of areas working either ‘in phase’ or ‘in counterphase’ most of the time, each one being constituted by areas working ‘in phase’ most of the time.
-Bottom panels: heatmaps illustrating the communities partitioning the projections of subject 857263857263857263, detected by minimising the frustration ($F$).
-The number of detected communities is at most 222 but speaking of modules is, now, a bit imprecise as the $F$-induced partition is dictated by the sole request of minimising the internal number of positive links and the external number of negative links.
+Figure 11: Heatmaps of the communities partitioning the bSCM-induced projections of four subjects.
+Matrices are coloured according to eq.
+97 and eq.
+98, i.e.
+blocks are blue (red) if the majority of links populating them is positive (negative); white blocks on the diagonal indicate singletons and white blocks off the diagonal indicate that $p_{rs}^{+}=p_{rs}^{-}$.
+Top panels: the communities partitioning the bSCM-induced projections of our subjects are detected by minimising the Bayesian Information Criterion (BIC); all subjects exhibit diagonal blocks with a majority of negative links and off-diagonal blocks with a majority of positive links, thereby satisfying the statistical variant of the Relaxed Balance Theory.
+Bottom panels: the communities partitioning the bSCM-induced projections of our subjects are detected by minimising the frustration ($F$); (left aside the white ones) all subjects admit blocks with a majority of negative links only off the main diagonal and blocks with a majority of positive links only on the main diagonal, hence aligning better with the statistical variant of the Traditional Balance Theory.
 VIII.3.1 A survey on individual projections
-Let us start by considering a pair of individual partitions, stressing that a large link density does not prevent BIC minimisation from detecting statistically significant mesoscopic structures: its sensitivity to both the number and the sign of links, in fact, makes it capable of partitioning both denser and sparser configurations.
+Let us start by considering individual partitions, stressing that a large link density does not prevent BIC minimisation from detecting statistically significant mesoscopic structures: its sensitivity to both the number and the sign of the links, in fact, makes it capable of partitioning both denser and sparser configurations.
 Naïve projections.
-As figs.
-10 and 11 show, BIC minimisation reveals the presence of less than 101010 (more precisely, from 555 to 666) communities: the vast majority of links within communities - if not all, as evident upon looking at some of the modules - is positive, a result implying that the series describing the activity of these areas are concordant for most of the time.
-For what concerns the links between communities, instead, the picture becomes much more interesting: the third and fourth community in fig.
-10 are, in fact, connected by mostly negative links, an evidence suggesting that the two blocks of areas are ‘in counterphase’ for most of the time; while this holds true for the first and third community as well, the first and fourth community are connected by mostly positive links, i.e.
+As fig.
+10 shows, BIC minimisation reveals the presence of $6$ communities: the vast majority of links within communities - if not all, as evident upon looking at some of the modules - is positive, a result implying that the series describing the activity of these areas are concordant most of the time.
+For what concerns the links between communities, instead, the picture becomes much more interesting: the fourth and fifth communities in fig.
+10 are, in fact, connected by mostly negative links, an evidence suggesting that the two blocks of areas are ‘in counterphase’ most of the time; while this holds true for the fifth and sixth communities as well, the first and second communities are connected by mostly positive links, i.e.
 even if these blocks of areas are separated, they still seem to work ‘in phase’.
 Validated projections.
-Filtering our projections with a homogeneous benchmark makes the picture even cleaner, as each block is, now, constituted by links of practically the same sign: more precisely, while positive links constitute the majority of links, negative links are found only between modules.
-Filtering our projections with a heterogeneous benchmark, instead, returns a picture which is halfway between the naïve one and the bSRGM-induced one, as the blocks are sparser than the naïve ones but less homogeneous than the ones induced by the bSRGM.
-The overall message conveyed by such an analysis, however, does not change: the brain seems to be constituted by groups of areas working either ‘in phase’ or ‘in counterphase’ most of the time, each one being constituted by areas working ‘in phase’ most of the time.
-Figure 12: Heatmaps of the communities partitioning the bSCM-induced projections of four subjects.
-The communities partitioning the bSCM-induced projections of subjects 106016106016106016, 857263857263857263, 115320115320115320 and 654754654754654754 are detected by minimising the Bayesian Information Criterion (BIC).
-Matrices are coloured according to eq.
-96 and eq.
-97, i.e.
-blocks are blue (red) if the majority of links populating them is positive (negative); white blocks on the diagonal indicate singletons - specifically, a cortical region for subject 106016106016106016 and a subcortical region for subject 654754654754654754.
-All subjects satisfy the statistical variant of the Relaxed Balance Theory, being characterised by blocks with a majority of positive links both on and off the main diagonal - subjects 115320115320115320 and 654754654754654754 are even characterised by one, diagonal block (i.e.
-a community) populated by a majority of negative links.
+Filtering our projections with a homogeneous benchmark makes the picture even clearer, as each block is now constituted by links of practically the same sign: more precisely, while positive links constitute the majority of links, (few) negative links are found only between modules.
+Filtering our projections with a heterogeneous benchmark, instead, returns a picture that is halfway between the naïve one and the bSRGM-induced one, as the blocks are sparser than the naïve ones, less homogeneous than those induced by the bSRGM, but richer in negative connections than both.
+The overall message conveyed by such an analysis, however, does not change: the brain seems to be constituted by groups of areas working ‘in phase’ most of the time and ‘in counterphase’ with those of other groups; still, BIC minimisation admits areas working ‘in counterphase’ even if belonging to the same group (see, for example, the patterns highlighted in figs.
+7 and 8), as well as areas working ‘in phase’ even if belonging to different groups.
 VIII.3.2 Traditionally or relaxedly balanced?
 Motivated by the last observation, we now employ our benchmarks to probe the patterns of structural (im)balance at the mesoscopic level.
-We also explicitly notice that speaking of balance theory for brain networks may appear a bit improper: still, this is the only existing framework to interpret our results.
+We also explicitly acknowledge that discussing balance theory for brain networks may seem somewhat inappropriate; nonetheless, this is the only existing framework to interpret our results.
 Traditional theory of balance.
-According to the strong variant of the traditional theory of balance (TBT), the optimal partition of the set of nodes of a network consists of two groups, whose internal (external) connections are dominated by positive (negative) signs; by contrast, the weak variant of the TBT allows for any number of groups.
-More formally, both versions of the TBT can be probed by minimising the frustration, defined as
+According to the strong variant of the Traditional Balance Theory (TBT), the optimal partition of a network set of nodes consists of two groups, whose internal (external) connections are positive (negative); by contrast, the weak variant of the TBT allows for any number of groups.
+More formally, both versions of the TBT can be probed by finding the partition $\bm{\sigma}$ that minimises the frustration, defined as
 $F(\bm{\sigma})=L_{\bullet}^{-}+L_{\circ}^{+},$ (94)
-i.e. as the number of negative links within modules (indicated with a filled dot) plus the number of positive links between modules (indicated with an empty dot).
-Figure 13: Basic statistics of the partitions across subjects.
-Top panels: histograms of the number of communities detected on the projections returned by each benchmark and optimisation criterion.
-While $F$ avoids creating frustrated patterns, by individuating a smaller number of larger communities, BIC singles out a larger number of smaller communities.
-Middle panels: histograms of the size of communities detected on the projections returned by each benchmark and optimisation criterion.
-These panels refine the picture provided by the previous ones, clarifying that $F$ tends to separate few singletons from the set of remaining nodes.
-BIC tends to individuate singletons as well, although the size of communities into which the set of remaining nodes is partitioned varies over a broader range (approximately, from 111 to 404040).
-Bottom panels: number of times each brain region is individuated as a singleton.
-As evident from the plot, this is mainly observed for subcortical regions, acting as ‘integrators’ of cortical ones.
-More quantitatively, the fraction of singletons that are identifiable as subcortical regions amounts at i) 0.9$\simeq$ 0.9≃ 0.9 according to $F$ and 0.9$\simeq$ 0.9≃ 0.9 according to BIC, on the naïve projections; ii) 0.9$\simeq$ 0.9≃ 0.9 according to $F$ and 0.7$\simeq$ 0.7≃ 0.7 according to BIC, on the bSRGM-induced projections; iii) 0.8$\simeq$ 0.8≃ 0.8 according to $F$ and 0.5$\simeq$ 0.5≃ 0.5 according to BIC, on the bSCM-induced projections.
-As figs.
-10 and 11 show, minimising $F$ always leads to a smaller number of communities with respect to the one revealed by minimising BIC: in fact, from 222 to 444, depending on the projection.
-The naïve and bSCM-induced ones are the most similar, as 222 mostly positive blocks (one being much larger than the others), separated by a vast majority of negative links, are individuated.
-The naïve, $F$-induced partition also illustrates the way the optimisation of $F$ works: the large number of positive links leads $F$ to collect them together, being more convenient to accommodate a small number of negative links within a larger, mostly positive, block than letting a large number of positive links be placed between blocks.
+i.e.
+as the number of negative links within modules (indicated with a filled dot) plus the number of positive links between modules (indicated with an empty dot).
+Let us also remark that minimising $F$ leads to results that are very similar to the ones obtained upon maximising the signed modularity gallo2024assessing.
+The naïve, $F$-induced partition in fig.
+10 illustrates a first way the optimisation of $F$ works: a large number of positive links leads $F$ to collect them together, being more convenient to accommodate a small number of negative links within a larger, mostly positive, block than letting a large number of positive links be placed between blocks.
+The heterogeneous, $F$-induced partition in fig.
+10, instead, illustrates a second way the optimisation of $F$ works: a large number of negative links leads $F$ to split them, even at the price of creating many singletons.
 In any case, the minimisation of $F$ further confirms that none of our projections obeys the traditional formulation of the Balance Theory.
@@
 Such a result motivates us to inspect the theory of balance best accommodating our observations.
-To answer such a question, let us, first, notice that the TBT can be formalised upon writing
+To answer such a question, let us first notice that the TBT can be formalised upon writing
 $p_{rr}^{-}=0,\quad\forall\>r$ (95)
@@
 with $p_{rr}^{+}$ indicating the probability that any two nodes belonging to the same block $r$ are connected by a positive link, $p_{rs}^{+}$ indicating the probability that any two nodes belonging to the different blocks $r$ and $s$ are connected by a positive link and analogously for their negative counterparts.
-As noticed in Gallo et al.
-(2024b), however, the positions defining the TBT can be replaced with the milder ones reading
+As noticed in gallo2024assessing, however, the positions defining the TBT can be replaced with the milder ones reading
 $\text{sgn}[p_{rr}^{+}-p_{rr}^{-}]=+1,\quad\forall\>r$ (97)
-which amounts at requiring $p_{rr}^{+}>p_{rr}^{-}$, $\forall\>r$ and
+which amounts to requiring $p_{rr}^{+}>p_{rr}^{-}$, $\forall\>r$ and
 $\text{sgn}[p_{rs}^{+}-p_{rs}^{-}]=-1,\quad\forall\>r<s$ (98)
-which amounts at requiring $p_{rs}^{+}<p_{rs}^{-}$, $\forall\>r<s$: in words, the deterministic rules firstly defined by Cartwright, Harary and Davis are replaced by a set of probabilistic criteria individuating ‘a tendency’ to obey, or not to obey, the TBT.
+which amounts to requiring $p_{rs}^{+}<p_{rs}^{-}$, $\forall\>r<s$: in words, the deterministic rules firstly defined by Cartwright, Harary and Davis are replaced by a set of probabilistic criteria individuating ‘a tendency’ to obey, or not to obey, the TBT.
 More formally, a configuration satisfying eqs.
-96 and 97 will be claimed to support the statistical variant of the TBT - specifically, its strong variant if $k=2$ and its weak variant if $k>2$; if, instead, $p_{rr}^{+}\leq p_{rr}^{-}$ for some diagonal blocks or $p_{rs}^{+}\geq p_{rs}^{-}$ for some off-diagonal blocks, it will be claimed to support the statistical variant of the relaxed balance theory (RBT).
-Let us, again, take a look at figs.
-10 and 11: although $p_{rs}^{+}\geq p_{rs}^{-}$ for all diagonal blocks (i.e.
-with $r=s$), $p_{rs}^{+}\geq p_{rs}^{-}$ for several, off-diagonal blocks (i.e.
-with $r\neq s$) as well: we are, thus, led to the conclusion that all projections obey the statistical variant of the RBT Gallo et al.
-(2024b) as a non-negligible number of positive links is found between - and not only within - clusters.
-To evaluate the alignment of our projections with the statistical variant of the TBT, let us count the number of subjects obeying eqs.
-96 and 97.
-When considering the BIC-induced partitions, we find that no subject obeys the statistical variant of the TBT, irrespectively from the benchmark employed: while a vast majority of diagonal blocks is, in fact, populated by a number of positive links that is larger than the number of negative ones - although subjects admitting diagonal blocks populated by a majority of negative links exist (see also fig.
-12) - also several, off-diagonal blocks are populated by a number of positive links that is larger than the number of negative ones - as the two examples in figs.
-10 and 11 clearly show, $p_{rs}^{+}\geq p_{rs}^{-}$ for many of them.
-When considering the $F$-induced partitions, instead, 65 naïve projections, 47 bSRGM-induced projections and 69 bSCM-induced projections obey the statistical variant of the TBT.
-Figure 14: Distributions of BIC and $F$ values across the partitions induced by their minimisation.
-Top panels: histograms of the BIC values characterising the partitions determined by minimising it and Yeo’s partitions.
-As expected, the former are characterised by a slightly smaller value of such a score than Yeo’s partitions.
-Middle panels: histograms of the $F$ values characterising the partitions determined by minimising it and Yeo’s partitions.
-In this second case, the former are characterised by a much smaller value of such a score than Yeo’s partitions.
-Bottom panels: histograms of the $F$ values characterising the partitions determined by minimising BIC and Yeo’s partitions.
-In this third case, both are characterised by a quite large value of $F$.
-Taken together, our results indicate that brain networks do not seek to minimise frustration although BIC minimisation seems (to embody a principle that is) capable of capturing a structure quite similar to the one individuated by Yeo’s task-positive regions.
-Kernel density estimation has been employed for better visualisation; bin widths are computed according to the Freedman-Diaconis rule.
+97 and 98 will be claimed to support the statistical variant of the TBT - specifically, its strong variant if $k=2$ and its weak variant if $k>2$; if, instead, $p_{rr}^{+}\leq p_{rr}^{-}$ for some diagonal blocks or $p_{rs}^{+}\geq p_{rs}^{-}$ for some off-diagonal blocks, it will be claimed to support the statistical variant of the Relaxed Balance Theory (RBT).
+In order to evaluate the alignment of our projections with either statistical variants, let us take a look at the top panels of fig.
+10, depicting the BIC-induced partitions: since $p_{rs}^{+}\leq p_{rs}^{-}$ for several, diagonal blocks (i.e.
+with $r=s$) and $p_{rs}^{+}\geq p_{rs}^{-}$ for several, off-diagonal blocks (i.e.
+with $r\neq s$), we are led to the conclusion that all projections obey the statistical variant of the RBT gallo2024assessing as a non-negligible number of negative (positive) links is found within (between) - and not only between (within) - clusters.
+As fig.
+11 further confirms, no subject obeys the statistical variant of the TBT, regardless of the benchmark employed, when considering the BIC-induced partitions.
+Let us, now, take a look at the bottom panels of fig.
+10, depicting the $F$-induced partitions: although eqs.
+97 and 98 are not strictly obeyed, our subject, now, admits blocks with a majority of negative links only off the main diagonal and blocks with a majority of positive links only on the main diagonal; we are, thus, led to the conclusion that all projections align better with the statistical variant of the TBT.
+As fig.
+11 further confirms, all subjects align better with the statistical variant of the TBT, regardless of the benchmark employed, when considering the $F$-induced partitions.
 VIII.4 Towards a representative brain?
-Let us, now, move from individual analyses to the aggregated one, the aim being that of understanding the extent to which the previous results can be generalised.
+Let us now move from individual analyses to the aggregated one, the aim being to understand the extent to which the previous results can be generalised.
+Figure 12: Basic statistics of the partitions across subjects.
+Top panels: histograms of the number of communities detected on the projections returned by each benchmark and optimisation criterion.
+For what concerns naïve and bSRGM-induced projections, $F$ avoids creating frustrated patterns by individuating a smaller number of larger communities, while BIC singles out a larger number of smaller communities; for what concerns bSCM-induced projections, instead, $F$ individuates a larger number of smaller communities than BIC.
+Middle panels: histograms of the size of communities detected on the projections returned by each benchmark and optimisation criterion.
+These panels refine the picture provided by the previous ones, clarifying that $F$ tends to separate the singletons from the set of remaining nodes; BIC tends to individuate singletons as well, although the size of communities into which the set of remaining nodes is partitioned varies over a broader range.
+Bottom panels: number of times each brain region is individuated as a singleton.
+As evident from the plot, this is almost solely observed for subcortical regions, especially on $F$-induced partitions; the most ‘integrated’ picture is, instead, returned by the bSCM-induced projections, partitioned by minimising BIC.
 VIII.4.1 Distribution of the number of communities
@@
 The results are shown in fig.
-13: the number of communities detected by minimising the frustration is practically always smaller than the number of communities detected by minimising BIC - the $F$-induced mode being 222 and the BIC-induced mode varying from 777 to 888; moreover, the $F$-induced distributions are narrower than the BIC-induced ones, a result hinting at a more pronounced inter-subject variability, in the second case.
+12: for what concerns naïve and bSRGM-induced projections, the number of communities detected by minimising the frustration is practically always smaller than the number of communities detected by minimising BIC - the $F$-induced mode being $2$ and the BIC-induced mode varying from $7$ to $8$; moreover, the $F$-induced distributions are narrower than the BIC-induced ones, a result hinting at a more pronounced inter-subject variability, in the second case.
+Concerning the bSCM-induced projections, instead, $F$ identifies a larger and more dispersed number of smaller communities than BIC.
 VIII.4.2 The role of singletons
-Figure 13 illustrates the histograms of the number of communities detected on the projections returned by each benchmark and optimisation criterion, revealing that $F$ individuates a smaller number of larger communities while BIC singles out a larger number of smaller communities.
-The histograms of the size of communities detected on the same projections refine the aforementioned picture, clarifying that $F$ tends to individuate (overall, too) simple structures such as few singletons detached from the rest of the nodes; although BIC individuates singletons as well, the size of communities into which the rest of the nodes is partitioned ranges from 111 to 404040.
+The histograms of the size of communities detected on the same projections refine the aforementioned picture, clarifying that $F$ tends to individuate simpler structures than BIC, such as fewer modules accompanied by several singletons detached from the rest of the nodes - although BIC may individuate singletons as well, the size of communities into which the rest of the nodes is partitioned varies over a broader range.
+Figure 13: Distributions of the BIC and $F$ values across the partitions induced by their minimisation.
+Top panels: histograms of the BIC values characterising the partitions determined by minimising it (in blue) and Yeo’s partitions (in orange).
+As expected, the former are characterised by a slightly smaller value of such a score than Yeo’s partitions.
+Middle panels: histograms of the $F$ values characterising the partitions determined by minimising it (in blue) and Yeo’s partitions (in orange).
+In this second case, the former are characterised by a (much) smaller value of such a score than Yeo’s partitions.
+Bottom panels: histograms of the $F$ values characterising the partitions determined by minimising BIC (in blue) and Yeo’s partitions (in orange).
+In this third case, both are characterised by a quite large value of $F$.
+Bin widths are computed according to the Freedman-Diaconis rule.
+Taken together, our results indicate that i) brain networks do not seek to minimise frustration - even when removing the subcortical areas; ii) BIC minimisation embodies a principle returning a structure that captures (at least) part of the features of the Yeo’s task-based one.
 Interestingly, singletons often coincide with the so-called subcortical regions, i.e.
-areas serving as hubs ‘integrating’ the cortical regions, among the other things: more quantitatively, the fraction of subcortical singletons amounts at i) 0.9$\simeq$ 0.9≃ 0.9 according to $F$ and 0.9$\simeq$ 0.9≃ 0.9 according to BIC, on the naïve projections; ii) 0.9$\simeq$ 0.9≃ 0.9 according to $F$ and 0.7$\simeq$ 0.7≃ 0.7 according to BIC, on the bSRGM-induced projections; iii) 0.8$\simeq$ 0.8≃ 0.8 according to $F$ and 0.5$\simeq$ 0.5≃ 0.5 according to BIC, on the bSCM-induced projections.
+areas serving as hubs ‘integrating’ the cortical regions, among the other things: more quantitatively, the fraction of subcortical singletons amounts to i) $\simeq 0.86$ according to $F$ and $\simeq 0.79$ according to BIC, on the naïve projections; ii) $\simeq 0.87$ according to $F$ and $\simeq 0.59$ according to BIC, on the bSRGM-induced projections; iii) $\simeq 0.74$ according to $F$ and $\simeq 0.03$ according to BIC, on the bSCM-induced projections.
+In words, carrying out a BIC-based community detection on bSCM-induced projections rarely leads to isolate subcortical areas, a result indicating that accounting for local properties such as the series- and the time-specific degrees leads to recover projections where all regions are (practically always) integrated.
 VIII.4.3 Distribution of BIC and $F$ values
-A related analysis is the one concerning the distributions of the BIC and $F$ values across the partitions determined by minimising each of them.
+A related analysis concerns the distributions of the BIC and $F$ values across the partitions determined by minimising each of them.
 As fig.
-14 confirms, the partitions retrieved by minimising BIC are characterised by a smaller value of such a score than Yeo’s partitions: while this is expected, it should be also noticed that these differences are not very large.
-Besides, both kinds of partitions are characterised by a similar, and quite large, value of $F$.
+13 confirms, the partitions retrieved by minimising BIC are characterised by a smaller value of such a score than Yeo’s partitions; besides, both kinds of partitions are characterised by a similar, and quite large, value of $F$.
 Inspecting the distribution of the values of $F$, instead, reveals that the partitions retrieved by minimising it are much less frustrated than Yeo’s partitions.
-Taken together, our results suggest that while brain networks do not seek to minimise frustration, BIC minimisation represents a recipe that is capable of capturing a structure de facto very close to the one captured by Yeo’s task-positive regions.
-VIII.4.4 The barycentrical partition
-Let us, now, ask ourselves if a criterion to condensate the information embodied by each, individual partition can be devised; naturally, the aforementioned results let us opt for solely focusing on the BIC-induced ones.
-A simple recipe could be that of averaging the individual projections in a benchmark-specific fashion; still, the presence of a vast majority of positive links and the spatial incoherence of the negative ones would cause most of the latter to be deleted.
-Alternatively, one could try to individuate the ‘most representative’ brain by individuating the ‘most representative’ mesoscale structure: finding an answer to this question ultimately amounts at finding a method to compare any two partitions, say $X\equiv\{x_{1}\dots x_{n}\}$ and $Y\equiv\{y_{1}\dots y_{m}\}$.
-Rand index (RI). One of the simplest metric is the RI, reading
-$\text{RI}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{FP}+\text{TN}+\text{FN}}=\frac{\text{TP}+\text{TN}}{N(N-1)/2}$ (98)
-[…]
-[…]
-Figure 15: Brain centroids individuated by our measures of distance between partitions.
-Top panels: brain centroids individuated by adopting the Rand Index (RI), the Normalised Mutual Information (NMI), the Rajski distance (RD), the Wallace Index (WI), the Jaccard Index (JI) and the Variation of Information (VI).
-Generally speaking, while RI, NMI and RD individuate projections with a number of communities lying on the far-right side of the spectrum (i.e.
-from 999 to 111111), WI, JI and VI individuate projections with a number of communities lying on the far-left side of the spectrum (i.e.
-444).
+Taken together, our results suggest that i) brain networks do not seek to minimise frustration (even when removing subcortical areas - as in this case, to carry out a fair comparison with Yeo’s classification); ii) BIC minimisation is capable of capturing a structure that shares several similarities with the one defined by Yeo’s task-based regions.
+Figure 14: Brain centroids individuated by the considered distances between partitions.
+Top panels: brain centroids individuated by adopting the Rand Index (RI), the Normalised Mutual Information (NMI), the Rajski distance (RD), the Wallace Index (WI), the Jaccard Index (JI), and the Variation of Information (VI).
+Generally speaking, while RI, NMI, and RD individuate projections with a number of predominantly positive communities lying on the right side of the spectrum (i.e.
+from $7$ to $10$), WI, JI, and VI individuate projections with a number of predominantly positive communities lying on the left side of the spectrum (e.g.
+55).
 Such a result may be explained by noticing that indices like WI and JI are solely defined in terms of true positives: as a consequence, identifying a smaller number of larger communities may enhance the detection of such events.
 Indices like RI, on the other hand, account for true negatives as well, whence the larger number of smaller communities.
-Bottom panels: different metrics induces a different overlap between our structure-based partition and Yeo’sfunction-based one: while the larger communities individuated by WI, JI and VI tend to merge large portions of different Yeo’s areas, the smaller ones individuated by RI, NMI and RD tend to be populated by (even a large portion of) single Yeo’s areas.
+Bottom panels: different metrics induce a different overlap between our structure-based partitions and Yeo’s task-based one.
+Overall, the communities identified by WI, JI, VI, and RD are able to capture up to $75\%$ of certain Yeo’s regions, while the smaller ones identified by RI and NMI rarely capture more than $50\%$ of a given Yeo’s region.
+VIII.4.4 The barycentric partition
+Let us, now, ask ourselves if a criterion to condense the information embodied by each individual partition can be devised.
+To this aim, one could try to single out the ‘most representative’ brain by individuating the ‘most representative’ mesoscale structure: finding an answer to this question ultimately amounts to finding a method to compare any two partitions, say $X\equiv\{x_{1}\dots x_{n}\}$ and $Y\equiv\{y_{1}\dots y_{m}\}$.
+Rand index (RI). One of the simplest metrics is the RI, reading
+$\text{RI}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{FP}+\text{TN}+\text{FN}}=\frac{\text{TP}+\text{TN}}{N(N-1)/2}$ (99)
+and computing the ratio between the sum of true positives (TP; the number of nodes that are ‘together’ in both partitions, i.e.
+in the same subset in $X$ and in the same subset in $Y$) and true negatives (TN; the number of nodes that are ‘separated’ in both partitions, i.e.
+in different subsets in $X$ and in different subsets in $Y$) and the total number of pairs of nodes.
+Naturally, $0\leq\text{RI}\leq 1$.
+Other variants are the Wallace Index (WI), defined as $\text{WI}=\text{TP}/(\text{TP}+\text{FP})$, and the Jaccard Index (JI), defined as $\text{JI}=\text{TP}/(\text{TP}+\text{FP}+\text{FN})$: notice, however, that both WI and JI measure the agreement between two partitions by solely considering the number of TP.
+Figure 15: Average brain, partitioned according to BIC minimisation and $F$ minimisation.
+As the presence of a vast majority of positive links causes most of the negative ones populating the naïve and bSRGM-induced projections to be deleted, only the bSCM-induced projections have been considered: while the average brain partitioned according to $F$ minimisation is characterised by $2$ predominantly positive communities that gather the vast majority of Yeo’s regions together and 66 predominantly negative modules that fragment the remaining ones, the average brain partitioned according to BIC minimisation is characterised by modules sharing similarities with the ones induced by NMI and those induced by VI, and capable to capture up to the $80\%$ of certain Yeo’s regions.
 Normalised mutual information (NMI). A second metric is the NMI, reading
@@
 $\text{MI}(X,Y)=\sum_{i=1}^{N}\sum_{j=1}^{N}h_{ij}\ln\left(\frac{h_{ij}}{f_{i}g_{j}}\right),$ (101)
-[…]
+with $h_{ij}=|X_{i}\cap Y_{j}|/N$ being the fraction of nodes common to subsets $X_{i}$ and $Y_{j}$, $S(X)=-\sum_{i=1}^{N}f_{i}\ln f_{i}$ being the Shannon entropy induced by the partition $X$ and $f_{i}=|X_{i}|/N$ being the fraction of nodes within the cluster $X_{i}$, $S(Y)=-\sum_{j=1}^{N}g_{j}\ln g_{j}$ being the Shannon entropy induced by the partition $Y$ and $g_{j}=|Y_{j}|/N$ being the fraction of nodes within the cluster $Y_{j}$.
+Naturally, $0\leq\text{NMI}\leq 1$.
 Rajski distance (RD). A third metric is the variation of information (VI), reading
@@
 $\text{RD}(X,Y)=\frac{\text{VI}(X,Y)}{S(X,Y)}=1-\frac{\text{MI}(X,Y)}{S(X,Y)};$ (104)
-[…]
+naturally, $0\leq\text{RD}\leq 1$.
 Finding the barycentre.
-Any of the metrics above can be employed to assess how different the partitions characterising the brains of any two subjects $u$ and $u^{\prime}$ are.
-Once such a matrix of distances, say $\mathbf{D}\equiv\{D_{uu^{\prime}}\}_{u,u^{\prime}=1}^{U}$, with $U=100$, has been obtained, a second method is needed to condensate such a wealth of information.
+Any of the metrics above can be employed to assess how different the partitions characterising the brains of any two subjects, say $u$ and $u^{\prime}$, are.
+Once such a matrix of distances, say $\mathbf{D}\equiv\{D_{uu^{\prime}}\}_{u,u^{\prime}=1}^{U}$, with $U=100$, has been obtained, a second method is needed to condense such a wealth of information.
 Our proposal is that of calculating the vector $\underline{\mathbf{D}}\equiv\{\underline{D}_{u}\}$, where
@@
 at the minimum, average one.
-Such a partition will be refereed to as to the barycentrical partition or brain centroid.
+Such a partition will be referred to as the barycentric partition or brain centroid.
 Let us explicitly notice that while VI and RD are proper distances over the space of possible partitions of a system, hence inducing the positions $D_{uu^{\prime}}^{\text{VI}}=\text{VI}_{uu^{\prime}}$ and $D_{uu^{\prime}}^{\text{RD}}=\text{RD}_{uu^{\prime}}$, RI and NMI quantify the similarity of any two partitions [^20] In fact, both RI and NMI could be directly maximised., hence inducing the positions $D_{uu^{\prime}}^{\text{RI}}=1-\text{RI}_{uu^{\prime}}$ and $D_{uu^{\prime}}^{\text{NMI}}=1-\text{NMI}_{uu^{\prime}}$.
-The observations about the importance of considering filtered projections and the prominent role played by heterogeneous benchmarks motivate us to focus on the partitions induced by the bSCM.
+The observations about the importance of considering filtered projections and the prominent role played by heterogeneous benchmarks motivate us to focus on the partitions obtained by minimising BIC on the projections induced by the bSCM.
 Upon looking at fig.
-15, we realise that i) RI individuates a barycentrical projection with 111111 communities; ii) NMI and RD individuate the same projection, in turn very close to the one individuated by RI - the only difference being the number of communities, now reading 999; iii) WI, JI and VI individuate a barycentrical projection with 444, major communities (plus a number of singletons).
-To sum up, while RI, NMI and RD individuate projections with a larger number of communities, WI, JI and VI individuate projections with a smaller number of communities.
+14, we realise that i) RI individuates a barycentric projection with $10$ predominantly positive communities; ii) NMI individuates a barycentric projection with $8$ predominantly positive communities; iii) RD individuates a barycentric projection with $7$ predominantly positive communities; iv) WI, JI, and VI individuate a barycentric projection with $5$, major communities defined by predominantly positive connections.
 The explanation of such a result may lie in the evidence that indices like WI and JI are solely defined in terms of true positives: as a consequence, identifying a smaller number of larger communities may enhance the detection of such events - although negative connections become under-represented; indices like RI, on the other hand, account for true negatives as well, whence the larger number of smaller communities - connected by a vast majority of negative links.
-As a last comment, let us notice that RI, NMI and RD retrieve partitions with a larger BIC than those retrieved by WI, JI and VI.
-Let us, now, compare the community structure of the different brain centroids with the partition defined by the Yeo’s regions introduced before.
+Let us, now, compare the community structure of the different brain centroids with the partition defined by Yeo.
 As fig.
-15 shows, each metrics induces a different overlap between our structure-based partition and Yeo’s function-based one: still, while the larger communities individuated by WI, JI and VI tend to merge (even) large portions of different Yeo’s areas, some of the smaller ones individuated by RI, NMI and RD tend to be populated by (even a large portion of) single Yeo’s areas.
-Moreover, WI, JI and VI tend to return singletons, represented by practically empty rows.
+14 shows, each metrics induces a different overlap between our structure-based partitions and Yeo’s task-based one: while the communities individuated by WI, JI, VI, and RD tend to capture (even) large portions of certain Yeo’s communities - up to the $75\%$ for the first three indices and up to the $60\%$ for the fourth one - the smaller ones individuated by RI and NMI rarely capture more than half of a given Yeo’s community.
+A simpler recipe to identify the representative brain could be that of averaging the individual projections in a benchmark-specific fashion; as the presence of a vast majority of positive links causes most of the negative ones populating the naïve and bSRGM-induced projections to be deleted, here we limit ourselves to consider the bSCM-induced projections.
+As fig.
+15 shows, both the BIC minimisation and the $F$ minimisation lead to single out $8$ communities: in the second case, these modules appear as quite uninformative as $6$ are very small groups of nodes fragmenting a single Yeo’s region in a quite unrealistic fashion, while the seventh and the eighth ones gather all the remaining regions together; in the first case, instead, these modules share similarities with the ones induced by NMI and those induced by VI, and are capable to capture up to the $80\%$ of certain Yeo’s regions.
 IX Discussion
-The study of the role played by negative connections within and between brain areas is a topic that has recently gained attention within the neuroscientific literature.
-The method that we have proposed here produces connections connotated by a negative sign as the result of a validation procedure that counts the number of times the corresponding series are found to be discordant [^22] Aggregating positive and negative contributions can, in fact, obscure the presence of anti-correlations.
-For instance, while the DMN and the DA network are connected by, approximately, $3\cdot 10^{5}$ positive links, they are also connected by, roughly, $9\cdot 10^{4}$ negative links: summing these numbers together would yield a net positive count, misleadingly hiding the presence of anti-correlated activity.
-and compare it with the prediction of a properly defined benchmark.
-Such a validation scheme allows us to answer a number of questions of neuroscientific interest.
-The first one concerns the frustration of the brain, at both the microscopic and the mesoscopic level.
-For what concerns the first one, the evidence that the median IBI value is positive - irrespectively from the benchmark employed to project our set of multivariate time series - lets us conclude that brain networks are indeed populated by a vast majority of (++)(+-+)( + - + ) triples - the ones reading ()(---)( - - - ) being almost absent.
-Interestingly, although the negative subgraph is more prone to inter-subject variability and less homogeneous than the positive one, the negative connections constituting the frustrated triples are mostly found between the groups of nodes constituting the DMN and the groups of nodes constituting task-positive regions such as the DA, the SM, the SVM and the VIS.
-As already mentioned, these results match the ones in Fox et al.
-(2005) as well as those discussed in Demertzi et al.
-(2022); Greicius et al.
-(2003); Hampson et al.
-(2010).
-Some contributions observe that the empirical level of (im)balance is smaller than the one characterising randomised configurations [^23] It should be, however, noticed that the results obtained in Saberi et al.
-(2021) depend on the chosen threshold, as also explicitly recognised by the authors.
-and deduce that the brain tends to stabilise itself Saberi et al.
-(2021).
-Our results, however, lead to a different conclusion: although it is true that the bSCM returns (slightly) less balanced configurations than the naïve ones, brain networks do not seem to minimise frustration: as confirmed by the distribution of $F$, the values assumed by it across the partitions characterising our and Yeo’s projections are quite larger than zero.
-It is, thus, of interest asking which principle is obeyed by the brain in order to self-organise its functioning.
+Understanding the role played by negative connections within and between brain areas is a topic that has recently gained attention within the neuroscientific literature.
+The method proposed here produces connections characterized by a negative sign as the result of a validation procedure that counts the number of times the corresponding series are found to be discordant and compares it with the prediction of a properly defined benchmark.
+Such a validation scheme allows us to address several neuroscientific questions.
+The first concerns the frustration of the brain at both the microscopic and mesoscopic levels.
+For the first one, the evidence that the median IBI value is positive - regardless of the benchmark employed to project our set of multivariate time series - lets us conclude that brain networks are populated by a non-negligible number of (++)(+-+) and ()(---) triads.
+Interestingly, the negative connections constituting the frustrated triads are mostly found within the groups of subcortical and limbic areas and between the groups of nodes constituting the DMN and those constituting task-positive regions such as the DA, the SM, the SVM, and the VIS; as already mentioned, these results match the ones discussed in Greicius2003functional; Fox2005Human; Hampson2010Functional; Demertzi2022.
+Some contributions observe that the empirical level of (im)balance is smaller than the one characterising randomised configurations [^21] It should be, however, noticed that the results obtained in Saberi2021 depend on the chosen threshold, as explicitly recognised by the authors themselves.
+and deduce that the brain tends to stabilise itself Saberi2021.
+Our results, however, lead to a different conclusion, i.e.
+that brain networks do not seem to minimise frustration: on the contrary, the values assumed by $F$ across Yeo’s partitions and those characterising our projections are quite larger than zero.
+It is, thus, of interest to wonder which principle is obeyed by the brain to self-organise its functioning.
 An answer seems to be provided by fig.
-14: task-based partitions align with those recovered by minimising BIC; in a sense, thus, some sort of optimisation principle seems to be at work at the mesoscopic scale.
-Still, individuating the partition ‘best representing’ a signed brain network remains a challenging task, as different metrics provide different answers; nevertheless, a message whose validity appear as quite general seems to emerge: brain networks align with the (statistical variant) of the Relaxed Balanced Theory, as a vast majority of positive signs populates off-diagonal blocks as well.
+13: Yeo’s partitions align with those returned by BIC minimisation, an advantage of which is that of acting in a maximally ‘agnostic’ way when coming to gather signed connections - differently from modularity maximisation, or closely related approaches Zhan2017, that seek for the group-wise structure best aligning with the Traditional Balance Theory.
+In a sense, thus, some sort of optimisation principle seems to be at work at the mesoscopic scale.
+Still, individuating the partition ‘best representing’ a signed brain network remains a challenging task, as different metrics provide different answers.
+Nevertheless, a message whose validity appear as quite general emerges: brain networks align with the (statistical variant) of the Relaxed Balanced Theory, as negative (positive) signs can populate diagonal (off-diagonal) blocks.
+For what concerns the physiological mechanisms behind the emergence of negative connections, instead, we redirect the interested reader to more technical contributions Harita2024.
 X Conclusions
-A critical, yet often overlooked, aspect of functional brain data is the role played by negative links: frequently ignored due to the interpretative challenges brought along, emerging evidence indicates that negative links play a fundamental role in shaping the (structural and dynamical) properties of resting-state networks, especially those related to self-stabilisation.
-As excluding negative links, thus, means underestimating the brain ability to function properly, we have proposed a validation technique allowing statistically significant signed links (be they positive or negative) to emerge out of the noise accompanying raw data.
-To this aim, two, different benchmarks have been proposed, filtering multivariate time series to a different extent and letting diverse patterns emerge out of the wealth of empirical observations: in particular, while the bSRGM emphasizes structural balance by levelling out the differences between time series, the bSCM preserves their heterogeneity, thus inducing sparser projections that retain more information about the original, bipartite structure.
-In particular, our analysis points out that brain networks are, indeed, frustrated, the patterns induced by signed connections overlapping to a quite large extent with those individuated by Yeo’s task-oriented partitions and returning the picture of a system aligning (better) with the statistical variant of the Relaxed Balance Theory.
-A future research direction is that of employing our algorithm to carry out both a ‘weighted-to-binary’ and a ‘weighted-to-weighted’ analysis, i.e.
-obtain statistically validated binary and weighted projections of multivariate weighted time series, defined by the ‘rule’ that any two nodes sharing a number of (either concordant or discordant) relationships characterised by a significantly large weight are connected by an edge ‘dressed’ with the validated weight.
+A critical, yet often overlooked, aspect of functional brain data is the role played by negative links: frequently ignored due to the interpretative challenges brought along, emerging evidence indicates negative links to be fundamental in shaping resting-state networks - especially the properties affecting self-stabilisation.
+As excluding negative links, thus, means underestimating the brain ability to function properly, we have proposed a validation technique allowing statistically significant signed links to emerge out of the noise accompanying raw data; two, different benchmarks have been employed, filtering multivariate time series to a different extent and letting diverse patterns emerge out of the wealth of empirical observations: while the bSRGM levels out the differences between time series, the bSCM preserves their heterogeneity, thus inducing sparser projections that retain more information about the original structure.
+In particular, our analysis points out that brain networks are, indeed, frustrated, the patterns induced by signed connections overlapping (in some cases, to a quite large extent) with Yeo’s task-based partitions and returning the picture of a system aligning (better) with the statistical variant of the Relaxed Balance Theory.
+A future research direction is that of employing our algorithm to obtain statistically validated binary and weighted projections of multivariate weighted time series, defined by the ‘rule’ that any two nodes involved in a number of relationships characterised by a significantly large weight are connected by an edge ‘dressed’ with such a validated weight.
 XI Acknowledgments
@@
 $S=-\sum_{\mathbf{B}\in\mathbb{B}}P(\mathbf{B})\ln P(\mathbf{B}),$ (108)
-$C_{0}\equiv\langle C_{0}\rangle\equiv 1$ sums up the normalisation condition and the remaining $M-1$ constraints represent proper, topological properties.
+$C_{0}\equiv\langle C_{0}\rangle\equiv 1$ sums up the normalisation condition and the remaining M1M-1 constraints represent proper, topological properties.
 Such an optimisation procedure defines the expression
@@
 Appendix B Numerical optimisation of likelihood functions
-In order to numerically solve the systems of equations defining the bSCM, we can follow the guidelines provided in Vallarano et al.
-(2021): more specifically, we will adapt the iterative recipe provided there to our setting.
+In order to numerically solve the system of equations defining the bSCM, we can follow the guidelines provided in vallarano2021: more specifically, we will adapt the iterative recipe provided there to our setting.
 First, let us notice that, upon posing $x_{i}\equiv e^{-\alpha_{i}}$, $y_{i}\equiv e^{-\gamma_{i}}$, $\forall\>i$ and $z_{t}\equiv e^{-\delta_{t}}$, $w_{t}\equiv e^{-\eta_{t}}$, $\forall\>t$, the system of equations defining the bSCM can be re-written as
@@
 $w_{t}$ $=\frac{\kappa_{t}^{-}(\mathbf{B^{*}})}{\sum_{i=1}^{N}\dfrac{y_{i}}{x_{i}z_{t}+y_{i}w_{t}}}\>\Longrightarrow\>w_{t}^{(n)}=\frac{\kappa_{t}^{-}(\mathbf{B^{*}})}{\sum_{i=1}^{N}\dfrac{y_{i}^{(n-1)}}{x_{i}^{(n-1)}z_{t}^{(n-1)}+y_{i}^{(n-1)}w_{t}^{(n-1)}}},\quad\forall\>t;$ (133)
-in order for our iterative recipe to converge, an appropriate vector of initial conditions needs to be chosen; here, we have opted the following one: $x_{i}=k_{i}^{+}(\mathbf{B}^{*})/\sqrt{L^{+}(\mathbf{B}^{*})}$, $\forall\>i$, $y_{i}=k_{i}^{-}(\mathbf{B}^{*})/\sqrt{L^{-}(\mathbf{B}^{*})}$, $\forall\>i$, $z_{t}=\kappa_{t}^{+}(\mathbf{B}^{*})/\sqrt{L^{+}(\mathbf{B}^{*})}$, $\forall\>t$ and $w_{t}=\kappa_{t}^{-}(\mathbf{B}^{*})/\sqrt{L^{-}(\mathbf{B}^{*})}$, $\forall\>t$.
+in order for our iterative recipe to converge, an appropriate vector of initial conditions needs to be chosen; here, we have opted for the following one: $x_{i}=k_{i}^{+}(\mathbf{B}^{*})/\sqrt{L^{+}(\mathbf{B}^{*})}$, $\forall\>i$, $y_{i}=k_{i}^{-}(\mathbf{B}^{*})/\sqrt{L^{-}(\mathbf{B}^{*})}$, $\forall\>i$, $z_{t}=\kappa_{t}^{+}(\mathbf{B}^{*})/\sqrt{L^{+}(\mathbf{B}^{*})}$, $\forall\>t$ and $w_{t}=\kappa_{t}^{-}(\mathbf{B}^{*})/\sqrt{L^{-}(\mathbf{B}^{*})}$, $\forall\>t$.
 Moreover, we have adopted two, different stopping criteria: the first one is a condition on the Euclidean norm of the vector of differences between the values of the parameters at subsequent iterations, i.e.
-$||\Delta\vec{\theta}||_{2}=\sqrt{\sum_{i=1}^{N}(\Delta\theta_{i})^{2}}\leq 10^{-8}$; the second one is a condition on the maximum number of iterations of our iterative algorithm, set to $10^{3}$.
+$\Delta\vec{\theta}||_{2}=\sqrt{\sum_{i=1}^{N}(\Delta\theta_{i})^{2}}\leq 10^{-8}$; the second one is a condition on the maximum number of iterations of our iterative algorithm, set to $10^{3}$.
 The accuracy of our method in estimating the constraints has been evaluated by computing the maximum absolute error (MAE), defined as
@@
 (i.e. as the infinite norm of the relative difference between the vector of the empirical values of the constraints and the vector of their expected values).
-Table 1, providing information on the basic statistics of 555, illustrative time series coming from the HCP-Young Adult dataset shows the time employed by our algorithm to converge as well as its accuracy in reproducing the constraints defining the bSCM: our method is, overall, fast and accurate, as the absolute errors are around $10^{-3}$ and the relative errors are around $10^{-6}$, the time employed to achieve such an accuracy being always less than a minute.
-Basic statistics of time series bSCM $N$ $T$ $B^{+}$ $B^{-}$ $c^{+}$ $c^{-}$ MAE MRE Time (s) HCP106016 116 2.400 142.407 135.993 0.512$\simeq$ 0.512≃ 0.512 0.488$\simeq$ 0.488≃ 0.488 $\simeq 1.10\cdot 10^{-3}$ $\simeq 9.94\cdot 10^{-7}$ 28.8$\simeq$ 28.8≃ 28.8 HCP857263 116 2.400 136.376 142.024 0.490$\simeq$ 0.490≃ 0.490 0.510$\simeq$ 0.510≃ 0.510 $\simeq 1.10\cdot 10^{-3}$ $\simeq 9.90\cdot 10^{-7}$ 33.3$\simeq$ 33.3≃ 33.3 HCP125525 116 2.400 139.054 139.346 0.499$\simeq$ 0.499≃ 0.499 0.501$\simeq$ 0.501≃ 0.501 $\simeq 1.26\cdot 10^{-3}$ $\simeq 1.17\cdot 10^{-6}$ 37.8$\simeq$ 37.8≃ 37.8 HCP100408 116 2.400 137.804 140.596 0.495$\simeq$ 0.495≃ 0.495 0.505$\simeq$ 0.505≃ 0.505 $\simeq 1.30\cdot 10^{-3}$ $\simeq 1.15\cdot 10^{-6}$ 37.8$\simeq$ 37.8≃ 37.8 HCP130316 116 2.400 137.196 141.204 0.493$\simeq$ 0.493≃ 0.493 0.507$\simeq$ 0.507≃ 0.507 $\simeq 1.97\cdot 10^{-3}$ $\simeq 1.89\cdot 10^{-6}$ 37.7$\simeq$ 37.7≃ 37.7
-Table 1: Performance of our fixed-point algorithm to solve the system of equations defining the bSCM on 555, illustrative time series coming from the HCP-Young Adult dataset (each pedex refers to a subject).
-$N$ is the number of brain regions, $T$ is the duration of the related time series, $B^{+}$ is the overall number of positive values, $B^{-}$ is the overall number of negative values, $c^{+}=B^{+}/(N\cdot T)$ is the density of positive values and $c^{-}=B^{-}/(N\cdot T)$ is the density of negative values.
+Table 1, providing information on the basic statistics of $5$, illustrative time series coming from the HCP-Young Adult dataset shows the time employed by our algorithm to converge as well as its accuracy in reproducing the constraints defining the bSCM: our method is, overall, fast and accurate, as the absolute error is steadily around $10^{-3}$ and the relative error is steadily around $10^{-6}$, the time employed to achieve such an accuracy being always less than a minute.
+Basic statistics of time series bSCM NN $T$ $^{+} BB^{-}$ $^{+} cc^{-}$ MAE MRE Time (s) HCP100408 116 2.400 137.804 140.596 $\simeq 0.495$ $\simeq 0.505$ $\simeq 1.30\cdot 10^{-3}$ $\simeq 1.15\cdot 10^{-6}$ $\simeq 37.8$ HCP106016 116 2.400 142.407 135.993 $\simeq 0.512$ $\simeq 0.488$ $\simeq 1.10\cdot 10^{-3}$ $\simeq 9.94\cdot 10^{-7}$ $\simeq 28.8$ HCP125525 116 2.400 139.054 139.346 $\simeq 0.499$ $\simeq 0.501$ $\simeq 1.26\cdot 10^{-3}$ $\simeq 1.17\cdot 10^{-6}$ $\simeq 37.8$ HCP130316 116 2.400 137.196 141.204 $\simeq 0.493$ $\simeq 0.507$ $\simeq 1.97\cdot 10^{-3}$ $\simeq 1.89\cdot 10^{-6}$ $\simeq 37.7$ HCP857263 116 2.400 136.376 142.024 $\simeq 0.490$ $\simeq 0.510$ $\simeq 1.10\cdot 10^{-3}$ $\simeq 9.90\cdot 10^{-7}$ $\simeq 33.3$
+Table 1: Performance of our fixed-point algorithm to solve the system of equations defining the bSCM on 55, illustrative time series coming from the HCP-Young Adult dataset (each pedex refers to a subject): NN is the number of brain regions, $T$ is the duration of the related time series, $B^{+}$ is the overall number of positive values, $B^{-}$ is the overall number of negative values, $c^{+}=B^{+}/(N\cdot T)$ is the density of positive values and $c^{-}=B^{-}/(N\cdot T)$ is the density of negative values.
 Our fixed-point algorithm leads to MAE values around $10^{-3}$ and MRE values of the order of $10^{-6}$ in less than forty seconds.
 Appendix C Degree distributions of cortical and subcortical areas
-As anticipated in the main text, subcortical areas tend to have a larger negative degree than cortical areas. This is true in both a subject-wise (see fig. 16) and a pooled fashion (see fig. 17)
-Figure 16: Complementary cumulative distribution functions (CCDFs) of positive ($k^{+}$) and negative ($k^{-}$) degrees for cortical and subcortical areas, per benchmark.
-Analyses have been carried out in a subject-wise fashion (i.e.
-for subjects 106016106016106016 and 857263857263857263).
-Figure 17: Complementary cumulative distribution functions (CCDFs) of positive ($k^{+}$) and negative ($k^{-}$) degrees for cortical and subcortical areas, per benchmark.
+As anticipated in the main text, subcortical areas tend to have a larger negative degree than cortical areas.
+Figure 16: Cumulative distributions of positive ($k^{+}$) and negative ($k^{-}$) degrees for cortical and subcortical areas, per benchmark.
 Analyses have been carried out in a pooled fashion (i.e.
-across the 100100100 subjects considered for the present analysis).
+across the $100$ subjects considered for the present analysis).